### v2

# The Statistical Fairness-Accuracy Frontier

Abstract
Machine learning models must balance accuracy and fairness, but these goals often conflict, particularly when data come from multiple demographic groups. A useful tool for understanding this trade-off is the fairness-accuracy (FA) frontier, which characterizes the set of models that cannot be improved in either fairness or accuracy without worsening the other. Prior analyses of the FA frontier provide a full characterization under the assumption of complete knowledge of population distributions—an unrealistic ideal. We study the FA frontier in the finite-sample regime, showing how it deviates from its population counterpart and quantifying the worst-case gap between them. In particular, we derive minimax-optimal estimators that depend on the designer’s knowledge of the covariate distribution. For each estimator, we characterize how finite-sample effects asymmetrically impact each group’s risk, and identify optimal sample allocation strategies. Our results transform the FA frontier from a theoretical construct into a practical tool for policymakers and practitioners who must often design algorithms with limited data.
1 Introduction
Across domains where predictive models guide consequential decisions—from lending and hiring to healthcare and education—society expects systems to be both accurate and fair. In supervised learning with multiple demographic groups, these objectives are often in tension: models that minimize overall error can yield unequal performance across groups, while interventions that reduce disparities may incur accuracy losses. A central, practically relevant challenge arises when a single predictive model must serve multiple groups, whether due to legal, logistical, or normative constraints. For instance, in the United States, the Equal Credit Opportunity Act (ECOA) prohibits lending decisions based on protected attributes such as race or sex and is commonly interpreted to disallow the use of different predictive models for different groups. Similarly, in hiring, automated screening systems typically apply a single predictive model to all applicants, since Title VII of the Civil Rights Act imposes legal constraints that, along with operational challenges, discourage the use of group-specific models (see raghavan2020mitigating for a discussion of U.S. employment discrimination law and the guidelines on “disparate treatment”).
In such cases—where group-specific models are disallowed or impractical—many existing models and algorithms are known to exhibit disparate performance across groups [bartlett2022consumer, fuster2022predictably, dastin2022amazon]. Thus, it becomes critical to characterize the best possible compromises between fairness and accuracy. A principled way to approach this tension is through study of the fairness–accuracy (FA) frontier. The FA frontier consists of models for which it is impossible to improve fairness without sacrificing accuracy for at least one group, or to improve a group’s accuracy without harming fairness or the other group’s accuracy. This frontier allows decision-makers to visualize the explicit trade-offs and select a model that aligns with their fairness and performance objectives.
Recent work, notably by liang2021algorithm, provides a valuable characterization of the FA frontier under the assumption that the population distributions for all groups are known exactly. In most real-world scenarios, however, this assumption is unlikely to hold—we are unlikely to possess such a distribution or to have access to the entire population. Instead, in practice, we must base our inferences and decisions regarding the FA frontier on a finite number of samples drawn from the underlying population. This paper tackles this finite-sample regime directly. We study how the empirical FA frontier deviates from its population counterpart, characterize the worst-case gap between them, and develop optimal estimators and sampling strategies. In short, we recast the FA frontier from an idealized object into a practical statistically-based mechanism design tool.
[…]
Figure 1: (a) Population FA frontier: The blue shaded area represents the set of risk pairs that are achievable by some linear model, and $\beta_{r}$ and $\beta_{b}$ are the error-minimizing models for groups $r$ and $b$, respectively. $\beta_{\text{fair}}$ is the error-equalizing model, which falls between $\beta_{r}$ and $\beta_{b}$ in the group-balanced case. The FA frontier corresponds to the purple region along the boundary of the feasible set of error pairs. (b) For any $\lambda\in[0,1]$, $\beta_{\lambda}$ is the first point of tangency between the line $\lambda x+(1-\lambda)y=c$ and the FA frontier as we increase $c>0$. This point moves along the frontier from $\beta_{r}$ to $\beta_{b}$ as $\lambda$ ranges from $1\to 0$.
Consider a two-group linear regression setting, with groups labeled “red” ($r$) and “blue” ($b$). For each individual in a group $g\in\{r,b\}$, the outcome variable is modeled as the inner product of the individual’s covariate vector and a group-specific parameter $\beta_{g}\in\mathbb{R}^{d}$, plus a mean-zero random noise term. The two groups may differ in both their underlying parameters, i.e., $\beta_{r}\neq\beta_{b}$, and in their covariate distributions. For any linear prediction model parameterized by $\beta$, we measure its group-specific accuracy as the risk—the expected squared loss under the distribution of group $g$. We measure fairness as the absolute difference in risk between the two groups, which is the equalized loss criterion [zhang2019group, khalili2023loss]. Following liang2021algorithm, we visualize the FA frontier in a two-dimensional plot, with the risks for the red and blue groups on the two axes. In this case, as shown in Figure 1(a), the FA frontier (under full knowledge of the distribution) traces a portion of the boundary of the set of all risk pairs (across the two groups) that are attainable by some linear model. For homoskedastic linear regression, the FA frontier is group-balanced in the sense of liang2021algorithm: the most fair point lies between the two group-optimal endpoints. Hence, as displayed in Figure 1(a), the two endpoints of the FA frontier lie at the group-$r$ optimum $\beta_{r}$ and the group-$b$ optimum $\beta_{b}$.
In this case, we can parametrize the FA frontier using a parameter $\lambda\in[0,1]$: for each $\lambda$, we define an objective function as a convex combination of group-specific risks, assigning weight $\lambda$ to the red group and $1-\lambda$ to the blue group. The pair of risks corresponding to the minimizer of this objective, denoted by $\beta_{\lambda}$, lies on the FA frontier and represents the first point of tangency between the frontier and a line of the form $\lambda x+(1-\lambda)y=c$ as $c>0$ increases, as illustrated in Figure 1(b). By varying $\lambda$ continuously from 0 to 1, we trace out the entire FA frontier.
The choice of $\lambda$ is a central design decision shaped by the model’s legal and ethical context, reflecting the social planner’s preferences over group errors. To illustrate this, suppose 90% of the population belongs to the red group. A natural choice would then be to set $\lambda=0.9$. However, if we are training a predictive model for patients or for hiring across different demographics, we might want to ensure that the smaller group is not overshadowed. In that case, we could increase its weight and even assign equal importance to both groups, i.e., set $\lambda=0.5$. This approach, often referred to as “weighting by inverse class frequency,” is common in the fairness literature [cui2019class, steininger2021density]. There are also smoothed variants that interpolate between these extremes, offering a context-dependent balance between accuracy and fairness [mikolov2013distributed]. In summary, the choice of $\lambda$ ultimately rests with the designer, and the role of the FA frontier is to ensure that, whatever choice is made, the resulting algorithm is optimal for that setting.
[…]
Figure 2: (a) Finite-sample estimation: The error pair corresponding to the empirical estimator $\widehat{\beta}_{\lambda}$ lies on the line $\lambda x+(1-\lambda)y=c^{\prime}$, where $c-c^{\prime}$ is the excess risk. (b) Asymmetric group-wise impact: The displacement from $\beta_{\lambda}$ to $\widehat{\beta}_{\lambda}$ decomposes into a vertical change in the blue-group error and a horizontal change in the red-group error; their unequal magnitudes show that the estimator affects the two groups asymmetrically.
Characterizing the FA frontier and a target model $\beta_{\lambda}$ requires complete knowledge of both the covariate distributions and the underlying parameters, which is not possible in a finite-sample regime. Further, as we will discuss, any finite-sample estimator of $\beta_{\lambda}$ will yield a risk pair that is strictly inside the achievable set. This raises a natural question: Given only $n_{r}$ and $n_{b}$ samples from groups $r$ and $b$, respectively, what is the best performance we can guarantee? To answer this, we must first define what “best” means.
A natural way to measure the performance of a finite-sample estimator is through its worst-case excess risk. Consider again lines of the form $\lambda x+(1-\lambda)y=c$; for a finite-sample estimator, define its excess risk as the amount by which $c$ must be increased (i.e., shifting the line upward and to the right) so that it intersects an attainable risk pair, in the worst-case sense over a family of distributions. This idea is illustrated in Figure 2(a). Our objective is to minimize the worst-case expected excess risk—where the expectation is taken over the randomness in the $n_{r}+n_{b}$ samples. Minimizing the worst-case expected excess risk corresponds to achieving the minimax rate, a standard benchmark for evaluating the finite-sample performance of statistical estimators.
We summarize our main contributions. First, we show that the minimax-optimal estimator differs in the setting in which we know the covariance matrices of the covariate but the parameters $\beta_{r}$ and $\beta_{b}$ are unknown and the setting in which neither the covariance matrices nor the parameters are known. We refer to the first scenario as the known covariance case and the second as the unknown covariance case. In the known covariance case, we establish that the minimax-optimal estimator preserves the structure of the oracle solution $\beta_{\lambda}$ (and depends explicitly on the true covariance matrices of the covariates in the two groups) while replacing the unknown group-specific parameters with their empirical analogs. In particular, we derive an upper bound on the excess risk of this estimator and a matching lower bound (see Theorem 1). We further simplify our bound under a “small-ball” condition, which—roughly speaking—bounds the lower tail of the sample covariance spectrum (and holds, for example, for Gaussian distributions). In this case, we show that the excess risk of the optimal estimator can be decomposed into a weighted sum of two group-specific error terms. Each term grows with the norm of the group’s covariance matrix and decreases with its sample size. This implies that higher-variance covariates require more samples to achieve the same accuracy. More importantly, increasing the sample size of only one group cannot reduce the overall error beyond a certain limit.
In the unknown covariance case, we consider the ordinary least squares (OLS) estimator, which minimizes the empirical version of the objective function (the weighted average of the risks of the two groups). We show that its excess risk consists of two components, which we refer to as the variance and the bias. The variance term arises from the estimation of the parameters $\beta_{r}$ and $\beta_{b}$, even when the covariance matrices are known, and thus mirrors the error from the known covariance case. The bias term, in contrast, represents the error incurred from the estimation of the covariance matrices and remains even when the group-wise parameters are known (see Proposition 3).
For subgaussian covariates satisfying the small-ball condition, we derive upper bounds for both the bias and variance terms. Our upper bound for the variance term naturally matches the bound in the known covariance case. For the bias term, we establish an upper bound that shows how it depends differently on the sample size, the covariance matrices, and the group heterogeneity which is defined as $\|\beta_{r}-\beta_{b}\|$. Moreover, for both terms, we establish matching lower bounds (up to constant factors), showing that the excess risk of any estimator must contain both the bias and variance terms, and that no estimator can improve upon the OLS estimator by more than a constant factor (see Theorem 2 and Theorem 3).
Our bounds also provide an optimal sampling rule for a fixed total sampling budget, $n_{r}+n_{b}$. In the known-covariance case, we show that the optimal allocation of the sampling budget depends on the norm of the covariance matrix as well as the weight of that group’s risk in the overall loss ($\lambda$ versus $1-\lambda$), with the optimal sample size increasing as either quantity grows. This result is a direct analogue of Neyman’s classical allocation in stratified sampling [neyman1934representative], where sample sizes are chosen in proportion to stratum variability and importance. The optimal sampling rule for the unknown covariance case is more nuanced. When the variance term dominates (e.g., when $\|\beta_{r}-\beta_{b}\|$ is small), the same allocation rule as in the known covariance case remains near-optimal. When heterogeneity is large and the bias term dominates, however, the optimal design shifts toward balancing the two sample sizes, $n_{r}\approx n_{b}$, to reduce covariance-estimation bias. Which regime governs depends on $\|\beta_{r}-\beta_{b}\|$, clarifying how problem structure should inform data collection when targeting a specific point on the FA frontier.
The results discussed thus far provide a complete characterization of how a given point on the FA frontier, parameterized by $\lambda$, is affected by the transition to the empirical setting. This characterization is particularly useful when the desired trade-off rate is known ex ante. However, policymakers may be interested in the entire FA frontier, as it illustrates the full range of feasible trade-offs. A key challenge in this case is that, because only a single dataset is available to derive all points on the frontier, the estimation errors are correlated across values of $\lambda$. We address this issue in Section 6 by deriving a uniform, high-probability bound on empirical error across the frontier; we simultaneously control how much the empirical pair of group risks deviates from the corresponding population risks on the frontier. This allows us to augment the empirical FA frontier with a high-probability confidence band for the entire frontier. We further show that the width of this confidence band decreases with sample size, being proportional to the maximum of the square roots of the inverse sample sizes across the two groups (see Theorem 4). This result provides a rigorous foundation for quantifying the uncertainty in policy decisions based on the empirical FA frontier.
Finally, while the preceding results bound the excess risk of the minimax-optimal empirical estimator with respect to the (weighted) average objective function, it is natural to ask: How does this estimator affect each group’s individual risk? As shown in Figure 2(b), we find that the empirical estimator does not affect the two groups equally. In particular, when the covariance is known, we show that the ratio of the increases in excess risk across the two groups is proportional to the ratio of the norms of their covariance matrices. In the unknown covariance case, beyond this difference arising from covariance scaling, the bias of the OLS estimator introduces an additional cross term. This term can shift the risks in opposite directions between groups and depends asymmetrically on $\lambda$ for the two groups. Interestingly, in both cases, the differing impact of the estimator on the two groups does not stem from differences in their sample sizes. These findings suggest that the fairness–accuracy trade-off implied by the empirical estimator may differ substantially in practice from the trade-off defined at the population level.
1.1 Related Work
We build on liang2021algorithm, who formalize the FA Pareto frontier for assessing fairness-accuracy trade-offs under the assumption of full knowledge of the covariate and outcome distributions. liu2024inference relax the full-knowledge assumption and introduce a consistent estimator of the FA frontier, deriving the asymptotic distribution of the estimator and using it to design statistical tests for properties such as the gap between a given algorithm and the fairest alternative. auerbach2024testing further contribute a statistical test for determining whether a model achieving a given accuracy admits a Pareto improvement in fairness. We likewise address the practical challenges of analyzing the FA frontier from finite data, but take a fully non-asymptotic approach, providing finite-sample guarantees for estimators that target a social planner’s fairness-accuracy preferences.
Our work contributes to the literature on fairness in regression, a topic that has traditionally received less attention than fair classification [see, e.g., donini2018empirical]. The most closely related work is that of chzhen2022minimax, who develop a minimax-optimal fair regression estimator for an explicit fairness constraint based on the Wasserstein distance between the loss distributions across groups. Beyond this, several other works have introduced various notions of fairness for the regression setting and have proposed corresponding optimal estimation procedures [calders2013controlling, berk2017convex, fitzsimons2019general, agarwal2019fair, chzhen2020fair, chzhen2020fairb, oneto2020general] and for estimating the minimal utility cost associated with achieving fairness in regression [zhao2021costs].
A growing body of work bridges perspectives on fairness-accuracy trade-offs in computer science and statistics with analyses of social welfare [hu2020fair, liang2024algorithmic, gans2024algorithmic, rosenfeld2025machine]. Our results on the FA frontier can also be interpreted in this context if we view each individual’s utility as the risk associated with their group. More broadly, in the algorithmic fairness literature, it is widely recognized that no single fairness criterion can capture the diverse preferences that decision-makers may have over these trade-offs [casacuberta2023augmenting, verma2018fairness, kleinberg2016inherent]. Economists make this point in a similar way by documenting the heterogeneity in how individuals weigh equity against efficiency [capozza2024should, lockwood2015gustibus]. Our framework builds on this connection by adopting a flexible, weight-based formulation that accommodates this kind of heterogeneity [saez2016generalized, hendren2020measuring, kasy2021fairness].
Finally, much attention has been devoted to the fundamental question of when fairness and accuracy are in conflict and when they are aligned [corbett2017algorithmic, wick2019unlocking, blum2019recovering, liang2021algorithm, andersen2020big, dutta2020there, cooper2021emergent]. In settings where these objectives are at odds, a variety of methods have been developed to identify Pareto improvements, where either fairness or accuracy can be improved while keeping the other one unchanged. One line of work designs algorithms for computing fairer alternatives to a given model for specific applications [fish2016confidence, coston2021characterizing, wei2022fairness, chohlas2024learning, dehdashtian2024utility]. A closely related line of work draws on the legal notion of a Less Discriminatory Alternative, or a fairer model that maintains the original model’s accuracy [gillis2024operationalizing, laufer2025constitutes, samad2025selecting]. Others take the reverse approach, focusing on accuracy improvements subject to fairness guarantees [agarwal2018reductions, wang2023aleatoric], or, more generally, methods for balancing multiple competing objectives in learning [rolf2020balancing, martinez2020minimax].
2 Framework
We consider a population where each individual is represented by a covariate vector $X\in\mathcal{X}\subset\mathbb{R}^{d}$, a response (outcome) variable $Y\in\mathbb{R}$, and a group label $g\in\mathcal{G}:=\{r,b\}$ (e.g., red, blue). For each group $g$, we observe $n_{g}$ independent data points, $\mathcal{S}_{g}:=\{(X_{i}^{(g)},Y_{i}^{(g)})\}_{i=1}^{n_{g}}$, each drawn from the conditional distribution $P(X,Y|G=g).$
Let $\mathcal{H}$ be a class of prediction functions $f:\mathcal{X}\to\mathbb{R}$. We quantify the error of a predictor $f(X)$ in estimating the output variable $Y$ using a nonnegative loss function, $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$. Specifically, the loss incurred on a single example is given by $\ell(f(X),Y)$. The group-wise population risk and its empirical counterpart for prediction function $f\in\mathcal{H}$ are given, respectively, by
$\mathcal{R}_{g}(f)\coloneqq\mathbb{E}_{(X,Y)\sim P(\cdot|G=g)}[\ell(f(X),Y)],\quad\widehat{\mathcal{R}}_{g}(f)\coloneqq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}\ell\left(f\left(X_{i}^{(g)}\right),Y_{i}^{(g)}\right).$
Each prediction function $f\in\mathcal{H}$ induces a pair of risks $(\mathcal{R}_{r}(f),\mathcal{R}_{b}(f))$. Here, $\mathcal{R}_{r}(f)$ and $\mathcal{R}_{b}(f)$ represent the prediction accuracy of $f$ over groups $r$ and $b$, respectively.
Following liang2021algorithm, we measure the fairness of a prediction function $f$ using the risk disparity $|\mathcal{R}_{r}(f)-\mathcal{R}_{b}(f)|$, which quantifies the difference in performance of the model between the two groups. For regression tasks with a continuous outcome variable YY—the primary focus of this paper—this notion coincides with the definition of Equalized Loss proposed by zhang2019group. [^1] For binary classifiers, i.e., when the outcome variable $Y$ is binary, liang2021algorithm provide a detailed discussion on how this definition relates to well-known fairness notions in the binary setting such as Equalized Odds [hardt2016equality].
2.1 Fairness-Accuracy Frontier
To investigate the fairness–accuracy trade-off, we begin by examining how one can maneuver between group accuracies and fairness by changing the prediction function. More formally, we start by defining the set of achievable population risk pairs over the class $\mathcal{H}$, given by
$\mathcal{E}(\mathcal{H})\coloneqq\{(\mathcal{R}_{r}(f),\mathcal{R}_{b}(f)):f\in\mathcal{H}\}.$
We next adopt the definition of the fairness-accuracy (FA) Pareto frontier from liang2021algorithm. To do so, we first revisit the definition of FA dominance.
Definition 1 (Fairness-Accuracy (FA) Dominance).
We say that a function $f^{\prime}\in\mathcal{H}$ FA-dominates a function $f\in\mathcal{H}$, denoted by $f^{\prime}\succ f$, if $\mathcal{R}_{r}(f^{\prime})\leq\mathcal{R}_{r}(f)$, $\mathcal{R}_{b}(f^{\prime})\leq\mathcal{R}_{b}(f)$, and $|\mathcal{R}_{r}(f^{\prime})-\mathcal{R}_{b}(f^{\prime})|\leq|\mathcal{R}_{r}(f)-\mathcal{R}_{b}(f)|$, with at least one of these three inequalities being strict.
In other words, a function $f^{\prime}$ FA-dominates a function $f$ if it achieves no higher risk on either group and no greater risk disparity between groups, with strict improvement for at least one. The FA frontier is then defined as the subset of achievable risk pairs that are not FA-dominated by any other point.
Definition 2 (Fairness-Accuracy (FA) Frontier).
The FA frontier, denoted by $\mathcal{F}(\mathcal{H})$, is defined as:
$\mathcal{F}(\mathcal{H})\coloneqq\left\{\left(\mathcal{R}_{r}(f),\mathcal{R}_{b}(f)\right)\in\mathcal{E}(\mathcal{H}):\nexists f^{\prime}\in\mathcal{H}:f^{\prime}\succ f\right\}.$
Insight into the shape of the fairness-accuracy (FA) frontier can be obtained by focusing on three key points: the two points corresponding to the prediction functions that achieve the best accuracy for each group, and the point that achieves the best fairness guarantee. More formally, for any $g\in\mathcal{G}$, let $f_{g}$ be defined as
$f_{g}:=\arg\min_{f\in\mathcal{H}}\mathcal{R}_{g}(f),$ (1)
and define $f_{f}$ as
$f_{f}:=\arg\min_{f\in\mathcal{H}}|\mathcal{R}_{r}(f)-\mathcal{R}_{b}(f)|.$ (2)
It is straightforward to verify that the pairs of risks corresponding to all three functions lie on the FA frontier. In fact, the endpoints of the FA frontier are two of the three points associated with $\{f_{r},f_{b},f_{f}\}$. liang2021algorithm define a distribution $P$ as group-skewed if the point corresponding to $f_{f}$ lies at one of the endpoints of the frontier. In contrast, $P$ is said to be group-balanced if the endpoints correspond to $f_{r}$ and $f_{b}$, and the most fair point $f_{f}$ lies between them. An illustration of group-skewed and group-balanced distributions is provided in Figure 3.
[…]
(a) Group-Balanced
[…]
(b) Group-Skewed
Figure 3: Group-Balanced vs. Group-Skewed. To avoid cluttering the figure, we label points by their corresponding predictors; for example, $f_{r}$ denotes the point $(\mathcal{R}_{r}(f_{r}),\mathcal{R}_{b}(f_{r}))$.
In the group-skewed scenario, the points associated with the risk pairs of both $f_{r}$ and $f_{b}$ lie on one side of the $\mathcal{R}_{b}=\mathcal{R}_{r}$ line. For instance, in Figure 3(b), $\mathcal{R}_{r}(f_{b})$ is lower than $\mathcal{R}_{b}(f_{b})$ even though $f_{b}$ is the optimal predictor for group $b$. In contrast, in the group-balanced case, the fair predictor $f_{f}$ falls between $f_{b}$ and $f_{r}$, indicating that each group achieves a lower risk at its respective optimal predictor. As a result, achieving fairness may require trading off some accuracy between the two groups. In the next section, we show that the regression problem, under the assumption of equal noise variance, falls into the group-balanced category. Therefore, we focus on this setting for the remainder of the paper.
For the group-balanced setting, the FA frontier can be traversed by varying a parameter $\lambda\in[0,1]$ which represents the designer’s preference in trading off between fairness and accuracy for each group. In particular, given a weight $\lambda$, the decision-maker seeks a decision policy $f_{\lambda}$ that achieves the desired tradeoff in errors between groups. This corresponds to solving the weighted risk minimization:
$f_{\lambda}=\arg\min_{f\in\mathcal{H}}\mathcal{R}_{\lambda}(f)\quad\text{with}\quad\mathcal{R}_{\lambda}(f)=\lambda\mathcal{R}_{r}(f)+(1-\lambda)\mathcal{R}_{b}(f).$ (3)
Here, $\lambda$ encodes the relative cost of errors affecting group $r$ to those affecting group $b$. The extreme cases $\lambda=1$ and $\lambda=0$ recover the group-optimal decision policies $f_{r}$ and $f_{b}$, respectively. In general, $f_{\lambda}$ is the first point in $\mathcal{E}(\mathcal{H})$ that intersects the line $\lambda\mathcal{R}_{r}+(1-\lambda)\mathcal{R}_{b}=c$ as $c$ increases—that is, as the line shifts upward and to the right. See Figure 4(a) for an illustration. This formulation thus allows us to target a point on the FA frontier and trade off between fairness and accuracy.
[…]
(a) Illustration of $f_{\lambda}$
[…]
(b) Empirical FA Frontier
Figure 4: Illustration of FA frontier, parametrized by $\lambda$, and its empirical version.
2.2 Empirical Fairness-Accuracy Frontier
In practice, the true distribution $P$ is unknown, so $f_{\lambda}$ is not available. Instead, we begin by obtaining an empirical estimate $\widehat{f}_{\lambda}\in\mathcal{H}$ based on the dataset $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. The error associated with this estimate, known as the excess risk, is given by $\mathcal{R}_{\lambda}(\widehat{f}_{\lambda})-\mathcal{R}_{\lambda}(f_{\lambda})$. It is nonnegative, since $f_{\lambda}$ minimizes $\mathcal{R}_{\lambda}(\cdot)$.
Note that the corresponding pair of risks $(\mathcal{R}_{r}(\widehat{f}_{\lambda}),\mathcal{R}_{b}(\widehat{f}_{\lambda}))$ will always lie in $\mathcal{E}(\mathcal{H})$, and may fall strictly in its interior. Moreover, for any estimator $\widehat{f}_{\lambda}$, the corresponding excess risk quantifies how much the frontier is pushed inward in the direction orthogonal to the line $\lambda\mathcal{R}_{r}+(1-\lambda)\mathcal{R}_{b}=\text{constant}$. An illustration of this phenomenon is shown in Figure 4(b).
Now, assuming that the distribution $P$ belongs to a class of distributions $\mathcal{P}$, we can proceed without knowledge of $P$ by finding an estimator $\widehat{f}_{\lambda}$ that minimizes the worst-case excess risk. More formally, we aim to solve the following problem:
$\inf_{\widehat{f}_{\lambda}\in\mathcal{H}}\sup_{P\in\mathcal{P}}\mathbb{E}\left[\mathcal{R}_{\lambda}(\widehat{f}_{\lambda})-\mathcal{R}_{\lambda}(f_{\lambda})\right],$ (4)
where the expectation is taken with respect to the draw of the dataset $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. In the remainder of the paper, we study this minimax problem in the setting of linear regression. We present a lower bound for the worst-case excess risk in (4), and develop an estimator that matches this lower bound with respect to many of the problem’s parameters. We also discuss how the risks for the two groups change when this optimal estimator is used.
2.3 Notation
For a vector $v$, denote by $\left\lVert v\right\rVert$ its Euclidean norm. For a positive-semidefinite (PSD) matrix $A$, we denote its trace by $\operatorname{Tr}\left(A\right)$, its minimum and maximum eigenvalues by $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$, respectively, and its operator (spectral) norm by $\left\lVert A\right\rVert$. For any matrix $A$ and vector $v$, define $\left\lVert v\right\rVert_{A}^{2}\coloneqq v^{\top}Av$. We write $I_{d}$ for the $d\times d$ identity matrix and use the Loewner order $A\preceq B$ to mean that BAB-A is PSD. For two probability measures $P,Q$, we write $\left\lVert P,Q\right\rVert_{\text{TV}}\coloneq\frac{1}{2}\int|dP-dQ|$ and $D_{\text{KL}}(P\;\|\;Q)\coloneq\int\log\frac{dP}{dQ}dP$ for their total variation and Kullback-Leibler divergences, respectively. For a subgaussian random variable $Z$, we define its $\psi_{2}$-norm as $\left\lVert Z\right\rVert_{\psi_{2}}\coloneqq\inf\{t>0:\mathbb{E}[\exp(Z^{2}/t^{2})]\leq 2\}$. For two functions $g(x)$ and $h(x)$, we write $g\lesssim h$ or $g=\mathcal{O}(h)$ if there exist a constant $C$ (independent of $x$) and a threshold $x_{0}$ such that $g(x)\leq Ch(x)$ for all $x\geq x_{0}$. Finally, $d_{h}$ denotes the Hamming distance on the hypercube and $\Xi^{d}\coloneq\{-1,1\}^{d}$ is the $d$-dimensional signed hypercube.
3 Linear Model Class
To study theoretical bounds on the empirical FA frontier, we restrict attention to the class of linear models:
Definition 3 (Linear Model Class).
The linear model class $\mathcal{P}_{\text{linear}}(\sigma^{2})$ consists of all distributions in which the response variable $Y$, conditional on covariate $X$ and group $g\in\{r,b\}$, satisfies
$Y=X^{\top}\beta_{g}+\varepsilon_{g},$
for some parameter $\beta_{g}\in\mathbb{R}^{d}$ and a random variable $\varepsilon_{g}$ such that $\mathbb{E}[\varepsilon_{g}|X]=0$ and $\mathbb{E}[\varepsilon_{g}^{2}|X]=\sigma^{2}$. If we further assume that the distribution of the covariate $X$ given group identity $g$ is known and equal to $P_{X}^{g}$, then we denote this subclass of $\mathcal{P}_{\text{linear}}(\sigma^{2})$ by $\mathcal{P}_{\text{linear}}(P_{X}^{r},P_{X}^{b},\sigma^{2})$.
A canonical example of a distribution in $\mathcal{P}_{\text{linear}}(\sigma^{2})$ is the well-specified linear model with homoskedastic Gaussian noise, $\varepsilon_{g}\sim\mathcal{N}(0,\sigma^{2})$, independent of $X$.
In what follows, we take $\mathcal{H}$ as the class of linear models, i.e., $\{f_{\beta}(x)=x^{\top}\beta:\beta\in\mathbb{R}^{d}\}$. In addition, we consider the squared loss function $\ell(f_{\beta}(X),Y)=(\beta^{\top}X-Y)^{2}$, and also simplify the notation of the risks $\mathcal{R}_{g}(f_{\beta})$ and $\mathcal{R}_{\lambda}(f_{\beta})$ to $\mathcal{R}_{g}(\beta)$ and $\mathcal{R}_{\lambda}(\beta)$, respectively. We write $\mathbb{E}_{g}$ and $\mathbb{P}_{g}$ to denote expectation and probability, respectively, under the conditional distribution $P(X,Y|G=g)$.
The group-$g$ population covariance is defined as $\Sigma_{g}\coloneqq\mathbb{E}_{g}[XX^{\top}]$, and the corresponding sample covariance is
$\widehat{\Sigma}_{g}\coloneqq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}^{(g)}X_{i}^{(g)\top}.$
For the subsequent analysis, we also find it helpful to define the cross-moment between $X$ and $Y$ for each group, along with its empirical counterpart:
$\nu_{g}\coloneqq\mathbb{E}_{g}\left[XY\right]=\Sigma_{g}\beta_{g},\quad\widehat{\nu}_{g}\coloneqq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}Y_{i}^{(g)}X_{i}^{(g)}.$ (5)
The following result shows that this model falls under the group-balanced scenario described in Section 2. The proof is provided in Section A.2.1.
Lemma 1 (Group-Balanced Structure).
The linear model described above exhibits a group-balanced structure. That is, each group’s risk-minimizing predictor achieves (weakly) lower prediction error on its own group than on the other:
$\mathcal{R}_{r}(\beta_{r})\leq\mathcal{R}_{b}(\beta_{r}),\quad\mathcal{R}_{r}(\beta_{b})\geq\mathcal{R}_{b}(\beta_{b}).$
3.1 Assumptions
We introduce the following assumptions to ensure that population-level quantities and their empirical estimators are well-defined and admit tractable risk bounds. Assumptions 1 and 2 are maintained throughout our analysis and guarantee identifiability and consistency of the least-squares estimator. Assumptions 3 and 4 provide control over the tails of the sample covariance spectrum, which we use to further simplify our risk bounds.
Assumption 1 (Invertible Covariance Matrix).
For each group $g\in\{r,b\}$, the population covariance matrix $\Sigma_{g}$ is invertible.
This assumption ensures that, for each group $g\in\{r,b\}$, $\beta_{g}$ is the unique minimizer of $\mathcal{R}_{g}(\cdot)$.
Assumption 2 (Invertible Sample Covariance Matrix).
For each group $g\in\{r,b\}$, the empirical covariance matrix
$\widehat{\Sigma}_{g}\coloneqq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}^{(g)}X_{i}^{(g)\top}$
is invertible almost surely.
This assumption requires that $n_{g}\geq d$. Given this condition, and as highlighted by mourtada2022exact, 2 is equivalent to either of the following two properties: (i) the distribution $P(X\mid G=g)$ does not assign positive probability to any linear hyperplane; and (ii) the ordinary least squares (OLS) estimator for each group, defined as
$\widehat{\beta}_{g}\coloneqq\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{d}}\widehat{\mathcal{R}}_{g}(\beta)=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{d}}\sum_{i=1}^{n_{g}}(X_{i}^{(g)\top}\beta-Y_{i}^{(g)})^{2},$ (6)
is uniquely defined almost surely, and is given by $\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}.$
Assumption 3 (Small-Ball Condition).
We say that the covariate of group $g$ satisfies the small-ball condition with parameters $C_{g}\geq 1$ and $\alpha_{g}\in(0,1]$ if, for all nonzero $\theta\in\mathbb{R}^{d}$ and all $t>0$,
$\mathbb{P}_{g}\left(|\theta^{\top}X|\leq t\left\lVert\theta\right\rVert_{\Sigma_{g}}\right)\leq\left(C_{g}t\right)^{\alpha_{g}}.$
This condition, adopted from prior work (e.g., koltchinskii2015bounding, mourtada2022exact), provides lower-tail control for the sample covariance spectrum, ensuring that the empirical covariance does not degenerate in directions of low variance of the covariates. This assumption holds for multivariate Gaussian distributions with $\alpha_{g}=1$, since in that case $\theta^{\top}X$ is Gaussian with variance $\|\theta\|_{\Sigma_{g}}^{2}$. Furthermore, rudelson2015small show that this condition holds with $\alpha_{g}=1$ for covariates with independent coordinates and bounded density.
Assumption 4 (Subgaussian Covariates).
We say that the covariate of group $g$ satisfies the subgaussian assumption if there exists a constant $K_{g}\geq 1$ such that, for any $u\in\mathbb{R}^{d}$,
$\mathbb{P}_{g}\left(\left|X^{\top}u\right|\geq t\left\lVert u\right\rVert_{\Sigma_{g}}\right)\leq 2\exp\left(\frac{-t^{2}}{K_{g}^{2}}\right).$
This condition ensures that the covariates have light tails. It is satisfied, for instance, by multivariate Gaussian distributions or bounded distributions.
Remark 1 (Role of Concentration and Anticoncentration).
Assumptions 3 and 4 are complementary and provide control over the spectrum of the sample covariance matrix. The subgaussian condition bounds the upper tail, limiting the growth of the largest eigenvalue and thereby the variance of our estimators. The small-ball condition controls the lower tail, preventing the collapse of the smallest eigenvalue such that inverse covariance terms are bounded. As discussed above, the two assumptions hold simultaneously for a broad class of distributions, including multivariate Gaussian distributions and covariates with independent subgaussian coordinates and bounded density.
3.2 Empirical Estimators
Let $\beta_{\lambda}$ denote the optimal predictor corresponding to the weighted loss $\mathcal{R}_{\lambda}(\beta)$:
$\beta_{\lambda}\coloneqq\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{d}}\mathcal{R}_{\lambda}(\beta).$ (7)
Under 1, it is straightforward to verify that $\beta_{\lambda}$ is given by $\beta_{\lambda}=\Sigma_{\lambda}^{-1}\nu_{\lambda}$, where
$\Sigma_{\lambda}\coloneqq\lambda\Sigma_{r}+(1-\lambda)\Sigma_{b}\quad\text{and}\quad\nu_{\lambda}:=\lambda\nu_{r}+(1-\lambda)\nu_{b}.$ (8)
The next proposition allows us to relate the excess risk of any estimator $\beta$ to that of $\beta_{\lambda}$, as well as to its estimation error on both groups. For details, see Section A.2.2.
Proposition 1.
For a given $\beta\in\mathbb{R}^{d}$, $\lambda\in[0,1]$, and $g\in\mathcal{G}$, the following identities hold:
$\mathcal{R}_{\lambda}(\beta)-\mathcal{R}_{\lambda}(\beta_{\lambda})$ $=\lambda\>\|\beta-\beta_{\lambda}\|_{\Sigma_{r}}^{2}+(1-\lambda)\|\beta-\beta_{\lambda}\|_{\Sigma_{b}}^{2},$ (9a)
$\mathcal{R}_{g}(\beta)-\mathcal{R}_{g}(\beta_{\lambda})$ $=\|\beta-\beta_{\lambda}\|_{\Sigma_{g}}^{2}+2(\beta-\beta_{\lambda})^{\top}\Sigma_{g}(\beta_{\lambda}-\beta_{g}).$ (9b)
Recall that our goal is to minimize the worst case expected excess risk, i.e.,
$\inf_{\beta}\sup_{P\in\mathcal{P}}\mathbb{E}_{\mathcal{S}_{r},\mathcal{S}_{b}}\left[\mathcal{R}_{\lambda}(\beta)-\mathcal{R}_{\lambda}(\beta_{\lambda})\right]=\inf_{\beta}\sup_{P\in\mathcal{P}}\mathbb{E}_{\mathcal{S}_{r},\mathcal{S}_{b}}\left[\lambda\|\beta-\beta_{\lambda}\|_{\Sigma_{r}}^{2}+(1-\lambda)\|\beta-\beta_{\lambda}\|_{\Sigma_{b}}^{2}\right],$ (10)
where $\beta$ is a function of the datasets $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. [^2] Moving forward, we omit the dependence of the expectation on the datasets when it is clear from the context. For the cases $\lambda=0$ and $\lambda=1$, which reduce to the single-group setting (say, group $g$), mourtada2022exact shows that the OLS estimator $\widehat{\beta}_{g}$, defined in (6), is the optimal worst-case estimator with respect to the minimax criterion in (10), when $\mathcal{P}$ is taken to be the linear model class $\mathcal{P}_{\text{linear}}(\sigma^{2})$. Interestingly, he also shows that knowledge of the covariate distribution $P_{X}^{g}$ does not affect this optimality: the OLS estimator remains minimax optimal even when the supremum is taken over the more restricted class $\mathcal{P}_{\text{linear}}(P_{X}^{g},\sigma^{2})$.
This naturally leads us to consider the empirical version of $\mathcal{R}_{\lambda}(\cdot)$, denoted by $\widehat{\mathcal{R}}_{\lambda}(\beta)$, which is given by
$\widehat{\mathcal{R}}_{\lambda}(\beta)\coloneqq\lambda\widehat{\mathcal{R}}_{r}(\beta)+(1-\lambda)\widehat{\mathcal{R}}_{b}(\beta).$ (11)
Under 2, and via first-order optimality conditions, it is straightforward to see the minimizer of (11) is given by
$\widehat{\beta}_{\lambda}:=\widehat{\Sigma}_{\lambda}^{-1}\widehat{\nu}_{\lambda}\quad\text{with}\quad\widehat{\Sigma}_{\lambda}=\lambda\widehat{\Sigma}_{r}+(1-\lambda)\widehat{\Sigma}_{b}\,\,\text{ and }\,\,\widehat{\nu}_{\lambda}:=\lambda\widehat{\nu}_{r}+(1-\lambda)\widehat{\nu}_{b}.$ (12)
As we will see next, this estimator is nearly minimax optimal when the distribution class $\mathcal{P}$ in (10) is taken to be the linear model class $\mathcal{P}_{\text{linear}}(\sigma^{2})$. However, unlike in the one-group case, under additional knowledge of the covariate distribution—or even just the covariance matrices $\Sigma_{r}$ and $\Sigma_{b}$—a different estimator becomes minimax optimal. This marks a sharp departure from the classical linear regression setting, in which the OLS estimator is optimal regardless of knowledge of the covariate distribution.
4 Estimation with Known Covariances
We first consider the setting in which the population covariance matrices $\Sigma_{r}$ and $\Sigma_{b}$ are known. Given samples from both groups, we propose the following estimator:
$\tilde{\beta}_{\lambda}=(\Sigma_{\lambda})^{-1}(\lambda\Sigma_{r}\widehat{\beta}_{r}+(1-\lambda)\Sigma_{b}\widehat{\beta}_{b}),$ (13)
with $\Sigma_{\lambda}=\lambda\Sigma_{r}+(1-\lambda)\Sigma_{b}$, and $\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}$, as defined in (8) and (6), respectively. This estimator maintains the structure of the optimal predictor $\beta_{\lambda}$ but replaces the true cross-moment $\nu_{g}=\Sigma_{g}\beta_{g}$ with the empirical quantity $\Sigma_{g}\widehat{\beta}_{g}$ for each group $g$.
In what follows, we first establish that this estimator is minimax optimal when the covariance matrices (or even the full distribution of the covariates) are known, and then characterize its estimation error for each group as a function of the distribution’s parameters. Finally, we discuss the implications of this result for algorithm design and optimal sampling strategies.
4.1 The Optimal Estimator
As Proposition 1 shows, the excess risk of an estimator can be decomposed into the sum of its distances from the optimal predictor $\beta_{\lambda}$ under the Mahalanobis norms induced by $\Sigma_{r}$ and $\Sigma_{b}$. The following result shows that, for each group $g$, the worst-case distance $\|\beta-\beta_{\lambda}\|_{\Sigma_{g}}^{2}$ is minimized by $\tilde{\beta}_{\lambda}$, thereby implying the minimax optimality of this estimator.
Theorem 1.
Suppose Assumptions 1 and 2 hold. Then, for any group $g\in\{b,r\}$, we have
$\inf_{\beta}\sup_{P\in\mathcal{P}_{\text{linear}}(P_{X}^{r},P_{X}^{b},\sigma^{2})}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ (14)
$\quad\quad=\lambda^{2}\frac{\sigma^{2}}{n_{r}}\mathbb{E}\left[\operatorname{Tr}\left(\Sigma_{g}\Sigma_{\lambda}^{-1}\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\Sigma_{r}\Sigma_{\lambda}^{-1}\right)\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\mathbb{E}\left[\operatorname{Tr}\left(\Sigma_{g}\Sigma_{\lambda}^{-1}\Sigma_{b}\widehat{\Sigma}_{b}^{-1}\Sigma_{b}\Sigma_{\lambda}^{-1}\right)\right],$
and the infimum is achieved by setting $\beta$ equal to $\tilde{\beta}_{\lambda}$, given by (13).
The proof is provided in Section A.3.1. The proof involves two steps: first, characterizing the error rate of the estimator $\tilde{\beta}_{\lambda}$ for any distribution in $\mathcal{P}_{\text{linear}}(P_{X}^{r},P_{X}^{b},\sigma^{2})$; and second, showing that no other estimator can achieve a better rate in the worst-case sense. For the second part, we note that the worst-case error can be lower bounded by the expected error over the family $\mathcal{P}_{\text{linear}}(P_{X}^{r},P_{X}^{b},\sigma^{2})$ for any chosen prior distribution. This expected error is then minimized by the Bayes estimator. Thus, it remains to choose a prior that yields the best lower bound. We do so by selecting a specific Gaussian prior over the group predictors $\beta_{r}$ and $\beta_{b}$, letting its variance tend to infinity, and applying the monotone convergence theorem.
Remark 2.
Notice that, while we take the supremum over the family $\mathcal{P}_{\text{linear}}(P_{X}^{r},P_{X}^{b},\sigma^{2})$, which assumes full knowledge of the distributions of the covariates, the optimal estimator $\tilde{\beta}_{\lambda}$ only uses the covariance matrices. In other words, any knowledge of the distribution of $X$ conditioned on group identity $g$ beyond the covariance matrix $\Sigma_{g}$ does not lead to the design of a better estimator.
As stated in the discussion before Theorem 1, this result, together with Proposition 1, implies the following corollary on the minimax optimality of $\tilde{\beta}_{\lambda}$ with respect to the excess risk of $\mathcal{R}_{\lambda}(\cdot)$.
Corollary 1.
Suppose Assumptions 1 and 2 hold. Then, $\tilde{\beta}_{\lambda}$ is the minimizer of the worst-case excess risk (10) when $\mathcal{P}$ is set as $\mathcal{P}_{\text{linear}}(P_{X}^{r},P_{X}^{b},\sigma^{2})$.
4.2 Bounding the excess risk
In this subsection, to illustrate the implications of Theorem 1, we further simplify the error bound under additional assumptions on the distribution. First, note that when the covariance matrices are known, it is without loss of generality to assume spherical covariances, since we can apply the transformation
$\tilde{X}^{(g)}\coloneqq\Sigma_{g}^{-1/2}X^{(g)},\quad\tilde{\beta}_{g}\coloneqq\Sigma_{g}^{-1/2}\beta_{g},$
which yields a spherical covariance matrix for the covariate vectors. We thus assume that each group has a spherical covariance structure, i.e., $\Sigma_{g}=\rho_{g}^{2}I_{d}$ for known $\rho_{g}>0$.
Corollary 2.
Suppose that for each group $g\in\mathcal{G}$, the covariance matrix satisfies $\Sigma_{g}=\rho_{g}^{2}I_{d}$ for known $\rho_{g}>0$. Furthermore, suppose 2 holds and the small-ball condition (3) also holds with constants $(C_{g},\alpha_{g})$. Also suppose that $n_{g}\geq 6d/\alpha_{g}$ and that $d\geq 2$. Then we have
$\mathbb{E}\left[\left\lVert\tilde{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\leq\frac{2\sigma^{2}d\rho_{g}^{2}}{(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2})^{2}}\left(\lambda^{2}C_{r}^{\prime}\frac{\rho_{r}^{2}}{n_{r}}+(1-\lambda)^{2}C_{b}^{\prime}\frac{\rho_{b}^{2}}{n_{b}}\right),$ (15)
where $C_{g}^{\prime}=3C_{g}^{4}\exp(1+9/\alpha_{g}).$ Furthermore, when the covariates satisfy 4 with parameter $K_{g}$ and $n_{g}\geq\min\left\{6\alpha_{g}^{-1}d,12\alpha_{g}^{-1}\log(12\alpha_{g}^{-1})\right\}$, the bound (15) holds with $\left(1+\frac{8C_{g}^{\prime}\zeta\rho_{g}^{8}K_{g}^{4}d}{n_{g}}\right)$ in place of $C_{g}^{\prime}$ for $g\in\{r,b\}$, where $\zeta>0$ is a universal constant.
See Section A.3.2 for the proof.
Remark 3.
The constant $C’_{g}$ arises from the bound on the minimum eigenvalue of the sample covariance, which is used here to bound the trace of $\widehat{\Sigma}_{r}^{-1}$ and $\widehat{\Sigma}_{b}^{-1}$, as given in mourtada2022exact. However, as highlighted in mourtada2022exact—based on results from wu2012optimal and edelman1988eigenvalues—this constant can be significantly reduced for Gaussian distributions. In particular, if the covariate distributions are Gaussian, $C’_{g}$ can be replaced by a constant whose limit, when ${d}/{n_{g}}\to h\in(0,1)$, is upper bounded by
$\left(\frac{1}{h}\right)^{3h}\left(\frac{\sqrt{e}}{1-h}\right)^{3(1-h)},$
which, in turn, can be shown to be upper bounded by $(1+\sqrt{e})^{3}$.
The following corollary combines the differences in matrix norms to explicitly characterize the excess risk bound for the optimal estimator $\tilde{\beta}_{\lambda}$.
Corollary 3.
Under the premise of Corollary 2, we have
$\mathbb{E}\left[\mathcal{R}_{\lambda}(\tilde{\beta}_{\lambda})-\mathcal{R}_{\lambda}(\beta_{\lambda})\right]\leq\frac{2\sigma^{2}d}{\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}}\left(\lambda^{2}C_{r}^{\prime}\frac{\rho_{r}^{2}}{n_{r}}+(1-\lambda)^{2}C_{b}^{\prime}\frac{\rho_{b}^{2}}{n_{b}}\right).$ (16)
We conclude this section with a few remarks on the insights we draw from these results.
Optimal allocation of the sampling budget:
In many applications, we have a limited sampling budget, so it is natural to ask: how should we allocate our sampling budget across the two groups? In other words, suppose $n_{r}+n_{b}$ is fixed; what choice of $n_{r}$ and $n_{b}$ minimizes the excess risk? Corollary 2, along with a simple Cauchy–Schwarz inequality, suggests setting
$\frac{n_{r}}{n_{b}}=\frac{\lambda\rho_{r}}{(1-\lambda)\rho_{b}}.$ (17)
The dependence on $\lambda$ is intuitive: the more weight you put on each group in your objective function, the more samples you allocate to that group. To interpret the dependence on $\rho_{r}$ and $\rho_{b}$, suppose the covariates are zero-mean with per-coordinate variance $\rho_{g}^{2}$. Then, when we take an average over $n_{g}$ samples, the per-coordinate variance decreases to $\rho_{g}^{2}/n_{g}$. Thus, the proposed sampling weight (17) balances the variances of the two groups by equalizing the variance-weighted sample sizes.
Per-group estimation error—finite-sample effects can shift fairness–accuracy trade-offs:
Note that, for each group $g\in\{r,b\}$, the estimator $\widehat{\beta}_{g}$ is unbiased for $\beta_{g}$, since
$\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}=\beta_{g}+\widehat{\Sigma}_{g}^{-1}\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}^{(g)}\varepsilon_{i}^{(g)}.$ (18)
Consequently, the optimal estimator $\tilde{\beta}_{\lambda}$ is also an unbiased estimator of $\beta_{\lambda}$. As a result, and by Proposition 1, the estimation error of this estimator on group $g$ is given by
$\mathbb{E}\left[\mathcal{R}_{g}(\tilde{\beta}_{\lambda})-\mathcal{R}_{g}(\beta_{\lambda})\right]=\mathbb{E}\left[\|\tilde{\beta}_{\lambda}-\beta_{\lambda}\|_{\Sigma_{g}}^{2}\right].$ (19)
In other words, the bounds in Theorem 1 and Corollary 2 quantify the per-group estimation error caused by using the empirical estimator instead of the true parameter. Interestingly, when we compare these errors across the two groups, they differ (only) in the $\rho_{g}^{2}$ term. That is, while the empirical estimator affects the risks of the two groups differently, this difference arises not from unequal sample sizes per group but from differences in their covariance matrices. Intuitively, this happens because when one group’s sample size is low, it impacts the accuracy of the estimator $\tilde{\beta}_{\lambda}$, which, in turn, affects the error for both groups. That said, if a group’s covariate has higher variance, this error is amplified more, and thus that group’s risk increases more when using the empirical estimator. An important implication of this discrepancy is that finite-sample estimation can distort the fairness-accuracy trade-off encoded by $\lambda$, shifting it away from the planner’s intended balance.
Expectation vs. one-instance realization:
It is worth noting that the results above are computed in expectation over the random draw of datasets $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$, and are not for a particular realization. For the excess risk with respect to the weighted loss, i.e., $\mathcal{R}_{\lambda}(\tilde{\beta}_{\lambda})-\mathcal{R}_{\lambda}(\beta_{\lambda})$, however, each realization is also positive. This means that the empirical estimator always increases the loss function $\mathcal{R}_{\lambda}(\cdot)$, simply because $\beta_{\lambda}$ is its minimizer. By contrast, this monotonicity does not hold for per-group risks: one group’s estimation error $\mathcal{R}_{g}(\cdot)$ might decrease under the empirical estimator relative to $\beta_{\lambda}$, even though the other group’s risk increases enough to make the weighted objective worse. The following result formalizes this distinction.
Proposition 2.
Suppose Assumptions 13 hold, the covariance matrix $\Sigma_{g}$ is equal to $\rho_{g}^{2}I_{d}$ for some $\rho_{g}>0$, $\lambda\in(0,1)$, and $n_{g}\geq 48/\alpha_{g}$. Then, for any $\delta>0$, there exists $\bar{n}$ such that, for any $n_{r},n_{b}\geq\bar{n}$, with probability $1/2-\delta$ over the draw of samples, we have
$\mathcal{R}_{g}(\tilde{\beta}_{\lambda})<\mathcal{R}_{g}(\beta_{\lambda}).$
The proof is given in Section A.3.3. This result follows from the decomposition established in Proposition 1:
$\mathcal{R}_{g}(\beta)-\mathcal{R}_{g}(\beta_{\lambda})=\|\beta-\beta_{\lambda}\|_{\Sigma_{g}}^{2}+2~(\beta-\beta_{\lambda})^{\top}\Sigma_{g}(\beta_{\lambda}-\beta_{g}).$ (20)
The first term is nonnegative and, with high probability, bounded by $C_{1}/n$ for some constant $C_{1}$. The second term, when multiplied by $\sqrt{n}$, converges to a mean-zero Gaussian with constant variance (up to an error of order $\mathcal{O}(1/\sqrt{n})$). Therefore, for large enough $n$, the leading term is a mean-zero term of order $\mathcal{O}(1/\sqrt{n})$.
Thus, although the expected per-group estimation error increases when using an empirical estimator, for large enough $n$ it increases with a probability close to $\frac{1}{2}$. On the other hand, since the excess risk—defined as the $\lambda$-weighted sum of the two groups’ estimation errors—is always nonnegative under an empirical estimator, in any given realization at most one group can benefit from estimation error; any such gain must be offset by a larger loss for the other group.
5 Estimation with Unknown Covariance
We next consider the setting in which the distributions of the covariates are unknown and must be estimated from data. Here, our proposed estimator is the OLS estimator (12), which, as we recall, is given by:
$\widehat{\beta}_{\lambda}=\widehat{\Sigma}_{\lambda}^{-1}\widehat{\nu}_{\lambda}=(\lambda\widehat{\Sigma}_{r}+(1-\lambda)\widehat{\Sigma}_{b})^{-1}(\lambda\widehat{\nu}_{r}+(1-\lambda)\widehat{\nu}_{b}).$ (21)
In this section, we first present a result that characterizes the excess risk of this estimator and decomposes it into bias and variance terms. We then provide upper and lower bounds under additional distributional assumptions (namely, subgaussian and small-ball conditions), thereby establishing the near-optimality of the estimator under these assumptions.
5.1 The Bias-Variance Decomposition
Recall from Proposition 1 that the excess risk of $\widehat{\beta}_{\lambda}$ can be expressed as the weighted sum of the two distances $\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{r}}^{2}$ and $\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{b}}^{2}$. Thus, as in the previous section, we begin by characterizing these two distances.
Proposition 3.
Suppose Assumptions 1 and 2 hold. Then, for any group $g\in\{r,b\}$, we have:
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]=\mathcal{V}_{g}(\lambda)+\mathcal{B}_{g}(\lambda),$ (22)
where the variance term is given by
$\mathcal{V}_{g}(\lambda)\coloneq\lambda^{2}\frac{\sigma^{2}}{n_{r}}\mathbb{E}\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-1}\Sigma_{g}\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{r}\right)+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\mathbb{E}\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-1}\Sigma_{g}\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{b}\right),$ (23)
and the bias term is given by
$\mathcal{B}_{g}(\lambda)$ $\coloneq\mathbb{E}\left[\left\lVert\Sigma_{g}^{1/2}\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})\right)\right\rVert^{2}\right]$ (24)
$=\mathbb{E}\left[\left\lVert\Sigma_{g}^{1/2}\left[\left(I_{d}+\frac{1-\lambda}{\lambda}\widehat{\Sigma}_{r}^{-1}\widehat{\Sigma}_{b}\right)^{-1}-\left(I_{d}+\frac{1-\lambda}{\lambda}\Sigma_{r}^{-1}\Sigma_{b}\right)^{-1}\right]\right\rVert^{2}\right]\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$ (25)
See Section A.4.1 for the proof. The variance term closely resembles the error rate in the known-covariance case: it captures the irreducible sampling error within each group, as it is scaled by the squared fairness weights and the inverse sample sizes of each group. The bias term, however, represents a key departure from the known-covariance setting. It arises from heterogeneity in the true underlying coefficients $\beta_{r}$ and $\beta_{b}$, it depends explicitly on $\beta_{r}-\beta_{b}$, and it vanishes when $\beta_{r}=\beta_{b}$.
To interpret this bias term more precisely, note that, even if $\beta_{r}$ and $\beta_{b}$ were known exactly, the target parameter $\beta_{\lambda}$ cannot be recovered without also knowing the covariance matrices $\Sigma_{r}$, $\Sigma_{b}$. In fact, $\beta_{\lambda}$ is a covariance-weighted average of $\beta_{r}$ and $\beta_{b}$; it coincides with both when the group-specific parameters are equal, and the discrepancy increases as the parameters diverge. The bias term quantifies the additional error introduced by replacing the true covariance structure with empirical estimates, which explains why it vanishes in the known-covariance case.
5.2 Upper and lower bounds for the bias and variance terms
In this section, we investigate the optimality of the estimator $\widehat{\beta}_{\lambda}$ under additional assumptions. For our upper bounds, we assume the distribution of covariates is subgaussian (4) and also satisfies the small-ball condition (3). We further assume that, for any group $g\in\{r,b\}$,
$\frac{1}{2}\rho_{g}^{2}I_{d}\preceq\Sigma_{g}\preceq\frac{3}{2}\rho_{g}^{2}I_{d}.$ (26)
The constants $1/2$ and $3/2$ are chosen for simplicity; as we show in the appendix, our upper bounds extend to more general conditions on the eigenvalues of the covariance matrix $\Sigma_{g}$, but we adopt these specific values in the main text for clarity.
For the lower bounds, we must specify the class of distributions $\mathcal{P}$ in (10) over which the worst-case excess risk is taken. To match our upper bounds, we impose the same assumptions as above. In addition, we have to assume the boundedness of the group-specific predictors $\beta_{r}$ and $\beta_{b}$, since Proposition 3 (and later our lower bound) shows that the bias term grows with $\|\beta_{r}-\beta_{b}\|$ and would diverge if the groups’ predictors were unbounded.
Accordingly, we consider the following subclass of $\mathcal{P}_{\text{linear}}(\sigma^{2})$, consisting of Gaussian covariates satisfying the assumptions above:
$\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)$= $\left\{P\in\mathcal{P}_{\text{linear}}(\sigma^{2})~\big|~\forall g\in\mathcal{G}:|\beta_{g}|\leq B\text{ and }P(X|G=g)\text{ is}\right$. (27)
 $\left.\text{ Gaussian with }~\frac{1}{2}\rho_{g}^{2}I_{d}\preceq\Sigma_{g}\preceq\frac{3}{2}\rho_{g}^{2}\right\}.$
We treat the variance and bias terms separately. For the upper bound, we bound the quantities $\mathcal{V}_{g}(\lambda)$ and $\mathcal{B}_{g}(\lambda)$ from Proposition 3. For the lower bound, we decompose the worst-case error into bias and variance components by considering two complementary scenarios. Specifically, we lower bound
$\sup_{P\in\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ (28)
by the maximum of two restricted subproblems: (1) the case where the covariance matrices $\Sigma_{r}$ and $\Sigma_{b}$ are known but the group predictors $\beta_{r}$ and $\beta_{b}$ are unknown (corresponding to the variance term), and (2) the case where the group predictors are known but the covariance matrices are unknown (corresponding to the bias term). Formally, consider
$\max\left\{\sup_{\begin{subarray}{c}P\in\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)\\ \Sigma_{r},\Sigma_{b}\text{ are known.}\end{subarray}}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right],\sup_{\begin{subarray}{c}P\in\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)\\ \beta_{r},\beta_{b}\text{ are known.}\end{subarray}}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\right\},$ (29)
which can be further lower bounded by the average of these two terms. Our next two results establish upper and lower bounds for the bias and variance terms, respectively.
Theorem 2.
Suppose Assumptions 14 hold, the covariance matrices satisfy (26), and, for $g\in\mathcal{G}$, and $n_{g}\geq\max\{48/\alpha_{g},K_{g}^{4}d\}$. Then, we have:
$\mathcal{V}_{g}(\lambda)\lesssim\frac{\rho_{g}^{2}\sigma^{2}d}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}}\left(\frac{\lambda^{2}\rho_{r}^{2}}{n_{r}}+\frac{(1-\lambda)^{2}\rho_{b}^{2}}{n_{b}}\right),$ (30)
where $\mathcal{V}_{g}(\lambda)$ is the variance term, as defined in Proposition 3, and $C^{\prime}_{g}$ is given in Corollary 2 and Remark 3. Moreover, assuming $n_{g}\geq\sigma^{2}d/(B\rho_{g}^{2})$ for $g\in\mathcal{G}$, we have
$\inf_{\beta}\sup_{\begin{subarray}{c}P\in\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)\\ \Sigma_{r},\Sigma_{b}\text{ are known.}\end{subarray}}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\gtrsim\frac{\rho_{g}^{2}\sigma^{2}d}{(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2})^{2}}\left(\frac{\lambda^{2}\rho_{r}^{2}}{n_{r}}+\frac{(1-\lambda)^{2}\rho_{b}^{2}}{n_{b}}\right),$ (31)
where the infimum is taken over any estimator $\beta$ as a function of the datasets $(\mathcal{S}_{r},\mathcal{S}_{b})$.
See Section A.4.2 for the proof. The upper and lower bounds together show that, up to constant factors, the variance term indeed captures the error arising from not knowing the true predictors $\beta_{r}$ and $\beta_{b}$, even in the known-covariance case. Moreover, the estimator $\widehat{\beta}_{\lambda}$ is minimax-optimal, again up to constant factors.
It is worth noting that, while the lower bound is for the known-covariance case, the result in Theorem 1 from the previous section is not applicable here. This is because we now operate under the additional assumption that the group predictors $\beta_{r}$ and $\beta_{b}$ are bounded, whereas in that earlier result, the lower bound was constructed by defining a prior over a family of distributions with unbounded parameters.
The proof of the lower bound applies Assouad’s method. Specifically, we construct a finite subclass of $\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)$ of size $2^{2d}$, such that the concatenated vector $\beta_{r}^{\top},\beta_{b}^{\top}]^{\top}$ lies on the vertices of a $2d$-dimensional hypercube with a carefully chosen side length. The minimax risk is then bounded below by the worst-case error over this subclass. Assouad’s lemma, as recalled in the appendix, reduces this problem to one of testing multiple hypotheses, from which the stated bound follows.
We next derive upper and lower bounds for the bias term.
Theorem 3.
Suppose Assumptions 14 hold, the covariance matrices satisfy (26), and, for $g\in\mathcal{G}$, $n_{g}\geq 48/\alpha_{g}$. Then, we have:
$\mathcal{B}_{g}(\lambda)\lesssim\frac{\lambda^{2}(1-\lambda)^{2}~\rho_{g}^{2}~\rho_{r}^{4}~\rho_{b}^{4}~d}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}~\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{2}}~\left(\frac{K_{r}^{4}}{n_{r}}+\frac{K_{b}^{4}}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2},$ (32)
where $\mathcal{B}_{g}(\lambda)$ is the bias term, as defined in Proposition 3, and $C^{\prime}_{g}$ is given in Corollary 2 and Remark 3. Moreover, if $n_{g}\geq 16d^{2}$ for $g\in\mathcal{G}$, we have
$\inf_{\beta}\sup_{\begin{subarray}{c}P\in\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)\\ \beta_{r},\beta_{b}\text{ are known.}\end{subarray}}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\gtrsim\frac{\lambda^{2}(1-\lambda)^{2}~\rho_{g}^{2}~\rho_{r}^{4}~\rho_{b}^{4}~d}{\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{4}}~\left(\frac{1}{n_{r}}+\frac{1}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2},$ (33)
where the infimum is taken over any estimator $\beta$ as a function of the datasets $(\mathcal{S}_{r},\mathcal{S}_{b})$.
See Section A.4.3 for the proof. Recall that the constants $K_{r}$ and $K_{b}$, as defined in 4, are invariant to scaling and hence are independent of $\rho_{r}$ and $\rho_{b}$. As a result, the upper and lower bounds above match up to constant factors, which again shows that the bias term truly captures the error arising from not knowing the covariance matrices, even when the group predictors $\beta_{r}$ and $\beta_{b}$ are known. Moreover, Theorem 3, together with the result of Theorem 2, highlights that the OLS estimator $\widehat{\beta}_{\lambda}$ achieves the minimax excess risk, as given in (10) with $\mathcal{P}=\mathcal{P}_{\text{Gauss}}(\sigma^{2},\rho_{r}^{2},\rho_{b}^{2},B)$, up to constant factors.
The proof of the bias lower bound also uses Assouad’s method, but over a delicately-constructed family of perturbed covariance matrices. To this end, define $\{u_{i}u_{i}^{\top}\}_{i=1}^{d}$, each of which is rank-one and positive semi-definite, where $u_{i}$ is defined as
$u_{i}$ $=\begin{cases}\frac{e_{i}+v}{\left\|e_{i}+v\right\|_{2}},&\text{if }e_{i}^{\top}v\geq 0,\\ \frac{e_{i}-v}{\left\|e_{i}-v\right\|_{2}},&\text{if }e_{i}^{\top}v\leq 0,\end{cases}\quad\text{ with }\quad v=\frac{\beta_{r}-\beta_{b}}{\|\beta_{r}-\beta_{b}\|}.$ (34)
For each group $g$, we perturb the baseline covariance matrix $\rho_{g}^{2}I_{d}$ in $2^{d}$ different ways by either adding or subtracting $h_{g}u_{i}u_{i}^{\top}$ for each $i$ and for some scalar $h_{g}$. This yields $2^{d}$ distinct covariance matrices, which form the family of distributions over which we take the worst-case error and apply Assouad’s method.
The choice of the perturbations $\{u_{i}\}$ is key here. Since the bias term has the form $M(\beta_{r}-\beta_{b})$ for some matrix $M$, its magnitude is maximized when $\beta_{r}-\beta_{b}$ is closely aligned with an eigenvector of $M$ corresponding to its largest eigenvalue. Hence we choose the perturbations so that this alignment holds, i.e., the construction of $u_{i}$ ensures that it has a positive and constant inner product with both $\beta_{r}-\beta_{b}$ and the canonical basis vector $e_{i}$. This alignment guarantees that we perturb the covariance matrix along all coordinate directions while maintaining proximity to $\beta_{r}-\beta_{b}$.
The next corollary combines the bounds on the variance and bias terms across both groups, in accordance with Proposition 1, to present the excess risk bound for the estimator $\widehat{\beta}_{\lambda}$.
Corollary 4.
Suppose Assumptions 14 hold, the covariance matrices satisfy (26), and, for $g\in\mathcal{G}$, and $n_{g}\geq\max\{48/\alpha_{g},K_{g}^{4}d\}$. Then, we have:
$\mathbb{E}\left[\mathcal{R}_{\lambda}(\widehat{\beta}_{\lambda})-\mathcal{R}_{\lambda}(\beta_{\lambda})\right]\lesssim\sigma^{2}d~\frac{\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}}\left(\frac{\lambda^{2}\rho_{r}^{2}}{n_{r}}+\frac{(1-\lambda)^{2}\rho_{b}^{2}}{n_{b}}\right)$
$+\frac{\lambda^{2}(1-\lambda)^{2}~\rho_{r}^{4}~\rho_{b}^{4}~d}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}~\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)}~\left(\frac{K_{r}^{4}}{n_{r}}+\frac{K_{b}^{4}}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$
Optimal allocation of the sampling budget:
We can pose the same question as in the previous section: given a fixed budget on $n_{r}+n_{b}$, what is the optimal way to choose these two parameters to minimize the excess risk? Theorem 2 suggests that, for minimizing the variance term, the answer is similar to the case of known covariances. In fact, similar to (17), we should choose ${n_{r}}/{n_{b}}=(\lambda\rho_{r})/((1-\lambda)\rho_{b})$ to minimize the variance term. In contrast, Theorem 3 shows that minimizing the bias term requires a balanced design, namely, $n_{r}=n_{b}$ (under the assumption of $K_{r}=K_{b}$).
Thus, the optimal sampling allocation is more nuanced and depends on whether the variance term or the bias term dominates the error, which in turn depends on the heterogeneity between the groups, i.e., $\|\beta_{r}-\beta_{b}\|$. In fact, the more the two groups differ, the larger the bias term becomes, and the more we would prefer $n_{r}$ and $n_{b}$ to be closer.
5.3 Per-Group estimation errors
We next ask: what is the impact of using the OLS estimator $\widehat{\beta}_{\lambda}$ instead of the true estimator $\beta_{\lambda}$ on each group’s risk? Recall from Proposition 1 that the excess risk for group $g\in\mathcal{G}$ can be written as
$\mathcal{R}_{g}(\widehat{\beta}_{\lambda})-\mathcal{R}_{g}(\beta_{\lambda})=\|\widehat{\beta}_{\lambda}-\beta_{\lambda}\|_{\Sigma_{g}}^{2}+2(\widehat{\beta}_{\lambda}-\beta_{\lambda})^{\top}\Sigma_{g}(\beta_{\lambda}-\beta_{g}).$ (35)
In the known-covariance case (see Section 4), the expectation of the second term on the right-hand side is zero, so our bounds on $\|\widehat{\beta}_{\lambda}-\beta_{\lambda}\|_{\Sigma_{g}}^{2}$ translate directly to the per-group (expected) estimation error. However, that is not the case here, as $\widehat{\beta}_{\lambda}$ is not an unbiased estimate of $\beta_{\lambda}$. Therefore, we must also bound the second term.
So far, our results have provided bounds on the expectation of the first term on the right-hand side of (35), showing that its dependence on the sample size takes the form $\mathcal{O}(1/n_{r}+1/n_{b})$. Our next result characterizes the second term in the expectation of the right-hand side of (35) and shows that it also admits a similar bound.
Proposition 4.
Suppose Assumptions 14 hold, the covariance matrices satisfy (26), and, for $g\in\mathcal{G}$, $n_{g}\geq 48/\alpha_{g}$. Then, we have:
$\left|\mathbb{E}\left[(\widehat{\beta}_{\lambda}-\beta_{\lambda})^{\top}\Sigma_{r}(\beta_{\lambda}-\beta_{r})\right]\right|\lesssim$
$\frac{\lambda(1-\lambda)^{2}d\rho_{r}^{4}\rho_{b}^{4}~~\|\beta_{r}-\beta_{b}\|^{2}}{\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{3}\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)}~\left(\frac{K_{r}^{2}}{\sqrt{n_{r}}}+\frac{K_{b}^{2}}{\sqrt{n_{b}}}\right)\left(\lambda\frac{K_{r}^{2}\rho_{r}^{2}}{\sqrt{n_{r}}}+(1-\lambda)\frac{K_{b}^{2}\rho_{b}^{2}}{\sqrt{n_{b}}}\right),$
where $C^{\prime}_{g}$ is given in Corollary 2 and Remark 3.
The proof is provided in Section A.4.4. A similar result can also be stated for group $b$, with the only difference being that $\lambda(1-\lambda)^{2}$ in the bound is replaced by $\lambda^{2}(1-\lambda)$.
Proposition 4, together with the bounds from Theorem 2 and Theorem 3, shows that, much like in the known-covariance setting, there is an inherent asymmetry in how the empirical estimator affects the risks of the two groups. Specifically, if we examine the per-group estimation errors in (35), we see that the first term on the right-hand side differs between group $r$ and group $b$ only through the substitution of $\rho_{r}^{2}$ with $\rho_{b}^{2}$. This mirrors exactly what was observed in the known-covariance setting.
The second term, $(\widehat{\beta}_{\lambda}-\beta_{\lambda})^{\top}\Sigma_{g}(\beta_{\lambda}-\beta_{g})$, captures a new phenomenon arising from the bias term. First, note that the bound in Proposition 4 controls its absolute value, but the term itself may be positive or negative. Moreover, it is straightforward to verify that if we sum this term across the two groups with weights $\lambda$ and $1-\lambda$, the result is zero. Thus, its contribution to one group’s risk is always offset by the other’s, ensuring that the two groups experience it with opposite signs. Even if the absolute-value bound were symmetric, the term would still be a source of disparity between the two groups’ risks.
However, the bound is in fact not symmetric. As stated after the proposition, it changes from $\lambda(1-\lambda)^{2}$ for group $r$ to $\lambda^{2}(1-\lambda)$ for group $b$. This bias effect is larger for group $r$ when $\lambda$ is close to zero and for group $b$ when $\lambda$ is close to one, meaning that, at those endpoints, the term is larger in absolute value for the group that is not prioritized.
These observations highlight that the fairness-accuracy trade-off implied by the choice of $\lambda$ in the population objective is not necessarily the trade-off realized in finite samples. The empirical estimator introduces systematic differences in the per-group risk in a way that shifts the balance away from the intended allocation. At the same time, it is noteworthy that these differences are not driven by sample imbalances. As in the known-covariance case, all bounds remain similar in $n_{r}$ and $n_{b}$ across both groups.
6 Uniform Frontier Bounds
So far, our analysis has provided finite-sample guarantees for a single fixed fairness–accuracy preference $\lambda$. In practice, however, a designer may wish to consider many values of $\lambda$ in order to select a data-driven choice $\widehat{\lambda}$. Because the empirical errors for different $\lambda$ values are computed using the same datasets, these errors are highly correlated. One approach is to use a union bound, but even then the designer can only afford to check finitely many values of $\lambda$, since the error would accumulate and requesting additional $\lambda$ values would become “costly.”
In this section, we show how to circumvent this issue by obtaining a uniform guarantee over the entire empirical frontier. We achieve this by viewing the empirical error of the FA frontier as a stochastic process indexed by $\lambda\in[0,1]$. This result follows from the construction of a Doob martingale to show a bounded-difference property of the process followed by a chaining argument.
Defining the population and empirical error pairs across groups,
$\mathcal{R}_{\lambda}\coloneq\left(\mathcal{R}_{r}(\beta_{\lambda}),\mathcal{R}_{b}(\beta_{\lambda})\right),\qquad\widehat{\mathcal{R}}_{\lambda}\coloneq\left(\widehat{\mathcal{R}}_{r}(\widehat{\beta}_{\lambda}),\widehat{\mathcal{R}}_{b}(\widehat{\beta}_{\lambda})\right),$
we define the frontier error as a stochastic process:
$\Delta_{\lambda}\coloneq\widehat{\mathcal{R}}_{\lambda}-\mathcal{R}_{\lambda}\in\mathbb{R}^{2},\quad\lambda\in[0,1].$
We will assume that the noise in the linear model given by Definition 3 is subgaussian.
Assumption 5 (Subgaussian noise).
For both groups $g\in\{r,b\}$, the noise $\varepsilon$ is conditionally mean-zero and $\sigma_{\varepsilon}$-subgaussian, independent of $X$.
The following theorem establishes a high-probability uniform bound on the deviation of the empirical fairness-accuracy frontier from its population counterpart across all possible fairness-accuracy preferences $\lambda\in[0,1]$ for the unknown-covariance estimator $\widehat{\beta}$ defined in (21).
Theorem 4.
Suppose Assumptions 1-5 hold. Then, for any fixed $\alpha>2$, there exists a constant $C_{u}$, such that, for every $\delta\in(0,1)$, with probability at least $1-\delta-\eta$, we have
$\sup_{\lambda\in[0,1]}\left\lVert\Delta_{\lambda}\right\rVert_{2}\leq C_{u}\log(\frac{1}{\delta})\left(\frac{\log(n_{r})^{5}}{\sqrt{n_{r}}}+\frac{\log(n_{b})^{5}}{\sqrt{n_{b}}}\right),$
with
$\eta=(n_{r}+n_{b})^{-C\alpha}+\left(C_{r}^{\prime}/2\right)^{\alpha_{r}n_{r}/6}+\left(C_{b}^{\prime}/2\right)^{\alpha_{b}n_{b}/6}.$
The constant $C_{u}$ depends on the problem’s parameters, and the exact dependence is provided in the appendix.
Theorem 4 provides control over the empirical error process with high probability uniformly over the entire frontier, that is, for all $\lambda\in[0,1]$. This result demonstrates that a designer implementing the frontier in finite samples, will, with high probability, land within a confidence band around the population frontier. In particular, with probability at least $1-\delta-\eta$, the empirical curve lies within a distance that scales as $1/\sqrt{n_{r}}+1/\sqrt{n_{b}}$ with the sample sizes of both groups (ignoring logarithmic terms).
We sketch the proof of the uniform empirical frontier bound in Theorem 4 here; the full proof appears in Section A.5. First, we show that, with high probability, features and outcomes are bounded and empirical covariances are well-conditioned, ensuring stability of the estimator. Next, we show that $\widehat{\beta}_{\lambda}$ is Lipschitz in $\lambda$ which implies Lipschitzness in the empirical risks. We then analyze single sample perturbations (where we replace one observation with an i.i.d draw) to obtain a bounded-difference property. We use a Doob martingale argument with Hoeffding’s lemma to establish subgaussianity of the increments in $|\lambda-\lambda^{\prime}|$ for the scalarized process. Chaining via Dudley’s integral inequality converts the increment control to a uniform bound over all $\lambda$, and an $\varepsilon$-net argument over all unit vectors brings the argument back to the two-dimensional empirical error process across the frontier. Finally, we de-condition using the high-probability truncation event to yield the result.
Since the bound given by Theorem 4 holds simultaneously for all $\lambda\in[0,1]$, a designer may select $\widehat{\lambda}$ adaptively after inspecting the empirical frontier while enjoying the same high-probability guarantee. Moreover, the result strengthens the guarantees provided under a naïve union bound argument on a finite grid: the bound holds for the entire interval $[0,1]$ simultaneously, including values not explicitly evaluated, so the designer may refine their search at no additional cost.
7 Experiments
In this section, we conduct synthetic experiments to demonstrate how well the estimators $\tilde{\beta}_{\lambda}$ and $\widehat{\beta}_{\lambda}$, for the known and unknown covariance settings, respectively, perform relative to the population quantity across the FA frontier.
We construct a two-group Gaussian linear model in $d$ dimensions with isotropic covariances $\Sigma_{g}=\rho_{g}^{2}I_{d}$. For each group $g\in\{r,b\}$, features $X_{g}\sim\mathcal{N}(0,\Sigma_{g})$, and outcomes $Y_{g}=X_{g}^{\top}\beta_{g}+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$. The pair $(\beta_{r},\beta_{b})$ is constructed via a dense random perturbation scaled so that the heterogeneity $\left\lVert\beta_{b}-\beta_{r}\right\rVert_{2}$ is fixed across iterations. We evaluate (i) the deterministic population frontier from $\beta_{\lambda}$ using the true $\Sigma_{g}$ and $\beta_{g}$, (ii) the empirical frontier in the known $\Sigma_{g}$ case from $\tilde{\beta}_{\lambda}$ given in (13), and (iii) the empirical frontier in the unknown $\Sigma_{g}$ case from $\widehat{\beta}_{\lambda}$ given in (21).
[…]
Figure 5: We sweep $\lambda$ on a uniform grid of 5050 points in $[0,1]$. For each $\lambda$, we run 100 Monte Carlo repetitions; in each iteration, we draw data $\mathcal{S}_{g}=\{(X_{g},Y_{g})\}_{i=1}^{n_{g}}$ from the group-$g$ Gaussian linear model with $\sigma^{2}=1$, build the estimators $\tilde{\beta}_{\lambda},\widehat{\beta}_{\lambda}$, and calculate the corresponding risk pair $(\mathcal{R}_{r},\mathcal{R}_{b})$ for each estimator. We plot the mean empirical frontiers across repetitions and depict the empirical contraction for the choice of $\lambda=0.5$ for two regimes: (a) Equal samples: $n_{r}=n_{b}=30,d=12,\rho_{r}=1,\rho_{b}=2.5$, (b) Equal covariance: $n_{r}=50,n_{b}=30,d=12,\rho_{r}=\rho_{b}=1.75$.
Figure 5 isolates the effect of heterogeneous covariance and sample size. Figure 5(a) illustrates the unequal covariance, equal sample size regime ($\rho_{b}>\rho_{r}$, $n_{r}=n_{b}$). Here, we observe that, the segment connecting the population point $\left(\mathcal{R}_{r}(\beta_{1/2}),\mathcal{R}_{b}(\beta_{1/2})\right)$ to the corresponding empirical points $\left(\mathcal{R}_{r}(\tilde{\beta}_{1/2}),\mathcal{R}_{b}(\tilde{\beta}_{1/2})\right)$ and $\left(\mathcal{R}_{r}(\widehat{\beta}_{1/2}),\mathcal{R}_{b}(\widehat{\beta}_{1/2})\right)$ is not parallel to the $45^{\circ}$ line. Instead, we see that it is tilted towards group $b$, reflecting a larger inflation in $\mathcal{R}_{b}$ than $\mathcal{R}_{r}$, on average. This asymmetry is explained by the results of Corollary 2 and Proposition 4, which show that asymmetry in the covariance across groups drives asymmetric contraction of the empirical frontier estimated in finite samples.
In contrast, in the unequal sample size, equal covariance regime ($n_{r}>n_{b}$, $\rho_{r}=\rho_{b}$) illustrated in Figure 5(b), the same segment is approximately parallel to the $45^{\circ}$ line. In other words, with equal covariance, we see that the inflation in $\mathcal{R}_{b}$ than $\mathcal{R}_{r}$ is roughly symmetric for both the known and unknown covariance estimation regimes. Beyond $\lambda=1/2$, we observe that the same pattern holds across the frontier.
Moreover, for both figures, we observe that the two empirical FA frontiers for known and unknown covariances approach each other for $\lambda$ near zero or one, and coincide at the endpoints $\lambda\in\{0,1\}$. The latter is expected, as the endpoints represent the per-group cases in which the optimal estimator does not depend on the knowledge of the covariance. The former observation, that the difference between the two frontiers shrinks around the endpoints, directly reflects our result in Theorem 3, which shows that the bias term of the unknown-covariance estimator includes a factor of $\lambda^{2}(1-\lambda)^{2}$, and therefore vanishes quadratically at the endpoints.
8 Conclusion
Our analysis demonstrates that operationalizing the FA frontier with finite data poses unique challenges. Notably, departing from the population ideal introduces asymmetries across groups. These asymmetries grow with group heterogeneity, and such settings are precisely those in which fairness considerations are often most pressing. More specifically, we show that, when the covariance structure is known, the optimal sampling allocation prioritizes groups with greater variance or higher weight in the objective. When the covariance structure is unknown, however, a new challenge emerges: the bias term, which motivates balancing sample sizes across groups. In both settings, failing to account for these asymmetries can yield solutions that deviate meaningfully from the intended population-level trade-off.
These results open several avenues for further work. First, extending the finite-sample analysis to nonlinear and nonparametric models would provide valuable insight into other model classes. Second, incorporating multiple groups and group label uncertainty would broaden our ability to approach complex tradeoffs with data. Third, studying these questions in a dynamic setting by considering feedback loops where the model may itself shape future populations and data availability remains an important open question. Overall, our results take a step toward understanding how data limitations and statistical uncertainty shape our ability to effectively balance fairness and accuracy.
9 Acknowledgments
We are grateful to Annie Liang for insightful comments and discussion. Annie Ulichney’s work is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. We also acknowledge funding from the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the National Science Foundation, the European Union or the European Research Council Executive Agency.
Appendix A Proofs
A.1 Auxiliary Results
Lemma 2 (Exact risk of the OLS convex combination estimator).
Under Definition 3 and Assumptions 1, 2 the excess risk decomposes as:
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ $=\lambda^{2}\frac{1}{n_{r}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{r}}\sigma_{r}^{2}(X_{i})\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]$
$+(1-\lambda)^{2}\frac{1}{n_{b}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{b}}\sigma_{b}^{2}(X_{i})\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]$
$+\quad\mathbb{E}\left[\left\lVert\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right].$
Proof of Lemma 2
We begin by noting that
$\widehat{\nu}_{g}\coloneqq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}Y_{i}=\widehat{\Sigma}_{g}\beta_{g}+\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}\varepsilon_{i}X_{i}.$
Substituting into $\widehat{\nu}_{\lambda}$, we obtain
$\lambda\widehat{\nu}_{r}+(1-\lambda)\widehat{\nu}_{b}=\lambda\widehat{\Sigma}_{r}\beta_{r}+(1-\lambda)\widehat{\Sigma}_{b}\beta_{b}+\lambda\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}\varepsilon_{i}X_{i}+(1-\lambda)\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\varepsilon_{i}X_{i}.$
Define the symmetric matrix
$W\coloneqq\widehat{\Sigma}_{\lambda}^{-1}\Sigma_{g}\widehat{\Sigma}_{\lambda}^{-1}$
and let
$Z_{r}\coloneqq\lambda\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}\varepsilon_{i}X_{i},\quad Z_{b}\coloneqq(1-\lambda)\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}\varepsilon_{i}X_{i},\quad Z\coloneqq Z_{r}+Z_{b}$
$B=\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda}).$
The error in estimating $\beta_{\lambda}$ thus decomposes as
$\widehat{\beta}_{\lambda}-\beta_{\lambda}=\widehat{\Sigma}_{\lambda}^{-1}(Z+B).$
so that
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]=\mathbb{E}\left[\left\lVert Z+B\right\rVert_{W}^{2}\right]=\mathbb{E}\left[\left\lVert Z\right\rVert_{W}^{2}\right]+\mathbb{E}\left[\left\lVert B\right\rVert_{W}^{2}\right]+2\mathbb{E}\left[\langle Z,B\rangle_{W}\right].$
Since $Z$ is mean-zero conditional on $\{X_{i}\}$ and $B$ and $W$ are deterministic given $\{X_{i}\}$, we have
$\mathbb{E}\left[\mathbb{E}\left[\langle Z,B\rangle_{W}\right|\{X_{i}\}\right]=0.$
The cross-terms between $Z_{r}$ and $Z_{b}$ also vanish by independence and the fact that the noise is zero-mean:
$\mathbb{E}\left[\mathbb{E}\left[\left.\langle Z_{r},Z_{b}\rangle_{W}\right]\right|\{X_{i}\}\right]=0.$
Therefore,
$\mathbb{E}\left[\left\lVert Z\right\rVert_{W}^{2}\right]$ $=\mathbb{E}\left[\left\lVert Z_{r}\right\rVert_{W}^{2}\right]+\mathbb{E}\left[\left\lVert Z_{b}\right\rVert_{W}^{2}\right]$
$=\lambda^{2}\frac{1}{n_{r}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{r}}\sigma_{r}^{2}(X_{i})\left\lVert X_{i}\right\rVert_{W}^{2}\right]+(1-\lambda)^{2}\frac{1}{n_{b}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{b}}\sigma_{b}^{2}(X_{i})\left\lVert X_{i}\right\rVert_{W}^{2}\right]$
Combining terms and observing that
$\left\lVert X_{i}\right\rVert_{W}^{2}=\left\lVert\tilde{X}_{i}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}$
yields the claim. $\blacksquare$
Lemma 3 (mourtada2022exact Corollary 4 generalized to spherical covariance).
Suppose that $n_{g}\geq 48/\alpha_{g}$ and that $X^{(g)}$ satisfies 3 with parameters $C_{g}$ and $\alpha_{g}$. Also suppose that $\rho_{g}^{2}I_{d}\preceq\Sigma_{g}$. Then,
$\mathbb{E}\left[\lambda_{\min}(\widehat{\Sigma}_{g})^{-4}\right]\leq 2\rho_{g}^{-8}\cdot{{C}_{g}^{\prime}}^{4}$ (36)
where $C^{\prime}_{g}\coloneq 3C_{g}^{4}\exp(1+9/\alpha_{g})$.
Proof of Lemma 3
Define the whitened covariate vectors and whitened sample covariance, respectively, as
${\tilde{X}_{i}}^{(g)}\coloneqq\frac{1}{\rho_{g}}{X_{i}}^{(g)},\quad{\bar{\Sigma}}_{g}\coloneqq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}{\tilde{X}_{i}}^{(g)}{{\tilde{X}_{i}}^{(g)}}^{\top}.$
By mourtada2022exact Corollary 4 taking $q=4$, if $n_{g}\geq 48/\alpha_{g}$,
$\mathbb{E}\left[\lambda_{\min}(\bar{\Sigma}_{g})^{-4}\right]\leq 2\cdot\left(3C_{g}^{4}\exp(1+9/\alpha_{g})\right)^{4}.$
Hence,
$\mathbb{E}\left[\lambda_{\min}(\widehat{\Sigma}_{g})^{-4}\right]\leq 2\rho_{g}^{-8}\cdot\left(3C_{g}^{4}\exp(1+9/\alpha_{g})\right)^{4}.\blacksquare$ (37)
Lemma 4 (Subgaussian sample covariance moment bound).
Suppose Assumption 4 holds and $n_{g}\geq d$. Then, there exists a universal constant $C>0$ such that:
$\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert^{4}\right]\leq CK_{g}^{8}\left\lVert\Sigma_{g}\right\rVert^{4}\frac{d^{2}}{n_{g}^{2}}.$
Proof of Lemma 4
By standard subgaussian matrix concentration [see, e.g., vershynin2018high, Remark 4.7.3], for all $\delta\in(0,1)$, there exists an absolute constant $C>0$ such that:
$\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq CK_{g}^{2}\left(\sqrt{\frac{d+u}{n_{g}}}+\frac{d+u}{n_{g}}\right)\left\lVert\Sigma_{g}\right\rVert\right)\leq 2\exp(-u).$
To invert this tail bound, we seek $f(s)$ such that $\mathbb{P}(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s)\leq 2\exp(-f(s))$. In other words, using $2\max\{a,b\}\geq a+b$, we want
$s\leq CK_{g}^{2}\left(\sqrt{\frac{d+u}{n_{g}}}+\frac{d+u}{n_{g}}\right)\left\lVert\Sigma_{g}\right\rVert\leq 2CK_{g}^{2}\max\left\{\sqrt{\frac{d+u}{n_{g}}},\frac{d+u}{n_{g}}\right\}\left\lVert\Sigma_{g}\right\rVert$ (38)
We split this into two cases:
  1. 1.
    If $\sqrt{\frac{d+u}{n_{g}}}\geq\frac{d+u}{n_{g}},$ (38) is satisfied if $s\leq 2CK_{g}^{2}\sqrt{\frac{d+u}{n_{g}}}\left\lVert\Sigma_{g}\right\rVert$ which occurs when $u\geq\frac{s^{2}n_{g}}{4C^{2}K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}}-d$.
  2. 2.
    If $\sqrt{\frac{d+u}{n_{g}}}<\frac{d+u}{n_{g}},$ (38) is satisfied if $s\leq 2CK_{g}^{2}{\frac{d+u}{n_{g}}}\left\lVert\Sigma_{g}\right\rVert$ which occurs when $u\geq\frac{sn_{g}}{2CK_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}-d$.
Combining,
$\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)\leq 2\exp\left(d-C^{\prime}n_{g}\min\left\{\frac{s^{2}}{K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}},\frac{s}{K_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}\right\}\right).$
Define $s_{1}\coloneqq K_{g}^{2}\left\lVert\Sigma_{g}\right\rVert\sqrt{\frac{d}{C^{\prime}n_{g}}}$ and $s_{2}\coloneqq K_{g}^{2}\left\lVert\Sigma_{g}\right\rVert\frac{d}{C^{\prime}n_{g}}$, the critical values of $s$ that make the argument of the exponent positive for cases (1) and (2) defined above, respectively.
Next, we express the expectation as the integral of the tails, performing the change of variables $t=s^{4}$:
$\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert^{4}\right]$ $=\int_{0}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert^{4}\geq t\right)dt=4\int_{0}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds$
Define the critical point $s_{d}\coloneqq\max\{s_{1},s_{2}\}$ and split the integral
$\int_{0}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds$ $=\int_{0}^{s_{d}}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds+\int_{s_{d}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds$
For the first term, we may take the trivial upper bound of $1$ to bound the integral as follows:
$\int_{0}^{s_{d}}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq\int_{0}^{s_{d}}s^{3}ds\leq\frac{1}{4}\max\{s_{1}^{4},s_{2}^{4}\}=\frac{K_{g}^{8}\left\lVert\Sigma_{g}\right\rVert^{4}}{4}\max\left\{\left(\frac{d}{C^{\prime}n_{g}}\right)^{2},\left(\frac{d}{C^{\prime}n_{g}}\right)^{4}\right\}$
For the second term, we consider two cases of $s_{d}=s_{1}$ and $s_{d}=s_{2}$:
$\int_{s_{1}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq 2\int_{s_{1}}^{\infty}s^{3}\exp\left(d-\frac{C^{\prime}n_{g}s^{2}}{K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}}\right)ds=\frac{K_{g}^{8}(d+1)\left\lVert\Sigma_{g}\right\rVert^{4}}{C^{\prime 2}n_{g}^{2}}$
$\int_{s_{2}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq 2\int_{s_{2}}^{\infty}s^{3}\exp\left(d-\frac{C^{\prime}n_{g}s}{K_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}\right)ds=\frac{2K_{g}^{8}\left(d^{3}+3d^{2}+6d+6\right){\left\lVert\Sigma_{g}\right\rVert}^{4}}{C^{\prime 4}n_{g}^{4}}.$
Therefore, in either case, using $n_{g}\leq d$, we conclude that, for an absolute constant $C>0$,
$\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert^{4}\right]\leq CK_{g}^{8}\left\lVert\Sigma_{g}\right\rVert^{4}\frac{d^{2}}{n_{g}^{2}}$
as desired. $\blacksquare$
Lemma 5.
Under 4, there exists a constant $C>0$ such that
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{2}\right)\right]\leq\operatorname{Tr}\left(\Sigma_{g}^{2}\right)+\frac{CK_{g}^{4}}{n_{g}}\operatorname{Tr}\left(\Sigma_{g}\right)^{2}.$
Proof of Lemma 5
For notational convenience, we write $X_{i}$ in place of $X_{i}^{(g)}\sim X|G=g$ throughout the proof. Applying linearity of the trace and expectation and expanding the empirical covariance matrix product, we have:
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{2}\right)\right]=\frac{1}{n_{g}^{2}}\sum_{i=1}^{n_{g}}\operatorname{Tr}\left(\mathbb{E}[X_{i}X_{i}^{\top}X_{i}X_{i}^{\top}]\right)+\frac{1}{n_{g}^{2}}\sum_{i\neq j}\operatorname{Tr}\left(\mathbb{E}[X_{i}X_{i}^{\top}X_{j}X_{j}^{\top}]\right)$
For the cross-terms, by independence, for $i\neq j$,
$\mathbb{E}\left[X_{i}{X_{i}}^{\top}X_{j}{X_{j}}^{\top}\right]=\mathbb{E}\left[X_{i}{X_{i}}^{\top}\right]\mathbb{E}\left[X_{j}{X_{j}}^{\top}\right]=\Sigma_{g}^{2}.$
Thus,
$\sum_{i\neq j}\operatorname{Tr}\left(\mathbb{E}[X_{i}X_{i}^{\top}X_{j}X_{j}^{\top}]\right)=n_{g}(n_{g}-1)\operatorname{Tr}\left(\Sigma_{g}^{2}\right).$
For the diagonal terms, interchanging the trace and expectation and applying invariance under cyclic permutations of the trace, we have
$\operatorname{Tr}\left(\mathbb{E}\left[X_{i}{X_{i}}^{\top}X_{i}{X_{i}}^{\top}\right]\right)=\mathbb{E}\left[\left\lVert X_{i}\right\rVert_{2}^{4}\right].$
By our subgaussianity assumption (Assumption 4) and standard moment bounds for subgaussian vectors [see, e.g., vershynin2018high], we have, for some absolute constant $C>0$:
$\mathbb{E}\left[\left\lVert X_{1}\right\rVert^{4}\right]\leq CK_{g}^{4}\operatorname{Tr}\left(\Sigma_{g}\right)^{2}.$
Putting these results together yields the claim. $\blacksquare$
Lemma 6 (Assouad’s Lemma ).
Fix an integer $m\geq 1$ and define the $m$-dimensional hypercube $\Xi_{m}\coloneqq\{-1,1\}^{m}$. Let $\mathcal{P}=\left\{\mathbb{P}_{\xi}:\xi\in\Xi_{m}\right\}$ be a family of probability measures such that for each $\xi\in\Xi_{m}$ there is an associated parameter $\theta(\mathbb{P}_{\xi})$ that is the target of interest. Also suppose that $|\mathcal{P}|=2^{m}$. Take the loss function to be the squared $\ell_{2}$-distance. For $j\in[m]$, let $\xi^{(j)}$ denote the vector obtained by flipping the $j$-th coordinate of vector $\xi$, i.e., $\xi_{-j}=\xi^{(j)}_{-j}$, $\xi_{j}=-\xi^{(j)}_{j}$. Define the per-coordinate separation as
$\alpha\coloneqq\inf_{j\in[m],\xi\in\Xi_{m}}\left\langle e_{j},\theta(\mathbb{P}_{\xi})-\theta\left(\mathbb{P}_{\xi^{(j)}}\right)\right\rangle^{2}.$
Then, for every estimator $\widehat{\theta}$, the minimax risk is lower bounded by
$\sup_{\mathbb{P}_{\xi}\in\mathcal{P}}\mathbb{E}_{\mathbb{P}_{\xi}}\left[\left\lVert\widehat{\theta}-\theta(\mathbb{P}_{\xi})\right\rVert_{2}^{2}\right]\geq\frac{m}{2}\cdot\alpha\cdot\left(1-\sup_{\begin{subarray}{c}\xi,\xi^{\prime}\in\Xi_{m}\\ d_{h}(\xi,\xi^{\prime})=1\end{subarray}}\left\lVert\mathbb{P}_{\xi},\mathbb{P}_{\xi^{\prime}}\right\rVert_{\text{TV}}\right)$
where $d_{h}$ denotes the Hamming distance.
A.2 Proofs from Section 3
A.2.1 Proof of Lemma 1
For any group $g\in\{r,b\}$, the risk of the group-optimal predictor $\beta_{g}$ is:
$\mathcal{R}_{g}(\beta_{g})=\mathbb{E}_{g}\left[(X^{\top}\beta_{g}-Y)^{2}\right]=\mathbb{E}_{g}[\varepsilon_{g}^{2}]=\sigma^{2}.$
For the other group $g^{\prime}\neq g$, the same predictor incurs risk:
$\mathcal{R}_{g^{\prime}}(\beta_{g})=\mathbb{E}_{g^{\prime}}\left[(X^{\top}\beta_{g}-Y)^{2}\right]=\mathbb{E}_{g^{\prime}}\left[\left(X^{\top}\left(\beta_{g}-\beta_{g^{\prime}}\right)\right)^{2}\right]+\sigma^{2}.$
The latter is greater or equal, and strictly greater when $\beta_{g}\neq\beta_{g’}$ and $\Sigma_{g’}$ is invertible.
A.2.2 Proof of Proposition 1
Note that
$\mathcal{R}_{g}(\beta)=\mathbb{E}_{g}\left[(X^{\top}\beta-Y)^{2}\right]=\mathbb{E}_{g}\left[(X^{\top}(\beta-\beta_{g})-\varepsilon_{g})^{2}\right]=\|\beta-\beta_{g}\|_{\Sigma_{g}}^{2}+\sigma^{2},$ (39)
where the last equality follows from the fact that $\mathbb{E}[\varepsilon_{g}|X]=0$. As a result, we have
$\mathcal{R}_{g}(\beta)-\mathcal{R}_{g}(\beta_{\lambda})=\|\beta-\beta_{g}\|_{\Sigma_{g}}^{2}-\|\beta_{\lambda}-\beta_{g}\|_{\Sigma_{g}}^{2}=\|\beta-\beta_{\lambda}\|_{\Sigma_{g}}^{2}+2(\beta-\beta_{\lambda})^{\top}\Sigma_{g}(\beta_{\lambda}-\beta_{g}),$ (40)
which completes the proof of (9b). To establish (9a), simply sum both sides of (9b) for the red and blue groups, weighted by $\lambda$ and $1-\lambda$, respectively, and use the definition of $\beta_{\lambda}$.
A.3 Proofs from Section 4
A.3.1 Proof of Theorem 1
Upper bound
We first establish that the error given in the statement of the theorem is indeed achieved by setting $\beta=\tilde{\beta}_{\lambda}$. Notice that, we have
$\tilde{\beta}_{\lambda}=\beta_{\lambda}-\Sigma_{\lambda}^{-1}\left(\lambda\Sigma_{r}(\beta_{r}-\widehat{\beta}_{r})+(1-\lambda)\Sigma_{b}(\beta_{b}-\widehat{\beta}_{b})\right).$
Substituting $\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}$ and using the decomposition
$\widehat{\nu}_{g}=\widehat{\Sigma}_{g}\beta_{g}+\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}^{(g)}\varepsilon_{i}^{(g)},$
we obtain
$\tilde{\beta}_{\lambda}-\beta_{\lambda}=\Sigma_{\lambda}^{-1}\left(\lambda\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\cdot\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)\Sigma_{b}\widehat{\Sigma}_{b}^{-1}\cdot\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right).$
Let $A\coloneqq{\Sigma}_{\lambda}^{-1}\Sigma_{g}{\Sigma}_{\lambda}^{-1}$. By the mean-zero and independence properties of the noise, the cross-terms cancel in the squared norm, which yields
$\mathbb{E}$ $\left[\left\lVert\tilde{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ $=\lambda^{2}\cdot\frac{1}{n_{r}^{2}}\mathbb{E}\left[\left\lVert\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}\right\rVert_{A}^{2}\right]+(1-\lambda)^{2}\cdot\frac{1}{n_{b}^{2}}\mathbb{E}\left[\left\lVert\Sigma_{b}\widehat{\Sigma}_{b}^{-1}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right\rVert_{A}^{2}\right]$
Using the independence and mean-zero property of $\varepsilon_{i}^{(g)}$ and applying the trace identity for the expectation of quadratic forms, we obtain
$\mathbb{E}\left[\left\lVert\Sigma_{g}\widehat{\Sigma}_{r}^{-1}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}\right\rVert_{A}^{2}\right]$ $=\sigma^{2}n_{r}\cdot\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{r}^{-1}\Sigma_{r}A\Sigma_{r}\right)\right].$
Applying the same procedure for the equivalent term for group $b$ and substituting completes the proof.
Lower bound
Next, we establish the minimax optimality. In particular, we aim to show
$\sup_{P\in\mathcal{P}_{\text{linear}}(P_{x}^{r},P_{X}^{b},\sigma^{2})}$ $\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$
$\geq\lambda^{2}\frac{\sigma^{2}}{n_{r}}\mathbb{E}\left[\operatorname{Tr}\left(\Sigma_{g}\Sigma_{\lambda}^{-1}\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\Sigma_{r}\Sigma_{\lambda}^{-1}\right)\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\mathbb{E}\left[\operatorname{Tr}\left(\Sigma_{g}\Sigma_{\lambda}^{-1}\Sigma_{b}\widehat{\Sigma}_{b}^{-1}\Sigma_{b}\Sigma_{\lambda}^{-1}\right)\right],$
for any $\beta$. We derive this lower bound by the Bayes risk under an appropriate prior. More specifically, suppose we choose a prior over groups’ regressor $\beta_{r}$ and $\beta_{b}$. Then, the Bayes risk provides a valid lower bound:
$\inf_{\beta}\sup_{\beta_{\lambda}}\mathbb{E}\left[\left\lVert\beta-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\geq\mathbb{E}\left[\left\lVert\widehat{\beta}^{*}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}\right],$ (41)
where $\widehat{\beta}^{*}_{\lambda}$ denotes the Bayes estimator under a prior specified below.
Let $\tau>0$ and suppose that the group-specific parameters follow independent Gaussian priors:
$\beta_{r}\sim\mathcal{N}(0,\tau^{2}I_{d}),\quad\beta_{b}\sim\mathcal{N}(0,\tau^{2}I_{d}).$
Then, the induced prior on $\beta_{\lambda}$ is also Gaussian:
$\beta_{\lambda}\sim\mathcal{N}\left(0,\Sigma_{\lambda}^{-1}\left(\lambda^{2}\tau^{2}\Sigma_{r}^{2}+(1-\lambda)^{2}\tau^{2}\Sigma_{b}^{2}\right)\Sigma_{\lambda}^{-1}\right).$
Under the linear model $Y=X^{\top}\beta_{g}+\varepsilon_{g}$ with i.i.d noise $\varepsilon_{g}\sim\mathcal{N}(0,\sigma^{2})$, the posterior distribution for each parameter $\beta_{g}|\left\{X_{i}^{(g)},Y_{i}^{(g)}\right\}$ is Gaussian:
$\beta_{g}|\left\{X_{i}^{(g)},Y_{i}^{(g)}\right\}\sim\mathcal{N}\left(\mu_{g}(\tau),V_{g}(\tau)\right)$
with posterior mean and variance:
$M_{g}(\tau)\coloneqq\widehat{\Sigma}_{g}+\frac{\sigma^{2}}{n_{g}\tau^{2}}I_{d},\quad\mu_{g}(\tau)\coloneqq{M_{g}(\tau)}^{-1}\widehat{\nu}_{g},\quad V_{g}(\tau)\coloneqq\frac{\sigma^{2}}{n_{g}}{M_{g}(\tau)}^{-1}.$
Since $\beta_{\lambda}$ is a deterministic linear combination of $\beta_{r}$ and $\beta_{b}$, its posterior is also Gaussian:
$\beta_{\lambda}\mid\mathcal{S}_{n}\sim\mathcal{N}\left(\mu_{\lambda}(\tau),V_{\lambda}(\tau)\right)$
where:
$\mu_{\lambda}(\tau)\coloneqq\Sigma_{\lambda}^{-1}\left(\lambda\Sigma_{r}\mu_{r}(\tau)+(1-\lambda)\Sigma_{b}\mu_{b}(\tau)\right),\quad V_{\lambda}\coloneqq\Sigma_{\lambda}^{-1}\left(\lambda^{2}\Sigma_{r}V_{r}(\tau)\Sigma_{r}+(1-\lambda)^{2}\Sigma_{b}V_{b}(\tau)\Sigma_{b}\right)\Sigma_{\lambda}^{-1}.$
Since the Bayes estimator under squared loss is the posterior mean, the Bayes risk in the $\Sigma_{g}$-norm is:
$\mathbb{E}\left[\left\lVert\widehat{\beta}^{*}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}\right]=\mathbb{E}\left[\operatorname{Tr}\left(\Sigma_{g}V_{\lambda}(\tau)\right)\right]$
To obtain a lower bound, we consider the limit $\tau\to\infty$, which corresponds to an uninformative prior. In this limit:
$\mu_{g}(\tau)\to\widehat{\Sigma}^{-1}\widehat{\nu}_{g},\quad V_{g}(\tau)\to\frac{\sigma^{2}}{n_{g}}\widehat{\Sigma_{g}}^{-1}.$
Thus, $V_{\lambda}(\tau)$ increases in the Loewner order as $\tau\to\infty$, and by monotone convergence,
$\lim_{\tau\to\infty}\mathbb{E}\left[\operatorname{Tr}\left(\Sigma_{g}V_{\lambda}(\tau)\right)\right]=\lambda^{2}\frac{\sigma^{2}}{n_{r}}\mathbb{E}\operatorname{Tr}\left(\Sigma_{g}\Sigma_{\lambda}^{-1}\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\Sigma_{r}\Sigma_{\lambda}^{-1}\right)+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\mathbb{E}\operatorname{Tr}\left(\Sigma_{g}\Sigma_{\lambda}^{-1}\Sigma_{b}\widehat{\Sigma}_{b}^{-1}\Sigma_{b}\Sigma_{\lambda}^{-1}\right).$
The desired minimax lower bound follows from substitution into the inequality (41).
A.3.2 Proof of Corollary 2
Under the group-wise spherical covariance assumption, we have
$\Sigma_{\lambda}^{-1}=\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{-1}I_{d}.$
Substituting and applying linearity of the trace yields
$\mathbb{E}\left[\left\lVert\tilde{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ $\leq\lambda^{2}\frac{\sigma^{2}}{n_{r}}\frac{\rho_{g}^{2}\rho_{r}^{4}}{(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2})^{2}}\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{r}^{-1}\right)\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\frac{\rho_{g}^{2}\rho_{b}^{4}}{(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2})^{2}}\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{b}^{-1}\right)\right].$ (42)
Let ${\tilde{X}_{i}}^{(g)}$ denote the whitened covariate vectors, i.e., ${\tilde{X}_{i}}^{(g)}=\rho_{g}^{1/2}{X}_{i}^{(g)}$ such that $\mathbb{E}\left[{\tilde{X}_{i}}^{(g)}{\tilde{X}_{i}^{(g)}}^{\top}\right]=I_{d}$.
Also observe
$\operatorname{Tr}\left(\tilde{\Sigma}_{g}^{-1}\right)\leq d\cdot\lambda_{\min}\left(\tilde{\Sigma}_{g}\right)^{-1}.$
By Corollary 4 of mourtada2022exact,
$\mathbb{E}\left[\lambda_{\min}\left(\tilde{\Sigma}_{g}\right)^{-1}\right]\leq 2C_{g}^{\prime}.$
Then, noting $\operatorname{Tr}\left(\tilde{\Sigma}_{g}\right)=\rho_{g}^{2}\operatorname{Tr}\left(\widehat{\Sigma}_{g}\right),$ we reach the claim by substituting the resulting bound
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{-1}\right)\right]\leq 2C_{g}^{\prime}\rho_{g}^{-2}d$ (43)
into (42).
To show the second half of the claim, consider subgaussian $X\in\mathbb{R}^{d}$ satisfying 4 with parameter $K_{g}$ and with covariance $\rho_{g}^{2}I_{d}$. Each coordinate satisfies $\left\lVert X_{i}\right\rVert_{\psi_{2}}\leq K_{g}\rho_{g}$ and it follows that there exists an absolute constant $\zeta_{0}$ such that $\mathbb{E}\left[X_{i}^{4}\right]\leq\zeta_{0}K_{g}^{4}\rho_{g}^{4}$. By Cauchy-Schwarz, for $i\neq j$, $\mathbb{E}\left[X_{i}^{2}X_{j}\right]\leq\zeta_{0}K_{g}^{4}\rho_{g}^{4}$. Hence, there exists a constant $\zeta>0$ such that
$\mathbb{E}\left[\left\lVert X\right\rVert^{4}\right]=d\cdot\mathbb{E}\left[X_{1}^{4}\right]+2\binom{d}{2}\cdot\mathbb{E}\left[X_{1}^{2}X_{2}^{2}\right]\leq\zeta K_{g}^{4}\rho_{g}^{4}.$
Then, by whitening and applying mourtada2022exact, we replace (43) with the following bound:
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{-1}\right)\right]\leq\rho_{g}^{-2}\left(d+\frac{8C_{g}^{\prime}\zeta\rho_{g}^{8}d^{2}}{n_{g}}\right)$
and the claim follows. $\blacksquare$
A.3.3 Proof of Proposition 2
Let us define $n$ as the minimum of $n_{r}$ and $n_{b}$. Notice that the empirical estimator is given by
$\tilde{\beta}_{\lambda}=\beta_{\lambda}+\Sigma_{\lambda}^{-1}\left(\lambda\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)\widehat{\Sigma}_{b}^{-1}\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right).$ (44)
Next, using a variant of the Woodbury identity, $(A-B)^{-1}=A^{-1}+A^{-1}B(A-B)^{-1}$, we obtain
$\widehat{\Sigma}_{g}^{-1}=\Sigma_{g}^{-1}+\Sigma_{g}^{-1}(\Sigma_{g}-\widehat{\Sigma}_{g})\widehat{\Sigma}_{g}^{-1}.$ (45)
Plugging (45) into (44) yields
$\tilde{\beta}_{\lambda}-\beta_{\lambda}$ $=\Sigma_{\lambda}^{-1}\left(\lambda\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right)$ (46)
$+\Sigma_{\lambda}^{-1}\left(\lambda(\Sigma_{r}-\widehat{\Sigma}_{r})\widehat{\Sigma}_{r}^{-1}\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)(\Sigma_{b}-\widehat{\Sigma}_{b})\widehat{\Sigma}_{b}^{-1}\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right).$ (47)
Now, note that the expectation of the norm of the second term (47) is bounded by $\mathcal{O}(1/n)$. To see this, note that, for any $g\in\mathcal{G}$, we have
$\mathbb{E}\left[\left\|(\Sigma_{g}-\widehat{\Sigma}_{g})\widehat{\Sigma}_{g}^{-1}\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}^{(g)}\varepsilon_{i}^{(g)}\right\|\right]$
$\leq\quad\sqrt{\mathbb{E}\left[\left\|\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}X_{i}^{(g)}\varepsilon_{i}^{(g)}\right\|^{2}\right]}\left(\mathbb{E}\left[\left\|(\Sigma_{g}-\widehat{\Sigma}_{g})\right\|^{4}\right]\right)^{1/4}\left(\mathbb{E}\left[\left\|\widehat{\Sigma}_{g}^{-1}\right\|^{4}\right]\right)^{1/4},$ (48)
and the first term is bounded by $\mathcal{O}(1/\sqrt{n})$ since it has zero mean, the second term is bounded by $\mathcal{O}(1/\sqrt{n})$ by Lemma 4, and, finally, the last term is bounded by a constant due to Lemma 3. Now, this, along with Proposition 1 and Corollary 2 implies that we can write per-estimation group error
$\mathcal{R}_{g}(\tilde{\beta}_{\lambda})-\mathcal{R}_{g}(\beta_{\lambda})$ (49)
as
$Z_{1}+Z_{2}~~~\text{with }~~~Z_{1}=2\left(\lambda\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right)^{\top}\Sigma_{\lambda}^{-1}\Sigma_{g}(\beta_{\lambda}-\beta_{g}),$ (50)
and some variable $Z_{2}$ with $\mathbb{E}[|Z_{2}|]\leq\kappa/n$ for some constant $\kappa$.
Next, note that, by Markov’s inequality, the probability that $Z_{2}$ is greater than
$\frac{2\kappa}{n\delta}$
is at most $\delta/2$. Also, by the central limit theorem, $\sqrt{n}Z_{1}$ converges, in distribution, to a zero-mean normal distribution, and thus, there exists a constant $c_{1}>0$ and a threshold $\bar{n}_{1}$, such that for all $n\geq\bar{n}_{1}$,
$\mathbb{P}\left(Z_{1}\leq\frac{-c_{1}}{\sqrt{n}}\right)\geq\frac{1}{2}-\frac{\delta}{2}.$
As a result, with probability $1/2-\delta$ and for $n\geq\bar{n}_{1}$, we have both $Z_{2}\leq\frac{2\kappa}{n\delta}$ and $Z_{1}\leq\frac{-c_{1}}{\sqrt{n}}$. Finally, taking
$\bar{n}=\max\left\{\bar{n}_{1},\frac{4\kappa^{2}}{c_{1}^{2}\delta^{2}}\right\}$
completes the proof. $\blacksquare$
A.4 Proofs from Section 5
A.4.1 Proof of Proposition 3
From Lemma 2, the excess risk decomposes as:
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ $=\lambda^{2}\frac{\sigma^{2}}{n_{r}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{r}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{b}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]$
$\quad\quad+\mathbb{E}\left[\left\lVert\lambda\Sigma_{g}^{-1/2}\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\Sigma_{g}^{-1/2}\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]$
where we define:
$\tilde{X}_{i}=\Sigma_{g}^{-1/2}X_{i}\quad\text{and}\quad\tilde{\Sigma}_{g}=\Sigma_{g}^{-1/2}\widehat{\Sigma}_{\lambda}\Sigma_{g}^{-1/2}.$
We apply the identity:
$\sum_{i=1}^{n}\left\lVert Ax_{i}\right\rVert^{2}=\operatorname{Tr}\left(A^{\top}A\sum_{i=1}^{n}x_{i}x_{i}^{\top}\right)$
with $A=\tilde{\Sigma}_{g}^{-1}$ and $x_{i}=\tilde{X_{i}}$. This gives:
$\frac{\sigma^{2}}{n_{r}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{r}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]=\frac{\sigma^{2}}{n_{r}}\mathbb{E}\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-1}\Sigma_{g}\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{r}\right),\quad\frac{\sigma^{2}}{n_{b}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{b}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]=\frac{\sigma^{2}}{n_{b}}\mathbb{E}\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-1}\Sigma_{g}\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{b}\right)$
For the final bias term, observe that:
$\left\lVert\lambda\Sigma_{g}^{-1/2}\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\Sigma_{g}^{-1/2}\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}=\left\lVert\Sigma_{g}^{1/2}\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})\right)\right\rVert^{2},$
which gives the first characterization of the bias term. To see the other representation, note that
$\lambda\Sigma_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\Sigma_{b}(\beta_{b}-\beta_{\lambda})=0.$
Therefore, we have
$\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})=\lambda\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})+(1-\lambda)\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda}).$
Then, we can rearrange to see that
$\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})+(1-\lambda)\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda})\right)$ (51)
$\quad=\lambda\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{r}\beta_{r}+(1-\lambda)\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{b}\beta_{b}-\beta_{\lambda}$
$\quad=\lambda\left(\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{r}-\Sigma_{\lambda}^{-1}\Sigma_{r}\right)\beta_{r}+(1-\lambda)\left(\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{b}-\Sigma_{\lambda}^{-1}\Sigma_{b}\right)\beta_{b}.$ (52)
Next, note that, we have
$\lambda\left(\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{r}-\Sigma_{\lambda}^{-1}\Sigma_{r}\right)=\left(I_{d}+\frac{1-\lambda}{\lambda}\widehat{\Sigma}_{r}^{-1}\widehat{\Sigma}_{b}\right)^{-1}-\left(I_{d}+\frac{1-\lambda}{\lambda}{\Sigma}_{r}^{-1}{\Sigma}_{b}\right)^{-1}.$ (53)
Similarly, we can write
$(1-\lambda)\left(\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{b}-\Sigma_{\lambda}^{-1}\Sigma_{b}\right)=\left(I_{d}+\frac{\lambda}{1-\lambda}\widehat{\Sigma}_{b}^{-1}\widehat{\Sigma}_{r}\right)^{-1}-\left(I_{d}+\frac{\lambda}{1-\lambda}{\Sigma}_{b}^{-1}{\Sigma}_{r}\right)^{-1}.$ (54)
Next, using the Woodbury matrix identity, we have $(I_{d}+X)^{-1}=I_{d}-(I_{d}+X^{-1})^{-1}$ for any invertible $d\times d$ matrix $X$. Therefore, we can recast (54) as
$(1-\lambda)\left(\widehat{\Sigma}_{\lambda}^{-1}\widehat{\Sigma}_{b}-\Sigma_{\lambda}^{-1}\Sigma_{b}\right)=-\left(I_{d}+\frac{1-\lambda}{\lambda}\widehat{\Sigma}_{r}^{-1}\widehat{\Sigma}_{b}\right)^{-1}+\left(I_{d}+\frac{1-\lambda}{\lambda}{\Sigma}_{r}^{-1}{\Sigma}_{b}\right)^{-1}.$ (55)
Plugging (53) and (55) into (52) completes the proof. $\blacksquare$
A.4.2 Proof of Theorem 2
Upper bound
We show a slightly more general result for the upper bound. In fact, we relax the condition (26) to
$\rho_{g}^{2}I_{d}\preceq\Sigma_{g}\preceq\mathrm{P}_{g}^{2}I_{d},$ (56)
and also drop the assumption $n_{g}\geq K_{g}^{4}d$, and establish the upper bound
$\mathcal{V}_{g}(\lambda)\lesssim\frac{\mathrm{P}_{g}^{2}\sigma^{2}}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}}\left[\lambda^{2}\mathrm{P}_{r}^{2}\left(\frac{d}{n_{r}}+\frac{{K_{r}^{2}}{d}^{3/2}}{n_{r}^{3/2}}\right)+(1-\lambda)^{2}\mathrm{P}_{b}^{2}\left(\frac{d}{n_{b}}+\frac{{K_{b}^{2}}{d}^{3/2}}{{n_{b}}^{3/2}}\right)\right].$ (57)
To do so, first apply Proposition 3 under $\Sigma_{g}\preceq\mathrm{P}_{g}^{2}I_{d}$ to obtain the following bound:
$\mathcal{V}_{g}(\lambda)$ $\leq\mathrm{P}_{g}^{2}\left(\lambda^{2}\frac{\sigma^{2}}{n_{r}}\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-2}\widehat{\Sigma}_{r}\right)\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-2}\widehat{\Sigma}_{b}\right)\right]\right)$
Next, by Cauchy-Schwarz on the trace inner product and again on the expectation:
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-2}\widehat{\Sigma}_{g}\right)\right]\leq\mathbb{E}\left[\sqrt{\operatorname{Tr}\left(\ \widehat{\Sigma}_{\lambda}^{-4}\right)}\sqrt{\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{2}\right)}\right]\leq\sqrt{\mathbb{E}\left[\operatorname{Tr}\left(\ \widehat{\Sigma}_{\lambda}^{-4}\right)\right]\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{2}\right)\right]}.$
Under 4 that $X|G=g$ is $K_{g}$-subgaussian, by Lemma 5,
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{2}\right)\right]\leq\operatorname{Tr}\left(\Sigma_{g}^{2}\right)+\frac{CK_{g}^{4}}{n_{g}}\operatorname{Tr}\left(\Sigma_{g}\right)^{2}\lesssim\mathrm{P}_{g}^{4}\left(d+\frac{{K_{g}^{4}}d^{2}}{n_{g}}\right)$ (58)
where the second inequality uses that $\Sigma_{g}\preceq\mathrm{P}_{g}^{2}I_{d}$. Next, by Weyl’s inequality,
$\lambda_{\min}\left(\widehat{\Sigma}_{\lambda}\right)\geq\lambda\cdot\lambda_{\min}\left(\widehat{\Sigma}_{r}\right)+(1-\lambda)\cdot\lambda_{\min}\left(\widehat{\Sigma}_{g}\right).$ (59)
Therefore, we have
$\mathbb{E}\left[\lambda_{\min}\left(\widehat{\Sigma}_{\lambda}\right)^{-4}\right]\leq\min\left\{\lambda^{-4}\mathbb{E}\left[\lambda_{\min}\left(\widehat{\Sigma}_{r}\right)^{-4}\right],(1-\lambda)^{-4}\mathbb{E}\left[\lambda_{\min}\left(\widehat{\Sigma}_{b}\right)^{-4}\right]\right\}$ (60)
Now, by Lemma 3, we have
$\sqrt{\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-4}\right)\right]}\leq\sqrt{2}\sqrt{d}\min\left\{\lambda^{-2}{C_{r}^{\prime}}^{2}\rho_{r}^{-4},(1-\lambda)^{-2}{C_{b}^{\prime}}^{2}\rho_{b}^{-4}\right\}$ (61)
Combining the bounds (61) and (58) and using $\min\{1/a,1/b\}\leq 2/(a+b)$, we reach the bound:
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-2}\widehat{\Sigma}_{g}\right)\right]$ $\lesssim\mathrm{P}_{g}^{2}\frac{1}{\frac{\lambda^{2}}{{C}_{r}^{\prime 2}}\rho_{r}^{4}+\frac{(1-\lambda)^{2}}{{C}_{b}^{\prime 2}}\rho_{b}^{4}}\left(d+\frac{{K_{g}^{2}}d^{3/2}}{\sqrt{n_{g}}}\right)$ (62)
$\lesssim\mathrm{P}_{g}^{2}\frac{1}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}}\left(d+\frac{{K_{g}^{2}}d^{3/2}}{\sqrt{n_{g}}}\right)$ (63)
As a result, the variance terms satisfy the following bound:
$\mathrm{P}_{g}^{2}\left(\lambda^{2}\frac{\sigma^{2}}{n_{r}}\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-2}\widehat{\Sigma}_{r}\right)\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}}\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{\lambda}^{-2}\widehat{\Sigma}_{b}\right)\right]\right)$
$\lesssim\frac{\mathrm{P}_{g}^{2}\sigma^{2}}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}}\left[\lambda^{2}\mathrm{P}_{r}^{2}\left(\frac{d}{n_{r}}+\frac{{K_{r}^{2}}{d}^{3/2}}{n_{r}^{3/2}}\right)+(1-\lambda)^{2}\mathrm{P}_{b}^{2}\left(\frac{d}{n_{b}}+\frac{{K_{b}^{2}}{d}^{3/2}}{{n_{b}}^{3/2}}\right)\right].$
Under the assumption $n_{g}\geq K_{g}^{4}d$, the term $K_{g}^{2}d^{3/2}/n_{g}^{3/2}$ is bounded by $d/n_{g}$ which completes the proof of the upper bound.
Lower bound
We use the Assouad’s Lemma (Lemma 6). Here are the steps:
Perturbed parameter structure:
For $g\in\{r,b\}$, set
$h_{g}^{2}=\frac{\sigma^{2}}{4n_{g}\rho_{g}^{2}},\quad\beta_{g}^{(\xi)}\coloneqq h_{g}\xi_{g}$
where $\xi_{g}\in\Xi^{d}$. The condition on $n_{g}$ ensures $\|\beta_{g}^{(\xi)}\|\leq B$.
Consider the $2d$-dimensional hybercube $\Xi\coloneqq\Xi^{d}\times\Xi^{d}$. For $\xi\coloneqq(\xi_{r},\xi_{b})\in\Xi$, define by $\mathbb{P}^{(\xi)}$ the joint law of the data with mean $\beta^{(\xi)}\coloneqq(\beta_{r}^{(\xi_{r})},\beta_{b}^{(\xi_{b})})$.
KL divergence bound:
The following result provides an upper bound on the KL divergence between the distributions under parameters $\xi,\xi^{\prime}\in\Xi$ differing in only a single coordinate.
Claim 1.
Let $\xi$ and $\xi^{\prime}$ differ only in coordinate $i\in[d]$ (i.e., a coordinate corresponding to group $r$) where $\xi_{i}=-\xi_{i}^{\prime}$. Then
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})\leq\frac{2h_{r}^{2}\rho_{r}^{2}}{\sigma^{2}}$
Proof of 1.
To see this, take $\xi$ and $\xi^{\prime}$ to be neighbors that differ only in the $i$-th coordinate of $\xi_{r}$, i.e., $\xi_{r,i}=1,\xi_{r,i}=-1$ and $\xi_{r,-i}=\xi^{\prime}_{r,-i}$. Recalling well-known results for the KL divergence of multivariate Gaussians [see, e.g., zhang2023properties], conditional on observing a single observation $X_{i}$, the KL divergence between $\mathbb{P}_{\xi_{r}}=\mathcal{N}(h_{r}X_{i},\sigma^{2}),\mathbb{P}_{\xi^{\prime}_{r}}=\mathcal{N}(-h_{r}X_{i},\sigma^{2})$ is given by
$D_{\text{KL}}(\mathbb{P}_{\xi_{r}}\;\|\;\mathbb{P}_{\xi^{\prime}_{r}}\mid X_{i})=\frac{2h_{r}^{2}X_{i}^{2}}{\sigma^{2}}.$
Then, for $X_{i}\sim\mathcal{N}(0,\rho_{r}^{2})$ with $\mathbb{E}[X_{i}^{2}]=\rho_{r}^{2}$, we observe that
$D_{\text{KL}}(\mathbb{P}_{\xi_{r}}\;\|\;\mathbb{P}_{\xi^{\prime}_{r}})=\frac{2h_{r}^{2}\rho_{r}^{2}}{\sigma^{2}}.$
By 1, over $n_{r}$ samples,
$D_{\text{KL}}(\mathbb{P}^{(\xi)\otimes n_{r}}\;\|\;\mathbb{P}^{(\xi^{\prime})\otimes n_{r}})\leq\frac{2h_{r}^{2}\rho_{r}^{2}n_{r}}{\sigma^{2}}=\frac{1}{2}.$
By Pinsker’s inequality,
$\left\lVert\mathbb{P}_{\xi},\mathbb{P}_{\xi^{\prime}}\right\rVert_{\text{TV}}\leq\frac{1}{2}.$
Parameter separation:
Let $\xi$ and $\xi^{\prime}$ differ only in the same coordinate $i\in[d]$ as in 1 where $\xi_{i}=-\xi_{i}^{\prime}$. Then, we can express
$\langle e_{i},\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}\rangle=\frac{\lambda\rho_{r}^{2}\left(h_{r}-(-h_{r})\right)}{\rho_{\lambda}^{2}}=\frac{2\lambda\rho_{r}^{2}h_{r}}{\rho_{\lambda}^{2}},$ (64)
with
$\rho_{\lambda}^{2}=\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}.$ (65)
After an analogous computation for a single-coordinate group $b$ perturbation,
$\langle e_{i},\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}\rangle=\frac{2(1-\lambda)\rho_{b}^{2}h_{b}}{\rho_{\lambda}^{2}}.$ (66)
Substitute the definitions of $h_{r}$ and $h_{b}$ into (64) and (66), respectively, and define
$\alpha_{r}^{2}\coloneqq\frac{\sigma^{2}\lambda^{2}\rho_{r}^{2}}{n_{r}\rho_{\lambda}^{4}},\quad\alpha_{b}^{2}\coloneqq\frac{\sigma^{2}(1-\lambda)^{2}\rho_{b}^{2}}{n_{b}\rho_{\lambda}^{4}}.$
Assouad’s Lemma:
Over a $2d$-dimensional signed hypercube, an application of Assouad’s lemma (as stated in 6) and the fact $\max\{a,b\}\geq\frac{1}{2}(a+b)$ yields the bound
$\inf_{\beta}\sup_{\xi\in\Xi}\mathbb{E}_{\xi}\left[\left\lVert\beta_{\lambda}-\beta_{\lambda}^{(\xi)}\right\rVert_{2}^{2}\right]\geq\frac{1}{16}\frac{\sigma^{2}d}{\rho_{\lambda}^{4}}\left(\frac{\lambda^{2}\rho_{r}^{2}}{n_{r}}+\frac{(1-\lambda)^{2}\rho_{b}^{2}}{n_{b}}\right).$
and the final bound follows by observing
$\left\lVert\beta_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\geq\rho_{g}^{2}\left\lVert\beta_{\lambda}-\beta_{\lambda}\right\rVert_{2}^{2}.\quad\blacksquare$
A.4.3 Proof of Theorem 3
Upper Bound
We again show a slightly more general result for the upper bound. In fact, we relax the (26) to
$\rho_{g}^{2}I_{d}\preceq\Sigma_{g}\preceq\mathrm{P}_{g}^{2}I_{d},$ (67)
and show the upper bound
$\mathcal{B}_{g}(\lambda)\lesssim\frac{\lambda^{2}(1-\lambda)^{2}~\mathrm{P}_{g}^{2}~\mathrm{P}_{r}^{4}~\mathrm{P}_{b}^{4}}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}\rho_{\lambda}^{4}}\cdot d\cdot\left(\frac{K_{r}^{4}}{n_{r}}+\frac{K_{b}^{4}}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2},$ (68)
with
$\rho_{\lambda}^{2}=\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}.$ (69)
To do so, first apply Proposition 3 under $\Sigma_{g}\preceq\mathrm{P}_{g}^{2}I_{d}$ to obtain the following bound:
$\mathcal{B}_{g}(\lambda)\leq\mathrm{P}_{g}^{2}\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})\right)\right\rVert^{2}\right].$ (70)
Next, observe that, be definition of $\beta_{\lambda}$,
$\lambda\Sigma_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\Sigma_{b}(\beta_{b}-\beta_{\lambda})=0.$
Subtracting this identity from its empirical counterpart therefore yields,
$\tilde{A}\coloneqq\lambda\widehat{\Sigma}_{r}(\beta_{r}-\beta_{\lambda})+(1-\lambda)\widehat{\Sigma}_{b}(\beta_{b}-\beta_{\lambda})=\lambda(\widehat{\Sigma}_{r}-\Sigma_{r})(\beta_{r}-\beta_{\lambda})+(1-\lambda)(\widehat{\Sigma}_{b}-\Sigma_{b})(\beta_{b}-\beta_{\lambda}).$ (71)
By triangle inequality,
$\left\lVert\tilde{A}\right\rVert\leq\lambda\left\lVert\widehat{\Sigma}_{r}-\Sigma_{r}\right\rVert\left\lVert\beta_{r}-\beta_{\lambda}\right\rVert+(1-\lambda)\left\lVert\widehat{\Sigma}_{b}-\Sigma_{b}\right\rVert\left\lVert\beta_{b}-\beta_{\lambda}\right\rVert.$ (72)
Next, notice that we can express
$\beta_{r}-\beta_{\lambda}$ $=(1-\lambda)\,\Sigma_{\lambda}^{-1}\Sigma_{b}(\beta_{r}-\beta_{b})$ (73a)
$\beta_{b}-\beta_{\lambda}$ $=\lambda\,\Sigma_{\lambda}^{-1}\Sigma_{r}(\beta_{b}-\beta_{r}).$ (73b)
Substituting (73a) and (73b) into (72) and applying the assumptions $\left\lVert\Sigma_{g}\right\rVert\leq\mathrm{P}_{g}^{2}$ and $\lambda_{\min}\left({\Sigma_{\lambda}}\right)\geq\rho_{\lambda}^{2}$, we reach
$\left\lVert\tilde{A}\right\rVert\leq\frac{\lambda(1-\lambda)}{\rho_{\lambda}^{2}}\left(\mathrm{P}_{b}^{2}\left\lVert\widehat{\Sigma}_{r}-\Sigma_{r}\right\rVert+\mathrm{P}_{r}^{2}\left\lVert\widehat{\Sigma}_{b}-\Sigma_{b}\right\rVert\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert.$ (74)
Raising (74) to the fourth power and using the inequality $(a+b)^{4}\leq 8(a^{4}+b^{4})$,
$\left\lVert\tilde{A}\right\rVert^{4}\leq\frac{8\lambda^{4}(1-\lambda)^{4}}{\rho_{\lambda}^{8}}\left(\mathrm{P}_{b}^{8}\left\lVert\widehat{\Sigma}_{r}-\Sigma_{r}\right\rVert^{4}+\mathrm{P}_{r}^{8}\left\lVert\widehat{\Sigma}_{b}-\Sigma_{b}\right\rVert^{4}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{4}.$
Taking an expectation and square root successively then using that $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b},$
$\sqrt{\mathbb{E}\left[\left\lVert\tilde{A}\right\rVert^{4}\right]}\leq\frac{2\sqrt{2}\lambda^{2}(1-\lambda)^{2}}{\rho_{\lambda}^{4}}\left(\mathrm{P}_{b}^{4}\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{r}-\Sigma_{r}\right\rVert^{4}\right]^{1/2}+\mathrm{P}_{r}^{4}\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{b}-\Sigma_{b}\right\rVert^{4}\right]^{1/2}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$
Applying the subgaussian covariance estimation bound Lemma 4 to each term $\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert^{4}\right]$, we obtain
$\sqrt{\mathbb{E}\left[\left\lVert\tilde{A}\right\rVert^{4}\right]}\leq\frac{2\sqrt{2}C\lambda^{2}(1-\lambda)^{2}}{\rho_{\lambda}^{4}}\mathrm{P}_{r}^{4}\mathrm{P}_{b}^{4}d\left(\frac{K_{r}^{4}}{n_{r}}+\frac{K_{b}^{4}}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$ (75)
Next, recall the property
$\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\tilde{A}\right\rVert^{2}\leq\lambda_{\min}\left(\widehat{\Sigma}_{\lambda}\right)^{-2}\left\lVert\tilde{A}\right\rVert^{2}.$ (76)
By Cauchy-Schwarz and (76),
$\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\tilde{A}\right\rVert^{2}\right]\leq\sqrt{\mathbb{E}\left[\lambda_{\min}\left(\widehat{\Sigma}_{\lambda}\right)^{-4}\right]}\sqrt{\mathbb{E}\left\lVert\tilde{A}\right\rVert^{4}}.$
By Weyl’s inequality, convexity, and Lemma 3 as in the preceding analysis of the variance term, we have
$\sqrt{\mathbb{E}\left[\lambda_{\min}\left(\widehat{\Sigma}_{\lambda}\right)^{-4}\right]}\lesssim\frac{1}{\left({\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}\right)^{2}}.$ (77)
Combining the bounds (75) and (77), we conclude the proof of the upper bound on the bias term.
Lower bound
Fix $\rho_{r}^{2},\rho_{b}^{2}>0$ and $\lambda\in(0,1)$. For each coordinate $i\in[d]$, set
$v$ $\coloneq\frac{\beta_{r}-\beta_{b}}{\left\|\beta_{r}-\beta_{b}\right\|_{2}},$ (78a)
[…]
Perturbed covariance structure for group $r$:
For a Rademacher vector $\xi\in\{-1,1\}^{d}$, set
$\Sigma_{r}^{(\xi)}\coloneqq\rho_{r}^{2}I_{d}+h_{r}\sum_{i=1}^{d}\xi_{i}u_{i}u_{i}^{\top},\quad\Sigma_{b}\coloneqq\rho_{b}^{2}I_{d}$
where the group $r$ perturbation level $h_{r}$ is given by
$h_{r}=\frac{2\rho_{r}^{2}}{5\sqrt{n_{r}}}\leq\frac{\rho_{r}^{2}}{10d},$ (79)
where the inequality follows from $n_{r}\geq 16d^{2}$. Hence, for every $\xi$,
$0.9\cdot\rho_{r}^{2}I_{d}\preceq\Sigma_{r}^{(\xi)}\preceq 1.1\rho_{r}^{2}I_{d}$ (80)
for $d\geq 1$.
KL diverence bound:
Define $\mathbb{P}^{(\xi)}\coloneqq\mathcal{N}(0,\Sigma_{r}^{(\xi)})$. The following result provides an upper bound on the KL divergence between the distributions under parameters $\xi,\xi^{\prime}\in\Xi^{d}$ differing in only a single coordinate.
Claim 2.
Let $\xi$ and $\xi^{\prime}$ differ only in coordinate $i$ where $\xi_{i}=-\xi_{i}^{\prime}$. Then
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})\leq\frac{25}{16}\frac{h_{r}^{2}}{\rho_{r}^{4}}.$
Proof of 2.
Observe that
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})=\frac{1}{2}\left[\log\frac{\det\left(\Sigma_{r}-h_{r}u_{i}u_{i}^{\top}\right)}{\det\left(\Sigma_{r}+h_{r}u_{i}u_{i}^{\top}\right)}-d+\operatorname{Tr}\left(\left(\Sigma_{r}-h_{r}u_{i}u_{i}^{\top}\right)^{-1}\right)\left(\Sigma_{r}+h_{r}u_{i}u_{i}^{\top}\right)\right]$
Define $\alpha=h_{r}u_{i}^{\top}\Sigma_{r}^{-1}u_{i}$. By the matrix determinant lemma,
$\det\left(\Sigma_{r}-h_{r}u_{i}u_{i}^{\top}\right)=\left(1-\alpha\right)\det(\Sigma_{r}).$
By the Sherman-Morrison lemma, $\left(\Sigma_{r}-h_{r}u_{i}u_{i}^{\top}\right)^{-1}=\Sigma_{r}^{-1}\frac{h_{r}}{1-\alpha}(\Sigma_{r}^{-1}u_{i}u_{i}^{\top}\Sigma_{r}^{-1})$. Substituting and simplifying, we reach
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})=\frac{1}{2}\left[\log\frac{\left(1-\alpha\right)}{\left(1+\alpha\right)}+\frac{2\alpha}{1-\alpha}\right].$
Observe $\alpha\geq 0$ since $\Sigma_{r}$ is PSD. By (80), $\Sigma_{r}^{-1}\preceq\frac{1}{0.9\rho_{r}^{2}}I_{d}$ thus $\alpha\leq\frac{h_{r}}{0.9\rho_{r}^{2}}$. By (79), it follows that $0\leq\alpha\leq\frac{1}{9d}\leq\frac{1}{9}$ for $d\geq 1$. Define $\gamma\coloneqq\frac{2\alpha}{1-\alpha}$ where $\gamma\in[0,1/4]$. Using the fact that, for $\gamma\in[0,1/4]$, $\log(1+\gamma)+\gamma\leq\frac{\gamma^{2}}{2}$ and the fact that, for $\alpha\in(0,1/9)$, $\frac{\alpha^{2}}{(1-\alpha)^{2}}\leq\frac{81}{64}\alpha^{2}$,
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})=\frac{1}{2}\left[-\log(1+\gamma)+\gamma\right]\leq\frac{\gamma^{2}}{4}=\frac{\alpha^{2}}{(1-\alpha)^{2}}\leq\frac{81}{64}\alpha^{2}\leq\frac{81}{64}h_{r}^{2}\left\lVert\Sigma_{r}^{-1}\right\rVert_{2}^{2}\leq\frac{25}{16}\frac{h_{r}^{2}}{\rho_{r}^{4}}$
where the last inequality follows by (80). ∎
By 2, over $n_{r}$ independent samples from group $r$,
$D_{\text{KL}}(\mathbb{P}^{(\xi)\otimes n_{r}}\;\|\;\mathbb{P}^{(\xi^{\prime})\otimes n_{r}})\leq\frac{25}{16}\frac{h_{r}^{2}n_{r}}{\rho_{r}^{4}}.$
By (79) and an application of Pinsker’s inequality, we reach the bound
$\left\lVert\mathbb{P}^{(\xi)\otimes n_{r}},\mathbb{P}^{(\xi^{\prime})\otimes n_{r}}\right\rVert_{\text{TV}}\leq\frac{1}{2}.$ (81)
Parameter separation:
Define ${A^{(\xi)}}\coloneqq\left(\lambda\Sigma_{r}^{(\xi)}+(1-\lambda)\Sigma_{b}\right)$. The target parameter under $\xi$ can be expressed as
$\beta_{\lambda}^{(\xi)}=\left({A^{(\xi)}}\right)^{-1}\left(\lambda\Sigma_{r}^{(\xi)}\beta_{r}+(1-\lambda)\Sigma_{b}\beta_{b}\right)=\beta_{r}+\left({A^{(\xi)}}\right)^{-1}(1-\lambda)\Sigma_{b}(\beta_{b}-\beta_{r}).$
Claim 3.
Let $\xi$ and $\xi^{\prime}$ differ only in coordinate $i$ where $\xi_{i}=-\xi_{i}^{\prime}$. Then,
$\left|\left\langle e_{i},\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}\right\rangle\right|\gtrsim\lambda(1-\lambda)h_{r}\rho_{b}^{2}\rho_{\lambda}^{-4}\left\lVert\beta_{r}-\beta_{b}\right\rVert.$
Proof of 3.
By the Sherman-Morrison lemma,
${A^{(\xi)}}^{-1}-{A^{(\xi^{\prime})}}^{-1}=\frac{2\lambda h_{r}}{1+2\lambda h_{r}u_{i}^{\top}{A^{(\xi)}}^{-1}u_{i}}{A^{(\xi)}}^{-1}u_{i}u_{i}^{\top}{A^{(\xi)}}^{-1}.$
Hence, we can express
$\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}=\left({A^{(\xi)}}^{-1}-{A^{(\xi^{\prime})}}^{-1}\right)(1-\lambda)\Sigma_{b}(\beta_{b}-\beta_{r}).$
Then, substitution of $v$ defined in (78a),
$\langle e_{i},\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}\rangle=2\lambda(1-\lambda)h_{r}\left\lVert\Sigma_{b}\right\rVert\left\lVert\beta_{r}-\beta_{b}\right\rVert\frac{e_{i}^{\top}{A^{(\xi)}}^{-1}u_{i}u_{i}^{\top}{A^{(\xi)}}^{-1}v}{1+2\lambda h_{r}u_{i}^{\top}{A^{(\xi)}}^{-1}u_{i}}.$
By the Loewner order relationship (80),
$\left\lVert{A^{(\xi)}}\right\rVert\geq 0.9\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\geq 0.9\lambda\rho_{r}^{2}$
The denominator is bounded, therefore, by:
$|1+2\lambda h_{r}u_{i}^{\top}{A^{(\xi)}}^{-1}u_{i}|\leq\left|1+2\lambda h_{r}\left\lVert{A^{(\xi)}}^{-1}\right\rVert\right|\leq\left|1+\frac{2\lambda h_{r}}{\left(0.9\lambda\rho_{r}^{2}\right)}\right|\leq\frac{11}{9}$
where the last inequality follows by assuming $h_{r}\leq\frac{\rho_{r}^{2}}{10d}$ and $d\geq 1$.
Next, we lower bound the numerator. Fix a coordinate $i\in[d]$, and observe that we can write the decomposition
${A^{(\xi)}}^{-1}e_{i}=c_{1}e_{i}+c_{2}w_{i}$
where $w_{i}\in\mathbb{R}^{d}$ is orthogonal to $e_{i}$ and satisfies $\left\lVert w_{i}\right\rVert=1$. By definition, $c_{1}=e_{i}^{\top}{A^{(\xi)}}^{-1}e_{i}$, and, using (80),
$\frac{1}{1.1}\rho_{\lambda}^{-2}\leq\left(1.1\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{-1}\leq c_{1}\leq\left(0.9\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{-1}\leq\frac{1}{0.9}\rho_{\lambda}^{-2}.$ (82)
Moreover, by construction and again by (80),
$c_{1}^{2}+c_{2}^{2}=\left\lVert{A^{(\xi)}}^{-1}e_{i}\right\rVert^{2}\leq\left\lVert{A^{(\xi)}}^{-1}\right\rVert^{2}\leq\left(0.9\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{-2}\leq\frac{1}{0.9^{2}}\rho_{\lambda}^{-4}.$ (83)
Combining (82) and (83), we have
$c_{2}\leq\left(0.9^{-2}-1.1^{-2}\right)^{0.5}\rho_{\lambda}^{-2}.$ (84)
Next, noting that the vectors $u_{i}$ and $v$ satisfy $|u_{i}^{\top}e_{i}|,|u_{i}^{\top}v|\geq\frac{1}{\sqrt{2}}$, we have
$|e_{i}^{\top}{A^{(\xi)}}^{-1}u_{i}|=|c_{1}e_{i}^{\top}u_{i}+c_{2}w_{i}^{\top}u_{i}|\geq\frac{c_{1}}{\sqrt{2}}-|c_{2}|,\quad|u_{i}^{\top}{A^{(\xi)}}^{-1}v|\geq\frac{c_{1}}{\sqrt{2}}-|c_{2}|.$
Combining (82) and (84), we see that
$\frac{c_{1}}{\sqrt{2}}-|c_{2}|\geq\frac{1}{300}\rho_{\lambda}^{-2}.$
All together, this yields the parameter separation lower bound:
$\langle e_{i},\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}\rangle^{2}\geq\left(\frac{18}{11\cdot 300^{2}}\right)^{2}\lambda^{2}(1-\lambda)^{2}h_{r}^{2}\rho_{b}^{4}\rho_{\lambda}^{-8}\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$
Assouad’s Lemma:
Let $\widehat{\beta}$ be any estimator. Assouad’s lemma (see Lemma 6) applied to the $d$-dimensional hypercube $\Xi_{d}$, given the results 3 and (81), yields
$\sup_{\xi\in\Xi^{d}}\mathbb{E}_{\xi}\left[\left\lVert\widehat{\beta}-\beta\right\rVert_{2}^{2}\right]\geq\left(\frac{18}{11\cdot 300^{2}}\right)^{2}\frac{2}{25}\lambda^{2}(1-\lambda)^{2}\frac{\rho_{r}^{4}\cdot\rho_{b}^{4}\cdot d}{\rho_{\lambda}^{8}\cdot n_{r}}\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$ (85)
Symmetric perturbation for group $b$:
Repeating the procedure so far but instead perturbing the covariance structure of group $b$, i.e., taking, for a Rademacher vector $\zeta\in\{-1,1\}^{d}$, set
$\Sigma_{r}\coloneqq\rho_{r}^{2}I_{d},\quad\Sigma_{b}^{(\zeta)}\coloneqq\rho_{b}^{2}I_{d}+h_{b}\sum_{i=1}^{d}\xi_{i}u_{i}u_{i}^{\top},\quad h_{b}=\frac{2\rho_{b}^{2}}{5\sqrt{n_{b}}}.$
This construction yields the analogous lower bound
$\sup_{\zeta\in\Xi^{d}}\mathbb{E}_{\xi}\left[\left\lVert\widehat{\beta}-\beta\right\rVert_{2}^{2}\right]\geq\left(\frac{18}{11\cdot 300^{2}}\right)^{2}\frac{2}{25}\lambda^{2}(1-\lambda)^{2}\frac{\rho_{r}^{4}\cdot\rho_{b}^{4}\cdot d}{\rho_{\lambda}^{8}\cdot n_{b}}\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.$ (86)
Since we may perturb either group, we may take the maximal lower bound and use the fact that $\max(a,b)\geq\frac{1}{2}(a+b)$ to obtain the bound:
$\sup_{\xi,\zeta\in\Xi^{d}}\left[\left\lVert\widehat{\beta}-\beta\right\rVert_{2}^{2}\right]\geq C\lambda^{2}(1-\lambda)^{2}\frac{\rho_{r}^{4}\cdot\rho_{b}^{4}\cdot d}{\rho_{\lambda}^{8}}\left(\frac{1}{n_{r}}+\frac{1}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}$
where $C=\left(\frac{18}{11\cdot 300^{2}}\right)^{2}\frac{1}{25}$.
Finally, using that $\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}\geq\rho_{g}^{2}\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{2}^{2}$, we reach the bound
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\geq\lambda^{2}(1-\lambda)^{2}\frac{\rho_{g}^{2}\cdot\rho_{r}^{4}\cdot\rho_{b}^{4}\cdot d}{\rho_{\lambda}^{8}}\left(\frac{1}{n_{r}}+\frac{1}{n_{b}}\right)\left\lVert\beta_{r}-\beta_{b}\right\rVert^{2}.\quad\blacksquare$
A.4.4 Proof of Proposition 4
First, note that, using (73) from the proof of Theorem 3, we have
$\left|\mathbb{E}\left[(\widehat{\beta}_{\lambda}-\beta_{\lambda})^{\top}\Sigma_{r}(\beta_{\lambda}-\beta_{r})\right]\right|\lesssim\frac{(1-\lambda)\rho_{r}^{2}\rho_{b}^{2}}{\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}}~\|\beta_{r}-\beta_{b}\|~\left\|\mathbb{E}\left[\widehat{\beta}_{\lambda}-\beta_{\lambda}\right]\right\|.$ (87)
Next, we recall from the proof of Lemma 2 that
$\mathbb{E}\left[\widehat{\beta}_{\lambda}-\beta_{\lambda}\right]=\mathbb{E}\left[\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\widehat{\Sigma}_{r}\beta_{r}+(1-\lambda)\widehat{\Sigma}_{b}\beta_{b}\right)-\beta_{\lambda}\right],$ (88)
which, as described in the proof of Proposition 3, can be further cast as
$\mathbb{E}\left[\widehat{\beta}_{\lambda}-\beta_{\lambda}\right]=\mathbb{E}\left[\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})+(1-\lambda)\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda})\right)\right].$ (89)
By the Woodbury matrix identity, we can write
$\widehat{\Sigma}_{\lambda}^{-1}=\Sigma_{\lambda}^{-1}-\Sigma_{\lambda}^{-1}(\widehat{\Sigma}_{\lambda}-\Sigma_{\lambda})\widehat{\Sigma}_{\lambda}^{-1}.$ (90)
By substituting the above identity into (89), and using the fact that $\mathbb{E}[\widehat{\Sigma}_{g}-\Sigma_{g}]=0$ for both groups, we obtain
$\mathbb{E}\left[\widehat{\beta}_{\lambda}-\beta_{\lambda}\right]=-\mathbb{E}\left[\Sigma_{\lambda}^{-1}(\widehat{\Sigma}_{\lambda}-\Sigma_{\lambda})\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})+(1-\lambda)\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda})\right)\right].$ (91)
Next, we bound the two terms on the right-hand side separately. First, notice that
$\left\|\mathbb{E}\left[\Sigma_{\lambda}^{-1}(\widehat{\Sigma}_{\lambda}-\Sigma_{\lambda})\widehat{\Sigma}_{\lambda}^{-1}\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})\right]\right\|$
$~~\lesssim\frac{(1-\lambda)\rho_{b}^{2}}{\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{2}}~\|\beta_{r}-\beta_{b}\|~\left\|\mathbb{E}\left[(\widehat{\Sigma}_{\lambda}-\Sigma_{\lambda})\widehat{\Sigma}_{\lambda}^{-1}\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)\right]\right\|$
$~~\lesssim\frac{(1-\lambda)\rho_{b}^{2}}{\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{2}}~\|\beta_{r}-\beta_{b}\|~\sqrt{\mathbb{E}\left[\left\|\widehat{\Sigma}_{\lambda}^{-1}\right\|^{2}\right]}\left(\mathbb{E}\left[\left\|\Sigma_{r}-\widehat{\Sigma}_{r}\right\|^{4}\right]\right)^{1/4}\left(\mathbb{E}\left[\left\|\Sigma_{\lambda}-\widehat{\Sigma}_{\lambda}\right\|^{4}\right]\right)^{1/4}$
$~~\lesssim\frac{(1-\lambda)\rho_{b}^{2}}{\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{2}}~\|\beta_{r}-\beta_{b}\|~\frac{K_{r}^{2}\rho_{r}^{2}{\sqrt{d}}/{\sqrt{n_{r}}}}{{\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}}\left(\lambda K_{r}^{2}\rho_{r}^{2}\frac{\sqrt{d}}{\sqrt{n_{r}}}+(1-\lambda)K_{b}^{2}\rho_{b}^{2}\frac{\sqrt{d}}{\sqrt{n_{b}}}\right),$ (92)
where the last inequality follows from Lemma 3 and Lemma 4. Similarly, we can show
$\left\|\mathbb{E}\left[\Sigma_{\lambda}^{-1}(\widehat{\Sigma}_{\lambda}-\Sigma_{\lambda})\widehat{\Sigma}_{\lambda}^{-1}\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda})\right]\right\|$
$~~\lesssim\frac{\lambda\rho_{r}^{2}}{\left(\lambda\rho_{r}^{2}+(1-\lambda)\rho_{b}^{2}\right)^{2}}~\|\beta_{r}-\beta_{b}\|~\frac{K_{b}^{2}\rho_{b}^{2}{\sqrt{d}}/{\sqrt{n_{b}}}}{{\lambda}\rho_{r}^{2}/{{C}_{r}^{\prime}}+(1-\lambda)\rho_{b}^{2}/{{C}_{b}^{\prime}}}\left(\lambda K_{r}^{2}\rho_{r}^{2}\frac{\sqrt{d}}{\sqrt{n_{r}}}+(1-\lambda)K_{b}^{2}\rho_{b}^{2}\frac{\sqrt{d}}{\sqrt{n_{b}}}\right).$ (93)
Plugging (A.4.4) and (A.4.4) into (91), and then substituting the whole term into (87) completes the proof. $\blacksquare$.
A.5 Proof of Theorem 4
We first provide a formal statement of Theorem 4 showing the full dependence on the parameters in our setting.
Theorem 4’. Suppose Assumptions 1-5 hold. Fix $\alpha>2$ and define the truncation thresholds
$T_{X}\coloneq C_{X}\max\{K_{r},K_{b}\}\max\{\rho_{r},\rho_{b}\}\sqrt{d+\alpha\log(n)},\quad T_{Y}\coloneq C_{Y}\left(BT_{X}+\sigma_{\varepsilon}\sqrt{\alpha\log(n)}\right)$
where $C_{X},C_{Y}>0$ are constants. Also define the polynomial
$K_{\Delta}\coloneq\sqrt{\frac{1}{n_{r}}+\frac{1}{n_{b}}}\left(12\kappa T_{X}^{2}T_{Y}^{2}+56\kappa^{2}T_{X}^{4}T_{Y}^{2}+76\kappa^{3}T_{X}^{6}T_{Y}^{2}+36\kappa^{4}T_{X}^{8}T_{Y}^{2}\right)$
where $\kappa\coloneq 2\min\{\lambda_{\min}(\Sigma_{r}),\lambda_{\min}(\Sigma_{b})\}^{-1}$ and, for each $g\in\mathcal{G}$, set
$\Gamma_{g}\coloneq 4B^{2}CK_{g}^{2}\left\lVert\Sigma_{g}\right\rVert\left(\sqrt{\frac{d}{n_{g}}}+\frac{d}{n_{g}}\right)+4BCK_{g}\sigma_{\varepsilon}\left\lVert\Sigma_{g}^{1/2}\right\rVert\sqrt{\frac{d}{n_{g}}}+C\sigma_{\varepsilon}^{2}\frac{1}{\sqrt{n_{g}}}$
where $C>0$ is an absolute constant. Fix any $\lambda_{0}\in[0,1]$ and fix
$\Gamma_{\lambda_{0}}\leq C\sum_{g\in\{r,b\}}\left(\kappa_{t}^{2}\left(T_{X}^{2}T_{Y}^{2}+B^{2}T_{X}^{4}\right)\left\lVert\Sigma_{g}^{1/2}\right\rVert^{2}\frac{d}{n_{g}}+B^{2}\left\lVert\Sigma_{g}^{1/2}\right\rVert^{2}(T_{X}T_{Y}+BT_{X}^{2})\sqrt{\frac{d}{n_{g}}}\right)+(\Gamma_{r}+\Gamma_{b})$
Then, for every $\delta\in(0,1)$, there exists a constant $C_{\Delta}>0$ such that, with probability at least $1-\delta-\eta$,
$\sup_{\lambda\in[0,1]}\left\lVert\Delta_{\lambda}\right\rVert_{2}\leq\Gamma_{\lambda_{0}}+2(\Gamma_{r}+\Gamma_{b})+C_{\Delta}K_{\Delta}\left(1.26+\log(100/\delta)\right)$
where
$\eta(\alpha,\alpha_{r},\alpha_{b})\coloneq n^{-C\alpha}+\left(C_{r}^{\prime}/2\right)^{\alpha_{r}n_{r}/6}+\left(C_{b}^{\prime}/2\right)^{\alpha_{b}n_{b}/6}.$
To show this claim, first fix $t\in(0,1)$ and set ${\kappa_{t}}\coloneq\left(t\cdot\min\{\lambda_{\min}(\Sigma_{r}),\lambda_{\min}(\Sigma_{b})\}\right)^{-1}$. We will show that,
$\mathbb{P}\left(\sup_{\lambda,\lambda^{\prime}\in[0,1]}\left\lVert\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\right\rVert_{2}\leq C_{\Delta}K_{\Delta}\left(1.26+\sqrt{\log\left(50/\delta\right)}\right)+2(\Gamma_{r}+\Gamma_{b})\right)\geq 1-\delta-{\eta}_{t}(\alpha,\alpha_{r},\alpha_{b})$ (94)
for every $\delta\in(0,1)$ where ${\eta}_{t}(\alpha,\alpha_{r},\alpha_{b})\coloneq n^{-C\alpha}+\left(C_{r}^{\prime}t\right)^{\alpha_{r}n_{r}/6}+\left(C_{b}^{\prime}t\right)^{\alpha_{b}n_{b}/6}$ and all other parameters are defined as in the statement of  Theorem 4. Choosing $t=1/2$ gives the theorem with $\kappa\coloneq\kappa_{1/2}$ and $\eta\coloneq\eta_{1/2}$.
A.5.1 Decomposition:
For any $\lambda,\lambda^{\prime}\in[0,1]$ and $u\in\mathbb{R}^{2}$, write
$\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle=S_{\lambda,\lambda^{\prime},u}(\mathcal{S})-D_{\lambda,\lambda^{\prime},u}$
with
$S_{\lambda,\lambda^{\prime},u}(\mathcal{S})\coloneq\sum_{g\in\mathcal{G}}u_{g}\left(\mathcal{R}_{g}(P_{n},\widehat{\beta}_{\lambda})-\mathcal{R}_{g}(P_{n},\widehat{\beta}_{\lambda^{\prime}})\right),\quad D_{\lambda,\lambda^{\prime},u}\coloneq\sum_{g\in\mathcal{G}}u_{g}\left(\mathcal{R}_{g}(P,\beta_{\lambda})-\mathcal{R}_{g}(P,\beta_{\lambda^{\prime}})\right).$
Since $D_{\lambda,\lambda^{\prime},u}$ is deterministic, it suffices to control $S_{\lambda,\lambda^{\prime},u}(\mathcal{S})$.
A.5.2 Risk gap bound
For each $g\in\mathcal{G}$, writing $\widehat{\beta}_{\lambda}$ = $\widehat{\beta}_{\lambda^{\prime}}+(\widehat{\beta}_{\lambda}-\widehat{\beta}_{\lambda^{\prime}})$, expanding the quadratic, and applying Cauchy-Schwarz gives
$\left|\mathcal{R}_{g}(\widehat{\beta}_{\lambda})-\mathcal{R}_{g}(\widehat{\beta}_{\lambda^{\prime}})\right|$ $\leq\left\lVert\widehat{\beta}_{\lambda}-\widehat{\beta}_{\lambda^{\prime}}\right\rVert_{2}^{2}V_{g}+\left\lVert\widehat{\beta}_{\lambda}-\widehat{\beta}_{\lambda^{\prime}}\right\rVert_{2}W_{g}+2\left\lVert\widehat{\beta}_{\lambda}-\widehat{\beta}_{\lambda^{\prime}}\right\rVert_{2}\left\lVert\widehat{\beta}_{\lambda^{\prime}}\right\rVert_{2}V_{g}.$ (95)
where
$V_{g}\coloneq\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}\left\lVert{X_{i}^{(g)}}\right\rVert_{2}^{2},\quad W_{g}\coloneq\frac{2}{n_{g}}\sum_{i=1}^{n_{g}}\lvert Y_{i}^{(g)}\rvert\left\lVert{X_{i}^{(g)}}\right\rVert_{2}.$ (96)
A.5.3 Truncation and empirical control
Fix $\alpha>2$ and absolute constants and define the truncation thresholds
$T_{X}\coloneq C_{X}\max\{K_{r},K_{b}\}\max\{\rho_{r},\rho_{b}\}\sqrt{d+\alpha\log(n)},\quad T_{Y}\coloneq C_{Y}\left(BT_{X}+\sigma_{\varepsilon}\sqrt{\alpha\log(n)}\right)$
where $B>0$ satisfies $\max_{g}\beta_{g}\leq B$ and $C_{X},C_{Y}$ are constants. Also define the event
$\mathcal{E}\coloneq\left\{\max_{i,g}\left\lVert X_{i}^{(g)}\right\rVert_{2}\leq T_{X},\max_{i,g}\left\lVert Y_{i}^{(g)}\right\rVert_{2}\leq T_{Y}\right\}.$
Under Assumptions 4, 3, and 5, by a standard $1/2$-net argument (e.g., vershynin2018high, vershynin2018high),there exists a constant $c>0$ such that
$\mathbb{P}\left(\mathcal{E}\right)\geq 1-n^{-c\alpha}.$ (97)
A.5.4 Empirical covariance control
Fix $t\in(0,1)$. Define the event
$\mathcal{A}(t)\coloneq\left\{\widehat{\Sigma}_{r}\succeq t\Sigma_{r},\widehat{\Sigma}_{b}\succeq t\Sigma_{b}\right\}.$
By 3, mourtada2022exact and the union bound,
$\mathbb{P}(\mathcal{A}(t))\geq 1-\left(C_{r}^{\prime}t\right)^{\alpha_{r}n_{r}/6}-\left(C_{b}^{\prime}t\right)^{\alpha_{b}n_{b}/6}.$ (98)
On $\mathcal{A}(t)$, $\widehat{\Sigma}_{\lambda}\succeq\lambda t\Sigma_{r}+(1-\lambda)t\Sigma_{b}\succeq t\Sigma_{\lambda}$, hence, on the same event,
$\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\right\rVert\leq\kappa_{t}\coloneq\left(t\cdot\min\{\lambda_{\min}(\Sigma_{r}),\lambda_{\min}(\Sigma_{b})\}\right)^{-1}.$ (99)
A.5.5 Lipschitz control for $\widehat{\beta}_{\lambda}$:
Recall $\widehat{\beta}_{\lambda}=\widehat{\Sigma}_{\lambda}^{-1}\widehat{\nu}_{\lambda}$. For any $\lambda,\lambda^{\prime}\in[0,1]$,
$\widehat{\beta}_{\lambda}-\widehat{\beta}_{\lambda^{\prime}}=\widehat{\Sigma}_{\lambda}^{-1}\left(\widehat{\nu}_{\lambda}-\widehat{\nu}_{\lambda^{\prime}}\right)+\left(\widehat{\Sigma}_{\lambda}^{-1}-\widehat{\Sigma}_{\lambda^{\prime}}^{-1}\right)\widehat{\nu}_{\lambda^{\prime}}$
For the first term, noting that $\widehat{\nu}_{\lambda}-\widehat{\nu}_{\lambda^{\prime}}=\left(\lambda-\lambda^{\prime}\right)(\widehat{\nu}_{r}-\widehat{\nu}_{b})$, we can bound
$\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\left(\widehat{\nu}_{\lambda}-\widehat{\nu}_{\lambda^{\prime}}\right)\right\rVert_{2}\leq\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\right\rVert\left\lVert\widehat{\nu}_{r}-\widehat{\nu}_{b}\right\rVert_{2}|\lambda-\lambda^{\prime}|.$ (100)
For the second term, using the identities $\widehat{\Sigma}_{\lambda}^{-1}-\widehat{\Sigma}_{\lambda^{\prime}}^{-1}=\widehat{\Sigma}_{\lambda}^{-1}\left(\widehat{\Sigma}_{\lambda}-\widehat{\Sigma}_{\lambda^{\prime}}\right)\widehat{\Sigma}_{\lambda^{\prime}}^{-1}$ and $\widehat{\Sigma}_{\lambda}-\widehat{\Sigma}_{\lambda^{\prime}}=\left(\lambda-\lambda^{\prime}\right)(\widehat{\Sigma}_{r}-\widehat{\Sigma}_{b})$, we have the bound
$\left\lVert\left(\widehat{\Sigma}_{\lambda}^{-1}-\widehat{\Sigma}_{\lambda^{\prime}}^{-1}\right)\widehat{\nu}_{\lambda^{\prime}}\right\rVert_{2}\leq\left\lVert\widehat{\Sigma}_{\lambda}^{-1}\right\rVert\left\lVert\widehat{\Sigma}_{\lambda^{\prime}}^{-1}\right\rVert\left\lVert\widehat{\Sigma}_{r}-\widehat{\Sigma}_{b}\right\rVert\left\lVert\widehat{\nu}_{\lambda^{\prime}}\right\rVert|\lambda-\lambda^{\prime}|.$ (101)
Combining (100) and (101) with (99), on $\mathcal{A}(t)$,
$\left\lVert\widehat{\beta}_{\lambda}-\widehat{\beta}_{\lambda^{\prime}}\right\rVert\leq\kappa^{*}|\lambda-\lambda^{\prime}|\quad\text{with}\quad\kappa^{*}\coloneq\kappa_{t}\left\lVert\widehat{\nu}_{r}-\widehat{\nu}_{b}\right\rVert_{2}+\kappa_{t}^{2}\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}\right\rVert\left\lVert\widehat{\Sigma}_{r}-\widehat{\Sigma}_{b}\right\rVert.$ (102)
A.5.6 Truncation event bounds:
On $\mathcal{E}$, for each $g\in\mathcal{G}$,
$V_{g}\leq{T_{X}^{2}},\quad W_{g}\leq 2T_{X}T_{Y},\quad\left\lVert\widehat{\Sigma}_{r}-\widehat{\Sigma}_{b}\right\rVert\leq 2T_{X}^{2},\quad\left\lVert\widehat{\nu}_{r}-\widehat{\nu}_{b}\right\rVert\leq 2T_{X}T_{Y},\quad\left\lVert\widehat{\nu}_{g}\right\rVert\leq T_{X}T_{Y}.$ (103)
By (102), on $\mathcal{A}(t)\cap\mathcal{E}$,
$\kappa^{\star}\leq 2\kappa_{t}T_{X}T_{Y}+2\kappa_{t}^{2}T_{X}^{3}T_{Y},\quad\left\lVert\widehat{\beta}_{\lambda}(\mathcal{S})\right\rVert\leq\kappa_{t}T_{X}T_{Y}.$ (104)
A.5.7 Lipschitzness to risk bound:
Plugging (102) into (95), we obtain that, on $\mathcal{A}(t)$,
$\left|S_{\lambda,\lambda^{\prime},u}(\mathcal{S})\right|\leq\left\lVert u\right\rVert_{2}L(\mathcal{S})|\lambda-\lambda^{\prime}|,\quad L(\mathcal{S})\coloneq\sqrt{L_{r}^{2}(\mathcal{S})+L_{b}^{2}(\mathcal{S})}$
with $L_{g}(\mathcal{S})\coloneq{\kappa^{*}}^{2}V_{g}(\mathcal{S})+\kappa^{*}W_{g}(\mathcal{S})+2\kappa_{t}\kappa^{*}\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}\right\rVert_{2}V_{g}(\mathcal{S}).$
A.5.8 One sample perturbation bounds:
Fix ${g^{*}}\in\mathcal{G}$, $j\in[n_{g^{*}}]$, and let $\mathcal{S}_{g^{*}}^{(j)}$ replace $(X_{j}^{({g^{*}})},Y_{j}^{({g^{*}})})$ by an i.i.d. copy $({X_{j}^{({g^{*}})}}^{\prime},{Y_{j}^{({g^{*}})}}^{\prime})$. With $w_{r}(\lambda)\coloneq\lambda,w_{b}(\lambda)\coloneq 1-\lambda$, by submultiplicativity,
$\left\lVert\widehat{\Sigma}_{\lambda}(\mathcal{S})-\widehat{\Sigma}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\frac{w_{g^{*}}(\lambda)}{n_{g^{*}}}\left(\left\lVert{X_{j}^{({g^{*}})}}^{\prime}\right\rVert_{2}^{2}+\left\lVert{X_{j}^{({g^{*}})}}\right\rVert_{2}^{2}\right),$ (105)
$\left\lVert\widehat{\nu}_{\lambda}(\mathcal{S})-\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\frac{w_{g^{*}}(\lambda)}{n_{g^{*}}}\left(\left\lVert{X_{j}^{({g^{*}})}}^{\prime}\right\rVert_{2}\Big|{Y_{j}^{({g^{*}})}}^{\prime}\Big|+\left\lVert{X_{j}^{({g^{*}})}}\right\rVert_{2}\Big|{Y_{j}^{({g^{*}})}}\Big|\right).$ (106)
On $\mathcal{A}(t)\cap\mathcal{E}$, by (105), (106), and (104),
$\left\lVert\widehat{\Sigma}_{\lambda}(\mathcal{S})-\widehat{\Sigma}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\frac{2}{n_{g^{*}}}T_{X}^{2},\quad\left\lVert\widehat{\nu}_{\lambda}(\mathcal{S})-\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\frac{2}{n_{g^{*}}}T_{X}T_{Y}.$ (107)
Then, using $\widehat{\beta}_{\lambda}=\widehat{\Sigma}_{\lambda}^{-1}\widehat{\nu}_{\lambda}$ and applying the identity $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$ to the following equation,
$\widehat{\beta}_{\lambda}(\mathcal{S})-\widehat{\beta}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})=\widehat{\Sigma}_{\lambda}^{-1}(\mathcal{S})\left(\widehat{\nu}_{\lambda}(\mathcal{S})-\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right)+\left(\widehat{\Sigma}_{\lambda}^{-1}(\mathcal{S}_{g^{*}}^{(j)})-\widehat{\Sigma}_{\lambda}^{-1}(\mathcal{S})\right)\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)}),$
on $\mathcal{A}(t)$,
$\left\lVert\left(\widehat{\Sigma}_{\lambda}^{-1}(\mathcal{S}_{g^{*}}^{(j)})-\widehat{\Sigma}_{\lambda}^{-1}(\mathcal{S})\right)\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\kappa_{t}^{2}\left\lVert\widehat{\Sigma}_{\lambda}(\mathcal{S})-\widehat{\Sigma}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\sup_{\lambda}\left\lVert\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert.$
Similarly, on $\mathcal{A}(t)$,
$\left\lVert\widehat{\Sigma}_{\lambda}^{-1}(\mathcal{S})\left(\widehat{\nu}_{\lambda}(\mathcal{S})-\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right)\right\rVert\leq\kappa_{t}\left\lVert\widehat{\nu}_{\lambda}(\mathcal{S})-\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert.$
On the event $\mathcal{A}(t)$,
$\left\lVert\widehat{\beta}_{\lambda}(\mathcal{S})-\widehat{\beta}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\kappa_{t}\left\lVert\widehat{\nu}_{\lambda}(\mathcal{S})-\widehat{\nu}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert+\kappa_{t}^{2}\left\lVert\widehat{\Sigma}_{\lambda}(\mathcal{S})-\widehat{\Sigma}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert.$
By (105) and (106),
$\left\lVert\widehat{\beta}_{\lambda}(\mathcal{S})-\widehat{\beta}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert$ $\leq\frac{w_{g^{*}}(\lambda)}{n_{g^{*}}}\left(\kappa_{t}\left(\left\lVert{X_{j}^{({g^{*}})}}^{\prime}\right\rVert_{2}\Big|{Y_{j}^{({g^{*}})}}^{\prime}\Big|+\left\lVert{X_{j}^{(g)}}\right\rVert_{2}\Big|{Y_{j}^{({g^{*}})}}\Big|\right)+\right$. (108)
$\left.\qquad\kappa_{t}^{2}\left(\left\lVert{X_{j}^{({g^{*}})}}^{\prime}\right\rVert_{2}^{2}+\left\lVert{X_{j}^{({g^{*}})}}\right\rVert_{2}^{2}\right)\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\right).$
On $\mathcal{A}(t)\cap\mathcal{E}$, by (103),
$\left\lVert\widehat{\beta}_{\lambda}(\mathcal{S})-\widehat{\beta}_{\lambda}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert\leq\frac{1}{n_{g^{*}}}\left(2\kappa_{t}T_{X}T_{Y}+2\kappa_{t}^{2}T_{X}^{3}T_{Y}\right).$ (109)
By the triangle inequality and the inequality $2ab\leq a^{2}+b^{2}$,
$\left|V_{g}(\mathcal{S})-V_{g}(\mathcal{S}_{g^{*}}^{(j)})\right|\leq\frac{1}{n_{g}}\left(\left\lVert{X_{j}^{(g)}}^{\prime}\right\rVert_{2}^{2}+\left\lVert{X_{j}^{(g)}}\right\rVert_{2}^{2}\right)$
$\left|W_{g}(\mathcal{S})-W_{g}(\mathcal{S}_{g^{*}}^{(j)})\right|\leq\frac{1}{n_{g}}\left(\left\lVert{X_{j}^{(g)}}\right\rVert^{2}_{2}+|{Y_{j}^{(g)}}|^{2}+\left\lVert{(X_{j}^{(g)})^{\prime}}\right\rVert^{2}_{2}+|{(Y_{j}^{(g)})^{\prime}}|^{2}\right)$
Next, we control $\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S})\right\rVert$ under perturbations. Using that for two real-valued functions $f,g$, $\sup(f-g)\geq\sup(f)-\sup(g)$,
$\left|\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S})\right\rVert_{2}-\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert_{2}\right|$ $\leq\max\left\{\left\lVert\widehat{\nu}_{r}(\mathcal{S})-\widehat{\nu}_{r}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert_{2},\left\lVert\widehat{\nu}_{b}(\mathcal{S})-\widehat{\nu}_{b}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert_{2}\right\}.$
Hence, by (106),
$\left|\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S})\right\rVert_{2}-\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S}_{g^{*}}^{(j)})\right\rVert_{2}\right|\leq\max_{{g^{*}}\in\mathcal{G}}\frac{1}{n_{g^{*}}}\left(\left\lVert{X_{j}^{({g^{*}})}}^{\prime}\right\rVert_{2}\Big|{Y_{j}^{({g^{*}})}}^{\prime}\Big|+\left\lVert{X_{j}^{({g^{*}})}}\right\rVert_{2}\Big|{Y_{j}^{({g^{*}})}}\Big|\right).$
A.5.9 Control of $S_{\lambda,\lambda^{\prime},u}(\mathcal{S})-S_{\lambda,\lambda^{\prime},u}(\mathcal{S}_{g^{*}}^{(j)})$:
For notational convenience, here we write $\mathcal{S}^{\prime}\coloneq\mathcal{S}_{g^{*}}^{(j)}$ and $R_{g}^{\mathcal{S}}(\beta)\coloneq\mathcal{R}_{g}(P_{n}(\mathcal{S}),\beta)$. Define
$\Delta_{g}^{(1)}\coloneq\left(R_{g}^{\mathcal{S}}(\widehat{\beta}_{\lambda}(\mathcal{S}))-R_{g}^{\mathcal{S}}(\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}))\right)-\left(R_{g}^{\mathcal{S}^{\prime}}(\widehat{\beta}_{\lambda}(\mathcal{S}))-R_{g}^{\mathcal{S}^{\prime}}(\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}))\right),$
$\Delta_{g}^{(2)}\coloneq\left(R_{g}^{\mathcal{S^{\prime}}}(\widehat{\beta}_{\lambda}(\mathcal{S}))-R_{g}^{\mathcal{S}^{\prime}}(\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}))\right)-\left(R_{g}^{\mathcal{S}^{\prime}}(\widehat{\beta}_{\lambda}(\mathcal{S}^{\prime}))-R_{g}^{\mathcal{S}^{\prime}}(\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}^{\prime}))\right),$
so that we can write
$S_{\lambda,\lambda^{\prime},u}(\mathcal{S})-S_{\lambda,\lambda^{\prime},u}(\mathcal{S}_{g^{*}}^{(j)})=\sum_{g\in\mathcal{G}}u_{g}\left(\Delta_{g}^{(1)}+\Delta_{g}^{(2)}\right).$
Note that $\Delta_{g}^{(1)}=0$ for $g\neq g^{*}$ since $\mathcal{S}$ and $\mathcal{S}^{\prime}$ on group $g$. For brevity, we drop the $g^{*}$ notation in what follows and write $X_{j},Y_{j}$ for samples in $\mathcal{S}$ and $\tilde{X_{j}},\tilde{Y_{j}}$ for the i.i.d. replacement sample in $\mathcal{S}^{\prime}$. Also let
$\Delta\beta\coloneq\widehat{\beta}_{\lambda}(\mathcal{S})-\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}),\quad q_{j}\coloneq Y_{j}-X_{j}^{\top}\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}),\quad\tilde{q}_{j}\coloneq\tilde{Y}_{j}-\tilde{X}_{j}^{\top}\widehat{\beta}_{\lambda^{\prime}}(\mathcal{S}).$
An expansion gives
$\Delta_{g^{*}}^{(1)}=\frac{1}{n_{g^{*}}}\left[\Delta\beta^{\top}\left(X_{j}X_{j}^{\top}-\tilde{X}_{j}\tilde{X}_{j}^{\top}\right)\Delta\beta-2\left(q_{j}X_{j}^{\top}-\tilde{q}_{j}\tilde{X}_{j}^{\top}\right)\Delta\beta\right].$
Hence, by Cauchy-Schwarz and the triangle inequality,
$|\Delta^{(1)}_{g^{*}}|\leq\frac{1}{n_{g^{*}}}\left[\left\lVert\Delta\beta\right\rVert^{2}\left(\left\lVert X_{j}\right\rVert^{2}+\left\lVert\tilde{X}_{j}\right\rVert^{2}\right)+2\left(|q_{j}|\left\lVert X_{j}\right\rVert+|\tilde{q}_{j}|\left\lVert\tilde{X}_{j}\right\rVert\right)\left\lVert\Delta\beta\right\rVert\right].$
By (99), (102), and using that $|\lambda-\lambda^{\prime}|\leq 1$, on $\mathcal{A}(t)$,
$|\Delta^{(1)}_{g^{*}}|\leq\frac{|\lambda-\lambda^{\prime}|}{n_{g^{*}}}A_{g^{\star},j}(\mathcal{S},\mathcal{S}^{\prime}),$
$A_{g^{\star},j}(\mathcal{S},\mathcal{S}^{\prime})$ $\coloneq{\kappa^{*}}^{2}(\left\lVert X_{j}\right\rVert^{2}+\|\tilde{X}_{j}\|^{2})$
$\qquad\qquad+2{\kappa^{*}}\left(|Y_{j}|\left\lVert X_{j}\right\rVert+|\tilde{Y}_{j}|\left\lVert X_{j}\right\rVert+\kappa_{t}\sup_{\xi\in[0,1]}\left\lVert\widehat{\nu}_{\xi}(\mathcal{S})\right\rVert\left(\left\lVert X_{i}\right\rVert+\left\lVert X_{j}\right\rVert\right)^{2}\right).$
On $\mathcal{A}(t)\cap\mathcal{E}$, recalling (104), one can show
[…]
Thus, on $\mathcal{A}(t)\cap\mathcal{E}$,
$|\Delta^{(1)}_{g^{*}}|\leq\frac{|\lambda-\lambda^{\prime}|}{n_{g^{*}}}\left(16\kappa_{t}T_{X}^{2}T_{Y}^{2}+48T_{X}^{4}T_{Y}^{2}+48\kappa_{t}^{3}T_{X}^{6}T_{Y}^{2}+16\kappa_{t}^{4}T_{X}^{8}T_{Y}^{2}\right).$ (111)
Next, we bound the second term $\Delta^{(2)}_{g^{*}}$. Fix $\xi\in[0,1]$. For the resampled dataset $\mathcal{S}^{\prime}$, define
$f_{g}(\beta)\coloneq\mathcal{R}_{g}(P_{n}(\mathcal{S}^{\prime}),\beta),\quad H_{g}(\xi)=f_{g}(\widehat{\beta}_{\xi}(\mathcal{S}))-f_{g}(\widehat{\beta}_{\xi}(\mathcal{S}^{\prime})),$
and see that, by the mean-value theorem,
$\Delta^{(2)}_{g^{*}}=H_{g}(\lambda)-H_{g}(\lambda^{\prime})=(\lambda-\lambda^{\prime})H^{\prime}_{g}(\xi)$ (112)
for some $\xi\in[\lambda,\lambda^{\prime}]$. It suffices to bound $H^{\prime}_{g}(\xi)$ uniformly in $\xi\in[0,1]$. Since $\Delta_{\beta}f_{g}(\beta)=2(\widehat{\Sigma}_{g}^{(j)}-\widehat{\nu}_{g}^{(j)})$,
$|H_{g}^{\prime}(\xi)|$ $\leq 2\left\lVert\widehat{\Sigma}_{g}^{(j)}\right\rVert\left\lVert\widehat{\beta}_{\xi}(\mathcal{S})-\widehat{\beta}_{\xi}(\mathcal{S}^{\prime})\right\rVert\left\lVert\partial_{\xi}\widehat{\beta}_{\xi}(\mathcal{S})\right\rVert$ (113)
$\qquad\qquad+2\left(\left\lVert\widehat{\Sigma}_{g}^{(j)}\right\rVert\left\lVert\widehat{\beta}_{\xi}(\mathcal{S}^{\prime})\right\rVert+\left\lVert\widehat{\nu}_{g}^{(j)}\right\rVert\right)\left\lVert\partial_{\xi}\widehat{\beta}_{\xi}(\mathcal{S})-\partial_{\xi}\widehat{\beta}_{\xi}(\mathcal{S}^{\prime})\right\rVert.$
By (103), (99), and (104), on $\mathcal{A}(t)\cap\mathcal{E}$,
$\left\lVert\partial_{\xi}\widehat{\beta}_{\xi}(\mathcal{S})\right\rVert$ $\leq 2\kappa_{t}T_{X}T_{Y}+2\kappa_{t}^{2}T_{X}^{3}T_{Y}.$ (114)
Using the identity $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$ and the triangle inequality, the truncation bounds (103), (99), (104), (109), and the single-sample perturbation bounds (107), we have that, on $\mathcal{A}(t)\cap\mathcal{E}$,
$\left\lVert\partial_{\xi}\widehat{\beta}_{\xi}(\mathcal{S})-\partial_{\xi}\widehat{\beta}_{\xi}(\mathcal{S}^{\prime})\right\rVert\leq\frac{1}{n_{g^{*}}}\left(2\kappa_{t}T_{X}T_{Y}+10\kappa_{t}^{2}T_{X}^{3}T_{Y}+12\kappa_{t}^{3}T_{X}^{5}T_{Y}\right).$ (115)
Define
$A^{*}(T_{X},T_{Y},\kappa_{t})\coloneq 16\kappa_{t}T_{X}^{2}T_{Y}^{2}+48\kappa_{t}^{2}T_{X}^{4}T_{Y}^{2}+48\kappa_{t}^{3}T_{X}^{6}T_{Y}^{2}+16\kappa_{t}^{4}T_{X}^{8}T_{Y}^{2},$
$P^{*}(T_{X},T_{Y},\kappa_{t})\coloneq 4\kappa_{t}T_{X}^{2}T_{Y}^{2}+32\kappa_{t}^{2}T_{X}^{4}T_{Y}^{2}+52\kappa_{t}^{3}T_{X}^{6}T_{Y}^{2}+28\kappa_{t}^{4}T_{X}^{8}T_{Y}^{2}.$
Simplifying and using (112), on $\mathcal{A}(t)\cap\mathcal{E}$,
$|\Delta^{(2)}_{g^{*}}|\leq|\lambda-\lambda^{\prime}|\frac{1}{n_{g^{*}}}P^{*}(T_{X},T_{Y},\kappa_{t}).$
Use $g^{\diamond}$ to denote the group that is not $g^{*}$ and observe $|u_{g^{*}}|,|u_{g^{\diamond}}|\leq\left\lVert u\right\rVert$. Combining the term-wise bounds, on $\mathcal{A}(t)\cap\mathcal{E}$,
$\left|S_{\lambda,\lambda^{\prime},u}(\mathcal{S})-S_{\lambda,\lambda^{\prime},u}(\mathcal{S}^{\prime})\right|\leq\frac{|\lambda-\lambda^{\prime}|}{n_{g^{*}}}\left\lVert u\right\rVert\left(A^{*}+2P^{*}\right).$ (116)
A.5.10 Bounded differences
Fix $\lambda,\lambda^{\prime}\in[0,1]$ and $u\in\mathbb{R}^{d}$. Let $f(\mathcal{S})\coloneq S_{\lambda,\lambda^{\prime},u}(\mathcal{S})$. Order the $i\in[n]$ (where $n=n_{r}+n_{b}$) samples $Z_{i}\coloneq(X_{i}^{(r)},Y_{i}^{(r)})$. Write $\mathcal{S}^{(j)}$ for the sample with $Z_{j}$ replaced by an i.i.d. copy. Define the data-dependent bounded difference radii:
$c_{i}(\mathcal{S})\coloneq\sup_{z,z^{\prime}}|f(Z_{1:i-1},z,Z_{i+1:n})-f(Z_{1:i-1},z^{\prime},Z_{i+1:n})|.$
Let $\mathcal{F}_{i}\coloneq\sigma(Z_{1:i})$, $g_{i}\coloneq\mathbb{E}\left[f((Z_{1:n}))|\mathcal{F}_{i}\right],Q_{i}\coloneq g_{i}-g_{i-1}$, and $g_{0}\coloneq\mathbb{E}[f(Z_{1:n})]$. Then, $\{g_{i}\}_{i=0}^{n}$ is the Doob martingale of $f$ with martingale differences $\{Q_{i}\}_{i=1}^{n}$ satisfying
$\mathbb{E}[Q_{i}|\mathcal{F}_{i-1}]=0,\qquad\sum_{i=1}^{n}Q_{i}=f(\mathcal{S})-\mathbb{E}[f(\mathcal{S})].$ (117)
A.5.11 Increment control
Define the endpoints
$A_{i}\coloneq\inf_{z}\left\{g_{i}(Z_{1:i-1},z)-g_{i-1}(Z_{1:i-1})\right\}\quad B_{i}\coloneq\sup_{z}\left\{g_{i}(Z_{1:i-1},z)-g_{i-1}(Z_{1:i-1})\right\}$
so that $Q_{i}\in[A_{i},B_{i}]$ almost surely. To bound the length of the interval, introduce the i.i.d sample $Z_{i+1:n}^{\prime}\coloneq(Z_{i+1}^{\prime},\ldots,Z_{n}^{\prime})$ independent of $Z_{1:i}$. Then, by independence,
$B_{i}-A_{i}\leq\mathbb{E}\left[\sup_{a,b}\left|f(Z_{1:i-1},b,Z^{\prime}_{i+1:n})-f(Z_{1:i-1},a,Z^{\prime}_{i+1:n})\right||\mathcal{F}_{i-1}\right]\leq c_{i}(Z_{-i}).$
By Hoeffding’s lemma, for $\gamma>0$,
$\mathbb{E}\left[\exp\left(\gamma\sum_{i=1}^{n}Q_{i}\right)\right]\leq\exp\left(\frac{\gamma^{2}}{8}\sum_{i=1}^{n}c_{i}(Z_{-i})^{2}\right).$ (118)
By (117), we obtain
$\mathbb{E}\left[\exp\left(\gamma\left(f(\mathcal{S})-\mathbb{E}[f(\mathcal{S})]\right)\right)\right]\leq\exp\left(\frac{\gamma^{2}}{8}\sum_{i=1}^{n}c_{i}(Z_{-i})^{2}\right).$ (119)
Index coordinates by $(g,j)$ for $j\in[n_{g}]$ for each $g\in\mathcal{G}$. On $\mathcal{A}(t)\cap\mathcal{E}$, by (116),
$c_{(g,j)}\leq\left\lVert u\right\rVert_{2}|\lambda-\lambda^{\prime}|\frac{1}{n_{g}}\left(A^{*}+2P^{*}\right).$
Consequently, on the same event,
$\sum_{i=1}^{n}c_{i}^{2}=\sum_{g\in\mathcal{G}}\sum_{j=1}^{n_{g}}c_{(g,j)}^{2}\leq\left\lVert u\right\rVert_{2}^{2}|\lambda-\lambda^{\prime}|^{2}(A^{*}+2P^{*})^{2}\left(\frac{1}{n_{r}}+\frac{1}{n_{b}}\right).$
Thus, by (119), on the same event,
$\mathbb{E}\left[\exp\left(\gamma\left(S_{\lambda,\lambda^{\prime},u}(\mathcal{S})-\mathbb{E}[S_{\lambda,\lambda^{\prime},u}(\mathcal{S})]\right)\right)\right]\leq\exp\left(\frac{\gamma^{2}}{8}\left\lVert u\right\rVert_{2}^{2}|\lambda-\lambda^{\prime}|^{2}(A^{*}+2P^{*})^{2}\left(\frac{1}{n_{r}}+\frac{1}{n_{b}}\right)\right).$ (120)
A.5.12 Centered subgaussian increments
We can write
$\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle=S_{\lambda,\lambda^{\prime},u}(\mathcal{S})-\mathbb{E}[S_{\lambda,\lambda^{\prime},u}(\mathcal{S})]+B_{\lambda,\lambda^{\prime},u}.$
where the deterministic bias term is
$B_{\lambda,\lambda^{\prime},u}\coloneq\mathbb{E}[S_{\lambda,\lambda^{\prime},u}(\mathcal{S})]-D_{\lambda,\lambda^{\prime},u}.$
By (120), for all $\lambda,\lambda^{\prime}\in[0,1]$ and all $u\in\mathbb{R}^{2}$,
$\mathbb{E}\left[\exp\left(\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle-\mathbb{E}\left[\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle\right]\right)\mid\mathcal{A}(t)\cap\mathcal{E}\right]\leq\exp\left(\frac{K_{\Delta}^{2}|\lambda-\lambda^{\prime}|^{2}}{2}\left\lVert u\right\rVert_{2}^{2}\right)$ (121)
where
$K_{\Delta}\coloneq\sqrt{\frac{1}{n_{r}}+\frac{1}{n_{b}}}\left(12\kappa_{t}T_{X}^{2}T_{Y}^{2}+56\kappa_{t}^{2}T_{X}^{4}T_{Y}^{2}+76\kappa_{t}^{3}T_{X}^{6}T_{Y}^{2}+36\kappa_{t}^{4}T_{X}^{8}T_{Y}^{2}\right).$
By (98), Equation 97, and De Morgan’s Law,
$\mathbb{P}\left((\mathcal{A}(t)\cap\mathcal{E})^{C}\right)\leq n^{-C\alpha}+\left(C_{r}^{\prime}t\right)^{\alpha_{r}n_{r}/6}+\left(C_{b}^{\prime}t\right)^{\alpha_{b}n_{b}/6}.$ (122)
All together, for any $u\in\mathbb{R}^{2}$, $\gamma\in\mathbb{R}$, and $\lambda,\lambda^{\prime}\in[0,1]$,
$\mathbb{E}\left[\exp\left(\gamma\left(\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle-\mathbb{E}\left[\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle\right]\right)\right)\mid\mathcal{A}(t)\cap\mathcal{E}\right]\leq\exp\left(\gamma^{2}\frac{(\lambda-\lambda^{\prime})^{2}}{2}\left\lVert u\right\rVert_{2}^{2}K_{\Delta}^{2}\right).$
Next, we obtain a uniform bound for the bias term $B_{\lambda,\lambda^{\prime},u}$. Let $b_{\lambda,g}\coloneq\mathbb{E}\left[\mathcal{R}_{g}(P_{n},\widehat{\beta}_{\lambda})-\mathcal{R}_{g}(P,\beta_{\lambda})\right]$ and observe
$|B_{\lambda,\lambda^{\prime},u}|=\sum_{g\in\mathcal{G}}|u_{g}|\left(|b_{\lambda,g}|+|b_{\lambda^{\prime},g}|\right)\leq 2\sum_{g\in\mathcal{G}}|u_{g}|\cdot\mathbb{E}\left[\sup_{\beta\in B_{2}(B)}\left|\mathcal{R}_{g}(P_{n},\beta)-\mathcal{R}_{g}(P,\beta)\right|\right].$
By the triangle inequality and Cauchy-Schwarz,
$\sup_{\beta\in B_{2}(B)}\left|\mathcal{R}_{g}(P_{n},\beta)-\mathcal{R}_{g}(P,\beta)\right|\leq 4B^{2}\left\lVert\Sigma_{g}-\widehat{\Sigma}_{g}\right\rVert+4B\left\lVert\frac{1}{n_{g}}\sum_{i:G_{i}=g}X_{i}^{(g)}\varepsilon_{i}^{(g)}\right\rVert+\left|{\frac{1}{n_{g}}\sum_{i:G_{i}=g}\left((\varepsilon_{i}^{(g)})^{2}-\sigma^{2}\right)}\right|.$
By standard subgaussian design sample covariance bounds (e.g., by integrating the tail bound as in Proof of Lemma 4), there exists a constant $C>0$ such that
$\mathbb{E}\left\lVert\Sigma_{g}-\widehat{\Sigma}_{g}\right\rVert\leq CK_{g}^{2}\left\lVert\Sigma_{g}\right\rVert\left(\sqrt{\frac{d}{n_{g}}}+\frac{d}{n_{g}}\right).$
Observing that, for all $i,g$, $(\varepsilon_{i}^{(g)})^{2}-\sigma^{2}$ is a centered subexponential random variable satisfying $\left\lVert(\varepsilon_{i}^{(g)})^{2}\right\rVert_{\psi_{1}}\leq K_{\varepsilon}^{2}\sigma^{2}$, by standard centering results [see, e.g., vershynin2018high, Exercise 2.7.10], there exists a constant $C>0$ such that $\left\lVert(\varepsilon_{i}^{(g)})^{2}-\sigma^{2}\right\rVert_{\psi_{1}}\leq CK_{\varepsilon}^{2}\sigma^{2}$. Thus, by Bernstein’s inequality [as in vershynin2018high, Corollary 2.8.3], there exists a constant $C>0$ such that
$\mathbb{E}\left|\frac{1}{n_{g}}\sum_{i:G_{i}=g}\left((\varepsilon_{i}^{(g)})^{2}-\sigma^{2}\right)\right|\leq C\frac{K_{\varepsilon}^{2}\sigma^{2}}{\sqrt{n_{g}}}.$
Similarly, observing $X_{i}^{(g)}\varepsilon_{i}^{(g)}$ is also subexponential, by a $\frac{1}{2}$-net argument, there exists a constant $C>0$ such that
$\mathbb{E}\left\lVert\frac{1}{n_{g}}\sum_{i:G_{i}=g}X_{i}^{(g)}\varepsilon_{i}^{(g)}\right\rVert\leq CK_{g}K_{\varepsilon}\sigma\left\lVert\Sigma_{g}^{1/2}\right\rVert\sqrt{\frac{d}{n_{g}}}.$
All together, there exists a constant $C>0$ such that, if we define for each $g\in\mathcal{G}$
$\Gamma_{g}\coloneq 4B^{2}CK_{g}^{2}\left\lVert\Sigma_{g}\right\rVert\left(\sqrt{\frac{d}{n_{g}}}+\frac{d}{n_{g}}\right)+4BCK_{g}K_{\varepsilon}\sigma\left\lVert\Sigma_{g}^{1/2}\right\rVert\sqrt{\frac{d}{n_{g}}}+CK_{\varepsilon}^{2}\sigma^{2}\frac{1}{\sqrt{n_{g}}},$
then
$\sup_{\beta\in B_{2}(B)}\left|\mathcal{R}_{g}(P_{n},\beta)-\mathcal{R}_{g}(P,\beta)\right|\leq\Gamma_{g}.$ (123)
Thus,
$\sup_{\lambda,\lambda^{\prime}\in[0,1]}|B_{\lambda,\lambda^{\prime},u}|\leq 2(\Gamma_{r}+\Gamma_{b})\left\lVert u\right\rVert_{2}.$
A.5.13 Uniform Control via Dudley’s Inequality
Take unit $u$. By Dudley’s tail bound with index set $[0,1]$ and metric $d(\lambda,\lambda^{\prime})=|\lambda-\lambda^{\prime}|$ [see, e.g., vershynin2018high, Theorem 8.1.6], for any $\delta_{0}>0$,
$\mathbb{P}\left(\sup_{\lambda,\lambda^{\prime}\in[0,1]}\left|\langle u,\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\rangle\right|\leq C_{\Delta}K_{\Delta}\left(1.26+\delta_{0}\right)+2(\Gamma_{r}+\Gamma_{b})\mid\mathcal{A}(t)\cap\mathcal{E}\right)\geq 1-2\exp(-\delta_{0}^{2}).$
where $\mathcal{N}$ denotes the covering number. Note the metric entropy bound
$\int_{0}^{1}\sqrt{\log\mathcal{N}([0,1],|\cdot|,\varepsilon)}d\varepsilon\leq\int_{0}^{1}\sqrt{\log(2/\varepsilon)}d\varepsilon\leq 1.26.$
Next, as in the proof of Equation 97, there exists an $1/2$-net $\mathcal{M}$ satisfying $|\mathcal{M}|\leq 25$ over the two-dimensional unit sphere. Apply the inequality $\left\lVert z\right\rVert_{2}\leq 2\sup_{v\in\mathcal{M}}\langle z,v\rangle$ [see vershynin2018high, Exercise 4.4.2] to $z=\Delta_{\lambda}-\Delta_{\lambda^{\prime}}$ and union bound over $v\in\mathcal{M}$ to obtain
$\mathbb{P}\left(\sup_{\lambda,\lambda^{\prime}\in[0,1]}\left\lVert\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\right\rVert\leq C_{\Delta}K_{\Delta}\left(1.26+\delta_{0}\right)+2(\Gamma_{r}+\Gamma_{b})\mid\mathcal{A}(t)\cap\mathcal{E}\right)\geq 1-2|\mathcal{M}|\exp(-\delta_{0}^{2}).$
Fix $\delta_{1}\in(0,1)$ and take $\delta_{0}=\sqrt{\log\left(2|\mathcal{M}|/\delta_{1}\right)}$. For any $\delta_{1}\in(0,1)$,
$\mathbb{P}\left(\sup_{\lambda,\lambda^{\prime}\in[0,1]}\left\lVert\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\right\rVert\leq C_{\Delta}K_{\Delta}\left(1.26+\sqrt{\log\left(50/\delta_{1}\right)}\right)+2(\Gamma_{r}+\Gamma_{b})\mid\mathcal{A}(t)\cap\mathcal{E}\right)\geq 1-\delta_{1}.$ (124)
A.5.14 Anchoring increments through a choice of $\lambda_{0}$
Fix $\lambda_{0}\in[0,1]$ and unit vector $u$. Define
$S_{\lambda_{0},u}(\mathcal{S})\coloneq\sum_{g\in\mathcal{G}}u_{g}\mathcal{R}_{g}\left(P_{n},\widehat{\beta}_{\lambda_{0}}(\mathcal{S})\right),\quad\Delta_{\lambda_{0}}\coloneq\widehat{\mathcal{R}}_{\lambda_{0}}-{\mathcal{R}}_{\lambda_{0}}\in\mathbb{R}^{2}.$
By the same bounded difference and one-sample perturbation steps as we employed to control $S_{\lambda,\lambda^{\prime},u}$, we reach that, on the event $\mathcal{A}(t)\cap\mathcal{E}$,
$\mathbb{E}\left[\exp\left(\gamma\left(S_{\lambda_{0},u}(\mathcal{S})-\mathbb{E}[S_{\lambda_{0},u}(\mathcal{S})]\right)\right)\right]\leq\exp\left(\frac{\gamma^{2}}{8}\left\lVert u\right\rVert_{2}^{2}(A^{*}+2P^{*})^{2}\left(\frac{1}{n_{r}}+\frac{1}{n_{b}}\right)\right).$
By the same $1/2$-net argument as before, for any $\delta_{2}\in(0,1)$, there exists a constant $C>0$ such that
$\mathbb{P}\left(\left\lVert\Delta_{\lambda_{0}}-\mathbb{E}[\Delta_{\lambda_{0}}|\mathcal{A}(t)\cap\mathcal{E}]\right\rVert\leq CK_{\Delta}\sqrt{\log(50/\delta_{2})}\right)\geq 1-\delta_{2}.$
As for the increment terms, write
$\langle u,\Delta_{\lambda_{0}}\rangle=S_{\lambda_{0},u}(\mathcal{S})-\mathbb{E}\left[S_{\lambda_{0},u}(\mathcal{S})\right]+B_{\lambda_{0},u},\quad B_{\lambda_{0},u}\coloneq\mathbb{E}\left[S_{\lambda_{0},u}(\mathcal{S})\right]-D_{\lambda_{0},u},\quad D_{\lambda_{0},u}\coloneq\sum_{g\in\mathcal{G}}u_{g}\mathcal{R}_{g}(P,\beta_{\lambda_{0}}).$
Let
$b_{\lambda_{0},g}\coloneq\mathbb{E}\left[\mathcal{R}_{g}(P_{n},\widehat{\beta}_{\lambda_{0}}-\mathcal{R}_{g}(P,\beta_{\lambda_{0}})\right]=\mathbb{E}\left[\mathcal{R}_{g}(P,\widehat{\beta}_{\lambda_{0}})-\mathcal{R}_{g}(P,\beta_{\lambda_{0}})\right]+\mathbb{E}\left[\mathcal{R}_{g}(P_{n},\widehat{\beta}_{\lambda_{0}}-\mathcal{R}_{g}(P,\widehat{\beta}_{\lambda_{0}})\right].$
By Cauchy-Schwarz and (123), we can bound
$b_{\lambda_{0},g}\leq U_{\lambda_{0},g}+2\sqrt{U_{\lambda_{0},g}}\left\lVert\Sigma_{g}^{1/2}(\beta_{\lambda_{0}}-\beta_{g})\right\rVert_{2}+\Gamma_{g},\quad U_{\lambda_{0},g}\coloneq{\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda_{0}}-\beta_{\lambda_{0}}\right\rVert_{\Sigma_{g}}^{2}\mid\mathcal{A}(t)\cap\mathcal{E}\right]}.$
On $\mathcal{A}(t)$ and under the $B$-bounded parameter assumption,
$\left\lVert\widehat{\beta}_{\lambda_{0}}-\beta_{\lambda_{0}}\right\rVert_{\Sigma_{g}}\leq\kappa_{t}\left\lVert\Sigma_{g}^{1/2}\right\rVert\left(\lambda_{0}\left\lVert\widehat{\nu}_{r}-\nu_{r}\right\rVert+(1-\lambda_{0})\left\lVert\widehat{\nu}_{b}-\nu_{b}\right\rVert\right)+\kappa_{t}B\left\lVert\Sigma_{g}^{1/2}\right\rVert\left(\lambda_{0}\left\lVert\widehat{\Sigma}_{r}-\Sigma_{r}\right\rVert+(1-\lambda_{0})\left\lVert\widehat{\Sigma}_{b}-\Sigma_{b}\right\rVert\right).$
By the truncation bounds on $\mathcal{E}$ and Matrix Bernstein ([vershynin2018high, Theorem 5.6.1]) , there exist constants $C_{1},C_{2}>0$ such that
$\mathbb{E}\left[\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\right]\leq C_{1}T_{X}^{2}\left(\sqrt{\frac{d}{n_{g}}}+\frac{d}{n_{g}}\right),\quad\mathbb{E}\left[\left\lVert\widehat{\nu}_{g}-\nu_{g}\right\rVert\right]\leq C_{2}T_{X}T_{Y}\sqrt{\frac{d}{{n_{g}}}}.$
Hence, on $\mathcal{A}(t)\cap\mathcal{E}$, there exists a constant $C>0$ such that
$U_{\lambda_{0},g}\leq C\left\lVert\Sigma_{g}^{1/2}\right\rVert^{2}\kappa_{t}^{2}\left(T_{X}^{2}T_{Y}^{2}+B^{2}T_{X}^{4}\right)\frac{d}{n_{g}}.$ (125)
Since $|B_{\lambda_{0},u}|\leq\left\lVert u\right\rVert_{2}\left\lVert b_{\lambda_{0}}\right\rVert_{2}$ for unit $u$ we reach that, by (125), there exists a constant $C>0$ such that
$\Gamma_{\lambda_{0}}\coloneq\left\lVert b_{\lambda_{0}}\right\rVert\leq C\sum_{g\in\{r,b\}}\left(\kappa_{t}^{2}\left(T_{X}^{2}T_{Y}^{2}+B^{2}T_{X}^{4}\right)\left\lVert\Sigma_{g}^{1/2}\right\rVert^{2}\frac{d}{n_{g}}+B^{2}\left\lVert\Sigma_{g}^{1/2}\right\rVert^{2}(T_{X}T_{Y}+BT_{X}^{2})\sqrt{\frac{d}{n_{g}}}\right)+(\Gamma_{r}+\Gamma_{b}).$
Combining, we obtain that, for any $\delta_{2}\in(0,1)$,
$\mathbb{P}\left(\left\lVert\Delta_{\lambda_{0}}\right\rVert\leq CK_{\Delta}\sqrt{\log(50/\delta_{2})}+\Gamma_{\lambda_{0}}\mid\mathcal{A}(t)\cap\mathcal{E}\right)\geq 1-\delta_{2}.$ (126)
Relating increment and anchor bounds for uniform bound:
By the triangle inequality,
$\sup_{\lambda\in[0,1]}\left\lVert\Delta_{\lambda}\right\rVert_{2}\leq\left\lVert\Delta_{\lambda_{0}}\right\rVert_{2}+\sup_{\lambda,\lambda^{\prime}\in[0,1]}\left\lVert\Delta_{\lambda}-\Delta_{\lambda^{\prime}}\right\rVert_{2}.$
Fix $\delta\in(0,1)$ and choose $\delta_{1},\delta_{2}>0$ such that $\delta_{1}=\delta_{2}=\delta/2$. Apply (124) and (126) with $\delta_{1},\delta_{2}$, respectively and apply (122) to conclude that there exists a constant $C>0$
$\mathbb{P}\left(\sup_{\lambda\in[0,1]}\left\lVert\Delta_{\lambda}\right\rVert_{2}\leq\Gamma_{\lambda_{0}}+2(\Gamma_{r}+\Gamma_{b})+CK_{\Delta}\left(1.26+\sqrt{\log(100/\delta)}\right)\right)\geq 1-\delta-\eta_{t}.$